A Complete Bibliography of ACM Transactions on Reconfigurable Technology and Systems

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

14 June 2025 Version 1.54

Title word cross-reference

+ [GL08]. 2 [BPCC09, LP15, RNTW22]. 3 [IAG23, JB15, SPS12, TZWZ15, WZY^{+}24]. K [LLS^{+}23, TK16]. N [MPK22]. QR [ZCL16].

-body [MPK22]. -D [SPS12]. -Means [TK16]. -Nearest [LLS^{+}23].

10 [TKB^{+}24]. 11th [AC14]. 15th [DH08]. 16-nm [TMLS21]. 19th [GC13]. 1D [SB25].

2 [YB18]. 2.0 [AZM^{+}19, GMZ^{+}23, WZL24, ZLY^{+}24]. 2.1 [JRHK15]. 200 [WBR16]. 2007 [AN09]. 2009 [Che11, WBAM10]. 2011 [Hüb12]. 2013 [CDM15]. 2014 [BAG15, DB15, GSCB15, RVHP16, SB15]. 2015 [CS17]. 2019 [Deh20]. 2020 [MST22, Sha22]. 2021 [GKK24, KZ23]. 2022 [Ien23, KS24, LH23]. 2023 [FB24]. 256 [MAK^{+}12].

5 [AKA09]. 5.0 [LKJ^{+}11].

7.0 [LGW^{+}14]. 7nm [SY24]. 7th [VG14].

8 [MPZ^{+}20]. 80 [YSC^{+}23].

A-Port [PVA^{+}09]. Abstraction [IBH^{+}15]. Abstractions [IPC14]. Academic [MWL^{+}15]. ACAP [ZLY^{+}24]. Accelerated [GKA^{+}24, MHS^{+}19, MCC10, SKJ22, WZY^{+}24, BE19, GKLLA23, SM25]. Accelerating [JLB^{+}08, LMD^{+}25, LLZ^{+}24, PFL22, SDGL^{+}22, TZWZ15, VL11, ZG16, ZVS20]. Acceleration [ACFm^{+}25, ABB^{+}23, BMC^{+}22, CKRG^{+}24, CAPA^{+}09, CZD^{+}25, CSS^{+}23, CBR^{+}14, CZ09, DFB^{+}22, HNM^{+}22, KLC11, KZB23, LUX^{+}21, LLS^{+}24, LAF24, MCD^{+}18, NG25, PFC15, PBPLA17, RRLW22, SCL24, TBS^{+}25, TK16, TYL^{+}23, WTS19, WMG^{+}10, WWC^{+}22, WSZ^{+}25, XCG^{+}09, YOY17, YBS16, ZBR12]. Accelerator [BHXJ24, CNZ^{+}18, CZZ23, CJH^{+}24, DCL^{+}22, HLW^{+}21, JS24, JZL^{+}24, KVV24, LML^{+}23, LXS^{+}24, LDJ^{+}17, LLS^{+}23, LLYC24, MYST25, QMG^{+}24, QNF^{+}23, SZKR22, SKW^{+}21, TWL^{+}23, TTY25, WHJ^{+}23, XRC23, XJXQ24, YHK^{+}21, YEC^{+}09, YGH^{+}18, ZZJB13, ZQBL25, YXC^{+}11]. Accelerators [BHF^{+}25, BSW^{+}23, BBO^{+}25, DCS24, FC24, GFC^{+}25, GZY^{+}18, HBXA23, JRHK15, KHK24, LCL^{+}24, NGM^{+}23, SLL^{+}20, SKJ22, UNBR14, ZLY^{+}24]. Access [AEM^{+}24, LYZ^{+}18]. Accesses [PFC15]. Accumulation [LLL^{+}23]. Accumulator [WS16]. Accuracy [DHL^{+}18, KY18, LP15, UNBR14]. Accurate [AVCP20, CSK17, DLBM18, JM14, SLH^{+}25]. ACE [HLW^{+}21]. ACE-GCN [HLW^{+}21]. ACM [Lee23]. Across [DCS24]. ACTS [JS24]. Adaptable [MLPK22]. Adaptation [FL20]. Adaption [BHI15]. Adaptive [CNE^{+}15, GRNW22, INF^{+}14, JCG^{+}12, LSP^{+}23, NNY12, OVI^{+}12, PMC^{+}14, SGW20, Tak17, ZCL15, ZMH^{+}23, Tak12, DGP^{+}15]. ADAS [CZZ23]. Adder [PBBP18]. Adders [HU10]. Adding [PSM^{+}14]. Addition [CAPA^{+}09, OBD13]. Addition-Related [OBD13]. Adjustable [FYZ^{+}25, ZWM19]. Adjustment [NW11]. Advantage [MPK22]. Advantages [WSDH23]. adventure [RD11]. AEKA [BHXJ24]. Aerial [CZ09]. Aerospace [WGGR16]. AES [DGP10, HF14]. Against [SRR23, LOM10]. Agent [GMBC17]. Aggregation [GS23]. Aging [CAG^{+}22]. Agreement [ADSH18]. AI [JZL^{+}24]. AIgean [TDH^{+}22]. Ain’t [RNTW22]. Algebra [HZS^{+}25]. Algorithm [AIH^{+}24, CBR^{+}14, EWL15, KVV24, MS25, RLY^{+}15, Ste10, SMN^{+}23, TL11, TK16, ZCK22]. Algorithm-Hardware [AIH^{+}24]. Algorithm/Architecture [EWL15]. Algorithmic [ZVS20, ZW25]. Algorithmic-Level [ZW25]. Algorithms [CW09, LRA13, NSS^{+}11]. Alignment [JLB^{+}08, MCC10, OBD13]. All-digital [LLS^{+}24]. Altera [SMOP15, TK16]. Altering [WML^{+}24]. AMD [HDN^{+}24]. Amenability [HNG09]. Analog [HH24]. Analyses [DRHM22]. Analysis [BPFD11, CCF^{+}18, CFBS15, CKG^{+}10, JCGW20, LLL^{+}23, MMMT09, PRV21, PPR^{+}10, RGGW10, RGCL16, RMSK16, SB08, TMLS21, WBCW24, GP13, Tak12]. Analytical [DLZ^{+}24, KSCC10, LAL13, YCV^{+}21, DW13, HGLS11]. Analytics [JS24, SZKR22]. Analyzing [GSJC13]. Anomaly [LBL23]. Application [ABCC09, BBND10, CM14, DDB^{+}10, GdLlG^{+}14, JSC14, KGS15, LJS11, MLPK22, MWK^{+}12, PMKM11, RUC11, SSK^{+}23, SLL^{+}20, VTN09, WYZ16, WMG^{+}10, YFW^{+}17, SSF^{+}13]. Application-Optimized [YFW^{+}17]. Application-Specific [PMKM11, LJS11, SLL^{+}20]. Applications [AZM^{+}19, AI22, CFBS15, CKG^{+}10, DFB^{+}22, GKM^{+}12, HDMN24, IAG23, KBM09, KCC^{+}14, LZF^{+}10, LBRS16, NJLW14, PSM^{+}14, PVB13, RRW^{+}22, SGC21, SML^{+}24, WHQ^{+}08, KSG11]. Applying [NSS^{+}11]. Approach [CM14, DCS24, FKM^{+}25, GKGS23, KM10, LYZ^{+}18, MS23, MWK^{+}12, NBS13, RKV23, SBC15, WSDH23]. Approaches [CHG22, MVGB15, SAD10]. Approximate [FAB22, LN23, SUK24, VVZL25]. ARC [BAG15, DB15, GSCB15, SB15, WBAM10]. ARC’08 [CWBD09]. Architecture [BYB23, ADSH18, ATJZ16, BCE^{+}10, CXG^{+}12, DS15, EWL15, FT17, GWS23, GMBC17, GPL24, IZO^{+}10, IGM^{+}20, IF23, IBH^{+}15, KLD16, KSCC10, KAL14, KD21, LNGP22, LLS^{+}24, LYC^{+}24, LVC24, LGW^{+}14, MPZ^{+}20, OWMZ11, PFC15, PB18, RRLW22, RNTW22, SBC10, SB15, SZZW23, SCL24, SKB^{+}22, Tak17, VL11, WS16, WSZ^{+}25, XJD^{+}16, ZCL16, ZLY^{+}24, DW13, LKJ^{+}11, Oli12]. Architectures [BBND10, BDX^{+}19, CBC^{+}12, DSB09, EBYB20, GC13, IAG23, JTLC09, KY18, LAL13, LFN^{+}18, NG25, SFT^{+}23, WCK21, YB18, ZFCZ25, HLL08]. Area [BHXJ24, DD15, KY18, SY24, Tho15, WCK21]. Area-Efficient [BHXJ24, DD15, Tho15]. ARISE [VTN09]. Arithmetic [SUK24, SCC10, TMLS21, WWC^{+}22, WSZ^{+}25]. ARM [GHWS22, PCC^{+}25]. ARM-based [PCC^{+}25]. Array [BCW21, SLH^{+}10, ZCL16]. Array-System [BCW21]. Arrays [DPHT19, HNM^{+}22, SCC10, ZCK22, ZH12]. Art [LCL^{+}24]. Artifact [Lee23]. Artificial [KAL14]. ASIC [BYB18, DE22]. ASIP [EWL15]. Assembly [BGSL17]. Assignment [SB08]. Associative [DD15]. Assurance [CHG22, KMK^{+}10]. Asymmetric [SDG12]. Atmospheric [GFL^{+}15]. Attack [SGM09]. Attacks [GER19, GTS23, KGT19, MDL^{+}23, TJCP25, ZQ19]. Authenticated [ADSH18]. Automata [BDX^{+}19, KB23, KD21, MHS09]. Automated [BBO^{+}25, CCZ^{+}24, DD18, GMZ^{+}23, HDMN24, LAF24, RMSK16, SCC10, XJXQ24]. Automatic [AZM^{+}19, APR^{+}22, SFT^{+}23, TYL^{+}23, YB18, YBS16]. Automatically [LP15]. Automating [NCJ^{+}15, YFW^{+}17]. Automation [KTB^{+}24, SV09]. Autonomous [BMR16, DVK15]. AutoScaleDSE [JYJC23]. Avionics [LZF^{+}10]. avoidance [RD11]. Aware [BAG15, BKT14, DLZ^{+}24, HNS^{+}10, LCS14, LN23, NJLW14, SB08, TTY25, CMT^{+}24, EA11, KSG11, ZVS20]. Awareness [AHL^{+}14, Bec14, GKLLA23, DGP^{+}15]. AWS [ES22]. AxOMaP [SUK24].

Bandwidth [AI22, BBND10, HNM^{+}22, SLH^{+}10, SXZ^{+}25a, USY17, BC11, SFT^{+}23, ZZB^{+}20]. Bandwidth-Bound [AI22]. Bandwidth-Reduction [SLH^{+}10]. Based [AL16, AIH^{+}24, BAG15, BHXJ24, CBFM14, CZ09, DGP^{+}15, DCL^{+}22, DLZ^{+}24, DL09, EWL15, FHL^{+}24, FYZ^{+}25, FZ25, GWPK20, GDHG11, GHO17, HLN^{+}10, JCG^{+}12, JTLC09, Kap16, KBT09, KD10, KGS^{+}12, LBRS16, LYC^{+}24, LZ19, LCL^{+}24, LT09, LVC24, LYH^{+}23, LL12, MVGB15, MWBL21, MZLS20, NNY12, OVI^{+}12, PRV21, PPR^{+}10, RKV23, RC10, SY24, SLH^{+}10, SB15, SP20, SCY^{+}23, SXZ^{+}25a, SXZ^{+}25b, TYB18, USY17, WAS^{+}23, WZY^{+}24, WGGR16, YOY17, ZCL16, ZBC^{+}09, ZNA^{+}18, ZBB^{+}16, AHAM^{+}19, AB23, BHF^{+}25, CZD^{+}25, CAG^{+}22, DLCJ20, EA11, FZ23, GWS23, GKGS23, GWXW21, GZY^{+}18, HBXA23, HLL08, KZB23, LZF^{+}10, LSP^{+}23, LZZ^{+}24, LFN^{+}18, LLS^{+}23, LLYC24, LBL23, MLFS22, MKSB22, MBJJ11, MS25, NGM^{+}23, PCC^{+}25, SUK24, SKW^{+}21, SLL^{+}20, SML^{+}24, Ste10, TWL^{+}23, TJCP25, TYL^{+}23, TTY25, TKB^{+}24, WTS19, WWC^{+}22, YXC^{+}11, YHK^{+}21, ZBR12, ZQ19, ZZB^{+}20, ZFCZ25, ZQBL25, KP14]. Based [TLC^{+}24, UNBR14, ZZJB13]. Bayesian [MYST25]. Behavior [PVA^{+}09]. Benchmarks [HH24, MKP23, MWL^{+}15, PB18]. Benefits [PSM^{+}14]. Bent [SZZW23]. BETA [TTY25]. between [LW08, MWL^{+}15, TOS17]. BFS [LXS^{+}24]. Big [RMSK16]. Binary [ADSH18, BHXJ24, BHF^{+}25, FAB22, HBXA23, PFC15, SCL24, YLF^{+}24]. Binary-Unary [FAB22]. Biomedical [KCC^{+}14, YBS16]. BISWSRBS [YHK^{+}21]. Bit [TTY25, UCR^{+}19, LYH^{+}23]. Bit-Level [TTY25]. Bit-Serial [UCR^{+}19]. Bits [DVK15, GKGS23]. Bitstream [BPFD11, HTG25, SMOP15]. Bitwidth [VDdSN23]. BlastFunction [DFB^{+}22]. BLASTP [JLB^{+}08, MH15]. Block [BDGH15, CBFM14, DE22, EBYB20, TYB18, ZCK22, ZQBL25]. Block-Based [CBFM14]. Blockchain [ILP24]. Blocks [AGM^{+}22, FK08, NI24, RRW^{+}22, SY24, TMLS21, PMKM11, dBVW^{+}24]. BLOOP [FZ23]. BNN [DCL^{+}22]. body [MPK22]. Boltzmann [KAL14]. Boolean [FZ23]. Bottleneck [SKJ22, KSG11]. Bound [AI22, MHS09, RLM^{+}17]. BRAM [AZM^{+}19]. BRAMs [DGP10]. Branch [RLM^{+}17]. Bringing [DFB^{+}22]. Broadcast [PSM^{+}14]. Buffer [HDMN24, JSG^{+}22]. Build [BF23]. Building [AGM^{+}22, DE22]. BurstZ [SKJ22]. Buses [HBA^{+}15].

C [ES22]. C/C [ES22]. Cache [AI22, KD19, SDG12]. Cache-coherent [KD19]. Cache-Unfriendly [AI22]. Caching [SLL^{+}20]. CAD [KA17, LKJ^{+}11, LGW^{+}14, MWL^{+}15, MPZ^{+}20]. calculation [DCL^{+}22]. Calculus [GWPK20]. Calculus-Based [GWPK20]. Calibration [TKB^{+}24]. Calling [XRC23]. CAM [SCL24]. Canalis [CNK^{+}24]. Cannot [BYB18]. Capabilities [GFBF12, HCGG24]. Capability [LZF^{+}10]. Capable [BMR16, KA21]. care [MBJJ11]. Carlo [TB10]. Carry [PABI09]. Cartography [SGM09]. Case [SFT^{+}23, NSS^{+}11]. CDO [KLC11]. Cell [IZO^{+}10, KA17, PABI09]. Cellular [KD21, MHS09]. Centers [ENPR22a, ENPR22b, KW22]. Centric [BFBN^{+}20, VG14]. CGRA [VL11, CCZ^{+}24, TJCP25, WCK21, dBVW^{+}24]. CGRA-based [TJCP25]. CGRA-EAM-Rapid [WCK21]. CGRAs [CM14, MS25, NZS^{+}23, WSDH23]. Chain [PABI09]. Challenge [MKP23]. Challenges [CLL^{+}22, NG25]. Channel [DSW^{+}24, GTS23, SG15, TJCP25, GER19]. channels [TES24]. Characterization [WMG^{+}10]. CHARM [ZLY^{+}24]. Check [DL09]. Checking [CPFM21, PD15]. Checkpointing [AB20, AB23]. Chief [Che19]. Chip [AB14, BCW21, CTH16, CSK17, DE22, GMBC17, GdLlG^{+}14, GS10, HBA^{+}15, JSC14, LXS^{+}24, LL12, VG14, ZW25, GNM^{+}15, LLS^{+}23]. CHIP-KNNv2 [LLS^{+}23]. CHIRP [BHF^{+}25]. Choose [RD11]. Choose-your-own-adventure [RD11]. Cibola [QRDC^{+}15]. Cipher [SCY^{+}23, YSC^{+}23]. Circuit [AHSS^{+}21, DVH^{+}15, EAAAA19, KS20, LL12, MKP23, US23, WSC09, GL08]. Circuit-Switched [LL12, MKP23]. Circuit-Switching [US23]. Circuitry [GGR^{+}18]. Circuits [BMR16, CBC^{+}12, CAG^{+}22, DL09, FC24, FZ25, GRNW22, JSG^{+}22, JMGI23, LSP^{+}23, SC08, SV09, Ste10, WBR16, WBR18]. Classification [MS23, ZWM19]. Clock [GRNW22, WSDH23, LW08]. Clocking [DB15]. Cloud [AKM^{+}25, BDX^{+}19, BE19, BMC^{+}22, DFB^{+}22, DE22, ES22, GTS23, KZB23, KGT19, LSP^{+}23, MKSB22, SM25, ZDS^{+}22, EAAAA19]. Cloud-native [DFB^{+}22]. Clouds [APR^{+}22, DE22, RHLK18]. Cluster [US23, GNM^{+}15]. Clustering [LRA13, ZMH^{+}23, EA11]. Clusters [FK08, OKA19, TDH^{+}22]. CNN [BSW^{+}23, JZL^{+}24, LFN^{+}18, MCD^{+}18, MZLS20, WWC^{+}22, YHK^{+}21, YGH^{+}18]. CNN-based [LFN^{+}18]. CNNs [dMdLC23]. Co [AIH^{+}24, AB23, EWL15, GCL^{+}23, MCL^{+}23, SMN^{+}23, ZHL^{+}21]. Co-Design [AIH^{+}24, ZHL^{+}21, MCL^{+}23]. Co-Exploration [EWL15]. Co-optimization [GCL^{+}23, SMN^{+}23]. Co-Simulation [AB23]. Coalesce [AI22]. Coarse [VL11, WCK21, XJD^{+}16]. Coarse-Grained [VL11, XJD^{+}16, WCK21]. Code [DVH^{+}15, DC16, ES22, GRG08, ZG16]. Codesign [JSLO24, CMT^{+}24, SC11]. Coding [BAG15]. Coefficient [FAB22]. CoEx [EWL15]. COFFE [YB18]. Coherency [SDG12]. Coherent [PRV21, KD19]. Collaborative [FKM^{+}25, MZLS20]. CollectiveHLS [FKM^{+}25]. Column [VL11]. Column-Oriented [VL11]. Combination [DRHM22]. Combinational [SFNP23]. Combinatorial [WSC09]. Combined [PP10]. CoMeFa [ABB^{+}23]. Commercial [FL16, MWL^{+}15, ZML^{+}22, PANBI11]. Communication [CNK^{+}24, HNS^{+}10, KLD16, PCC^{+}25, SKJ22, TL11, US23, VG14, HZW^{+}13]. Communication-Aware [HNS^{+}10]. Communication-centric [VG14]. Communications [BNW^{+}10]. Compact [BHF^{+}25, HBXA23]. Comparison [BNW^{+}10, LA17]. Compatible [LT09]. Compensation [DNL19]. Compilation [BPFD11, ES22, MWK^{+}12, SFNP23, UAS16, XPN^{+}24]. Compile [PPR^{+}10]. Compile-Time [PPR^{+}10]. Compiler [HLC^{+}15, ZG16]. Complex [YB18]. Complexity [FRS^{+}15]. Compliant [BCW21, KGC^{+}24]. Component [SCC10]. Components [ATJZ16, DC16, RDC^{+}21]. Composable [LBL23]. Composing [BSW^{+}23, LLO^{+}14, ZLY^{+}24]. Comprehensive [JCG^{+}12, MZLS20, GP13]. Compressed [DCL^{+}22, GS23]. Compressing [SB25]. Compression [GRG08, PP10, PBBP18, SKJ22, USY17, IYY^{+}11]. Compression/Decompression [PP10]. Compressor [CAPA^{+}09, HDN^{+}24, PABI09, PANBI11]. Computation [HH24, VDG^{+}24, IF23, MYST25, SXZ^{+}25b, dMdLC23]. Computational [CZZ23, HNM^{+}22, RGCL16, SFT^{+}23, TBS^{+}25]. Computations [RDC^{+}21]. Compute [ABB^{+}23, BBO^{+}25, LLS^{+}24, MHS09]. Compute-in-Memory [ABB^{+}23, LLS^{+}24]. Compute-Intensive [BBO^{+}25]. Computer [FHL^{+}24, LYS^{+}08, NSS^{+}11]. Computers [SPM^{+}10, THK12]. Computing [AJYH18, Bec14, CH10, CKG^{+}10, EAGEG09, FAB22, HNS^{+}10, IAG23, JCG^{+}12, MH15, MWBL21, NG25, RGGW10, RDB^{+}18, SKJ22, USY17, UAS16, UCR^{+}19, WGGR16, dDELVP13, KSG11]. Conference [AC14, LAA^{+}17]. Configurable [KGC^{+}24, LLS^{+}23, MCL^{+}23, PABI09, PRV21, WS16]. Configuration [AEM^{+}24, CAG^{+}22, DVK15, HBA^{+}15, KD10, SCY^{+}23]. Configuration-based [CAG^{+}22]. Configurations [MHK^{+}08]. CoNFV [ZSP^{+}21]. Congestion [AHAM^{+}19, CTH16]. Congestion-estimation [AHAM^{+}19]. Conjugate [RC10]. Connected [ATJZ16]. Connection [ZVS20]. Connection-aware [ZVS20]. Connections [WML^{+}24]. Connectivity [NI24, WAS^{+}23]. Consideration [TL11]. Considering [SC08]. Consolidation [DCL^{+}22]. Constant [FAB22, HCOB13]. Constant-Coefficient [FAB22]. constrained [MHS^{+}19, SMN^{+}23]. Constraint [LN23, MWK^{+}12]. Constraint-Aware [LN23]. Constraints [BAMR10, INF^{+}14, LP15]. Construction [LLZ^{+}24, YFW^{+}17]. Context [AB20, BMR16, KA21, NW11, SKW^{+}21]. Context-Switch [BMR16, KA21]. Continuous [GGR^{+}18]. Control [CJWC23, NW11, ZG16]. Control-Intensive [ZG16]. Controller [GdLlG^{+}14]. Converter [DNL19]. Converters [DSW^{+}24]. Convolution [WTS19]. Convolutional [BYB18, DWN^{+}22, LDJ^{+}17, LFN^{+}18, LLYC24, MHS^{+}19, PBBP18, TWL^{+}23, VDdSN23]. Coordination [ASGY12, PMC^{+}14]. Coprocessor [GS10]. CORDIC [RKV23, ZCL15, ZCL16]. CORDIC-Based [RKV23, ZCL16]. Core [IZO^{+}10, WPSI18, WMG^{+}10, BYB23, QNF^{+}23, SGNB08]. Correlation [GSJC13, LML^{+}23]. Correspondence [SM25]. Correspondence-free [SM25]. COSMIC [GGR^{+}18]. Cost [DPHT19, TL11, VVZL25, PDH11, ZH12]. Countermeasure [MMMT09]. Counters [LT09]. Counting [FK08, PBPLA17]. Covert [GER19, GTS23, TES24]. Covert- [GER19, GTS23]. Covert-channels [TES24]. CPU [CCF^{+}18, LLZ^{+}24, LAF24, MCD^{+}18]. CPU-FPGA [LAF24]. CPUs [TJCP25, TOS17]. Creating [DE22]. Creation [SFT^{+}23]. Creative [MCL^{+}13]. Cross [BDX^{+}19, DSW^{+}24, YGH^{+}18, GTS23]. Cross-layer [YGH^{+}18]. Cross-Platform [BDX^{+}19, DSW^{+}24]. Cross-VM [GTS23]. Cryo [TZWZ15]. Cryo-Electron [TZWZ15]. Crypto [GPL24]. Crypto-Puzzle [GPL24]. Cryptographic [BDGH15, SKW^{+}21, SGM09]. Cryptography [GFBF12, HBXA23, ILP24, KBM09, LYC^{+}24, SG15]. CRYSTALS [CKRG^{+}24, MCL^{+}23, ZHL^{+}21]. CRYSTALS-Dilithium [CKRG^{+}24, MCL^{+}23, ZHL^{+}21]. CRYSTALS-Kyber [CKRG^{+}24]. CSAIL2019 [GPL24]. CTScan [TJCP25]. Curve [ADSH18, GPP08, ILP24, KBM09, SG15]. Curve25519 [SG15]. Custom [GRG08, LCS14, PBBP18, TOS17]. Customizable [MPZ^{+}20]. Customized [LFN^{+}18, MLPK22]. cuts [KVK^{+}11]. Cyber [JSLO24]. Cyber-Physical [JSLO24]. Cybersecurity [PCFM23]. Cycle [CSK17]. Cycle-Accurate [CSK17]. Cyclostationary [LLL^{+}23].

D [BPCC09, IAG23, JB15, LP15, RNTW22, SPS12, TZWZ15, WZY^{+}24]. D-LSTM [RNTW22]. D-stacked [IAG23]. DANSEN [TBS^{+}25]. Data [ENPR22a, ENPR22b, GKLLA23, GKM^{+}12, HLW^{+}21, IABV15, KA21, KW22, LYZ^{+}18, PVB13, RMSK16, SZKR22, SWT^{+}22, SKH^{+}22, USY17, WAT15, dMdLC23, CA11, ZZB^{+}20]. Data-driven [HLW^{+}21]. Data-Flow [GKM^{+}12]. Data-Level [PVB13, CA11]. Database [LMD^{+}25, TBS^{+}25]. Databases [VL11]. Datacenters [BE19, BMC^{+}22]. Dataflow [GCL^{+}23, HDMN24, JSG^{+}22, JMGI23, MS25, YCV^{+}21, ZG16, ZFCZ25]. Datapath [SBC15, WHQ^{+}08]. Datapath-Oriented [WHQ^{+}08]. DBSCAN [SB15]. DCT [CA11]. Debug [JCGW20, WHQ^{+}08]. Debugging [AB23, IPC14, JCGW20, KS20]. Decision [CSS^{+}23, OKA19, TTY25]. Decoders [DLCJ20, CA11]. Decomposition [ZCL16]. Decompression [KBT09, PP10]. Deconvolutional [LFN^{+}18]. Dedekind [VDG^{+}24]. Dedicated [NZI22]. Deep [AHSS^{+}21, ABB^{+}23, BPF^{+}18, BATM22, CPW18, DWN^{+}22, EBYB20, GKLLA23, LLS^{+}24, LDJ^{+}17, MHT^{+}21, NGM^{+}23, RKV23, RVN24, RHLK18, XJXQ24, ZDS^{+}22, ZLY^{+}24, AGM^{+}22]. Deep-Learning [BPF^{+}18]. Deeply [LFN^{+}18]. defect [RD11]. Defenses [ZQ19]. defined [LUX^{+}21]. Deflection [KG17]. Deflection-Routed [KG17]. Defragmentation [FKS^{+}12]. Delay [LOM10, MHK^{+}08, SC08, WYZ16]. Delayed [FL20]. Delays [GNM^{+}15, WSC09]. DeLiBA2 [KHK24]. Demands [RUC11]. Demystifying [LFS22, ZFCZ25]. Dense [RC10, RMSK16]. Density [LML^{+}23]. DEpendability [KGS^{+}12, WGGR17]. Dependable [Ste10]. Dependency [GWXW21, JCGW20]. Deploying [ABB^{+}23, MKSB22, TDH^{+}22]. Deployment [BDX^{+}19]. Depth [CCF^{+}18]. Depthwise [LLYC24]. Derivatives [MWBL21]. Descriptor [SCL24]. Design [AIH^{+}24, BKT14, BMR16, CHG22, CCZ^{+}24, DLCJ20, DL09, EWL15, FHL^{+}24, FYZ^{+}25, FKM^{+}25, FZ25, GWPK20, GHO17, GCL^{+}23, IPC14, JSC14, JB15, JYJC23, KMK^{+}10, LP22, LVC24, LYH^{+}23, MKP09, MLFS22, MHS^{+}19, MZLS20, NBS13, PRV21, PCFM23, SJT09, SBC15, SCY^{+}23, SKB^{+}22, Tak12, TKB^{+}24, UNBR14, VDdSN23, WBCW24, ZMH^{+}23, ZHL^{+}21, ZW25, HLL08, HH13, MCL^{+}23, MAK^{+}12]. design-space [HLL08]. Designing [AHL^{+}14, FK08, KGC^{+}24, RVN24, SUK24]. Designs [BPCC09, DD18, DHL^{+}18, DLZ^{+}24, DB15, GMZ^{+}23, MKP23, NEPV24, RLM^{+}17, WYZ16]. Desktop [LYS^{+}08]. Detailed [NZI22]. Detection [AVCP20, AIH^{+}24, ATJZ16, GKA^{+}24, LBL23, PD15, SMN^{+}23, YGH^{+}18, KSG11]. Detector [MFOM23]. Deterministic [MS23, KB23]. Development [VTN09, DW13]. Device [CXG^{+}12]. Devices [FHL^{+}24, FKS^{+}12, RGCL16, WMG^{+}10]. DF [APR^{+}22, TTY25]. DF-BETA [TTY25]. Diagrams [CSS^{+}23]. Dictionary [GRG08]. Difference [NJLW14, SLH^{+}10]. Differential [MMMT09]. Digital [BNW^{+}10, DNL19, DSW^{+}24, LP15, MCN12, SFNP23, SSC16, LLS^{+}24]. Digital-Signal [SSC16]. Digitizer [TKB^{+}24]. Dilithium [CKRG^{+}24, MCL^{+}23, ZHL^{+}21]. Direct [WZL24, ZBR12]. Directional [KG17]. Directive [DLZ^{+}24]. Disaggregated [ASPP22]. Discovery [MCL^{+}13]. Discrete [GdLlG^{+}14, GPP08, LYC^{+}24]. Distributed [KHK24, LZZ^{+}24, LP22, OKA19, HZW^{+}13]. Div [MLFS22]. Divider [MLFS22]. DL [GZY^{+}18]. DMA [SSK^{+}23]. DNN [ACFm^{+}25, APR^{+}22, RRLW22]. Do [BYB18]. Domain [DDH^{+}11, SFT^{+}23, NSS^{+}11]. Domain-Specific [DDH^{+}11, SFT^{+}23]. Domains [NZS^{+}23]. DONGLE [WZL24]. don’t [MBJJ11]. don’t-care-based [MBJJ11]. DORGA [NW11]. Dot [VDdSN23]. Double [LGD^{+}14]. Double-Precision [LGD^{+}14]. DPA [LOM10]. DRAM [SLL^{+}20]. DRAM-based [SLL^{+}20]. Drift [GKLLA23]. Drift-awareness [GKLLA23]. Driven [DSK15, LRA13, MWL^{+}15, Ste10, YGH^{+}18, EA11, HLW^{+}21, ZML^{+}22]. driver [LKJ^{+}11]. Drone [SMN^{+}23]. DSL [SP20]. DSL-Based [SP20]. DSP [CBFM14]. DSP48E2 [NEPV24]. DSPs [DGP10]. Dual [HF14, LYH^{+}23]. Dual-Rail [HF14]. Dual-ring [LYH^{+}23]. Dynamic [AZM^{+}19, BHI15, CZZ23, CTH16, CJWC23, CW09, DVH^{+}15, DRHM22, ERO25, FKS^{+}12, JS24, KP14, KS20, LP15, TWG^{+}20, VPPK20, VVZL25, VMV15, NSS^{+}11, JS24]. Dynamic-ACTS [JS24]. Dynamic-scheduling [CJWC23]. Dynamically [BBND10, DGP^{+}15, HHSC10, MSSM10, NNY12, QMG^{+}24, TL11, ZBB^{+}16, HLL08, HH13, IYY^{+}11]. Dynamics [CH10, HNM^{+}22, SFT^{+}23]. DyRecMul [VVZL25].

EAM [WCK21]. Early [TTY25]. ECC [DL09, GS10, WHJ^{+}23]. Eciton [CJH^{+}24]. Edge [BF23, CJH^{+}24, DCL^{+}22, KZB23, TJCP25]. Edge-skip-calculation [DCL^{+}22]. Edition [DH08]. editor [AN09, Che19]. Editor-in-Chief [Che19]. Editorial [CDM15, Che19, DH08, GSCB15, WBAM10]. Editors [SJT09]. Edwards [ADSH18]. Effect [HLC^{+}15]. Efficiency [BYB18, DPHT19, NZS^{+}23, PBBP18, YLF^{+}24, HDN^{+}24, LYH^{+}23]. Efficient [BHXJ24, BMR16, BGSL17, CCT^{+}22, DD15, DLBM18, EBYB20, FT17, FK08, HH24, HU10, ILP24, KSCC10, KVV24, KD19, LYZ^{+}18, LLYC24, MCD^{+}18, PCC^{+}25, PBPLA17, RLY^{+}15, RLM^{+}17, SLH^{+}10, SSK^{+}23, SFNP23, Tho15, VDdSN23, WML^{+}24, XJXQ24, dBVW^{+}24, CA11, SMN^{+}23]. EiC [Che22]. Elastic [APR^{+}22]. Elastic-DF [APR^{+}22]. Electrical [KGT19]. Electrical-level [KGT19]. Electromagnetic [SGM09]. Electron [TZWZ15]. Element [MVGB15]. Elementary [LGD^{+}14]. Eliminating [SKJ22]. Elimination [NCJ^{+}15]. Elliptic [GPP08, ILP24, KBM09, SG15]. Embedded [BHI15, Kap16, KBT09, RRW^{+}22, SML^{+}24, WPSI18, WHQ^{+}08]. Embedding [HLW^{+}21]. Empirical [RKV23]. Emulation [CSK17, TJCP25, ZZB^{+}20]. EmulatoR [KGS^{+}12]. Enable [AZM^{+}19, AB20, RDB^{+}18]. Enabled [HNM^{+}22, JS24, BYB23, TES24]. Enabling [MWL^{+}15, OVI^{+}12]. encoded [KVK^{+}11]. Encryption [HTG25, SMOP15]. End [BSW^{+}23, BPF^{+}18, BF23, CMT^{+}24]. End-to-End [BPF^{+}18, BSW^{+}23, BF23, CMT^{+}24]. Energy [DSK15, DLBM18, DPHT19, KLD16, LP15, NZS^{+}23, SMN^{+}23, WCK21, dBVW^{+}24, CA11]. Energy-Efficient [DLBM18, dBVW^{+}24, SMN^{+}23, CA11]. Energy-Reliability [DSK15]. Engine [JZL^{+}24, JYJC23, MS23, SXZ^{+}25b]. Engines [RVN24, XCG^{+}09, YXC^{+}11]. Enhance [RKV23]. Enhanced [JCCM09, ZCL15, ZCL16, LXS^{+}24, MS23]. Enhancement [ABCC09]. Enhancements [ZVS20]. Enhancing [GKM^{+}12, GHWS22, MCN12, RDC^{+}21, TYB18]. Enough [RNTW22]. Ensemble [LBL23]. Ensembles [OKA19]. Entropy [FK08]. Enumerative [SXZ^{+}25b]. Environment [MCL^{+}13]. Environmentally [SGW20]. Equations [GFL^{+}15]. Era [AGM^{+}22]. Error [DVK15, PD15, ZNA^{+}18]. Espresso [SCY^{+}23]. Estimation [DHL^{+}18, DNL19, LML^{+}23, WPSI18, WCK21, WSDH23, YCV^{+}21, AHAM^{+}19]. ETA [PEM^{+}09]. Evaluating [LAL13, PFL22, PB18, SY24, WGGR16]. Evaluation [BDX^{+}19, KY18, Lee23, LOM10, NJLW14, PCC^{+}25, SKB^{+}22, SMOP15, TKB^{+}24]. Evaluations [KGS^{+}12]. Event [GWS23]. Event-based [GWS23]. Evolution [CBC^{+}12]. Evolutionary [ZCK22]. Evolvable [DS15]. Evolving [MLPK22]. Exact [OROS^{+}19]. Example [SP20]. ExaNeSt [PCC^{+}25]. Execution [DSK15]. ExHiPR [XPN^{+}24]. Exotic [FT17]. Expanding [HCGG24]. Experiment [QRDC^{+}15]. Explicitly [NI24]. Exploitation [BBO^{+}25, INF^{+}14, MAK^{+}12]. Exploiting [BDGH15, CA11, EAGEG09, GER19, LCS14, LZF^{+}10, MCD^{+}18, PVB13, SKH^{+}22, TBS^{+}25]. Exploration [BPF^{+}18, CCZ^{+}24, EWL15, JYJC23, MZLS20, RRLW22, SCY^{+}23, UNBR14, HLL08, LKJ^{+}11]. Exploring [ACFm^{+}25, JTLC09, MWL^{+}15, NI24, SPS12]. Exponentiation [dDELVP13]. Expression [SXZ^{+}25b]. Expressions [LT09]. Extended [DGP10, XPN^{+}24]. Extending [GdLlG^{+}14]. Extensible [WS16]. Extension [GB11, GFBF12, MWK^{+}12]. Extraction [GNM^{+}15].

Fabric [BHB14, SDGL^{+}22, WHQ^{+}08, SPS12]. Fabrics [KA17]. Factor [LRA13]. Factored [KAL14]. FADO [DLZ^{+}24]. Fantastic [FZ25]. Fast [AVCP20, BAG15, BPF^{+}18, BDX^{+}19, CSK17, HU10, HLW^{+}21, ILP24, JCGW20, JM14, NW11, UNBR14, VVZL25, XPN^{+}24, YGH^{+}18, ZCK22, ZWM19, SSF^{+}13, SP20]. Fault [BKT14, JCG^{+}12, RLY^{+}15, WBCW24]. Fault-Tolerant [BKT14, RLY^{+}15]. FAXID [GKA^{+}24]. FCCM [Deh20, LH23]. FCCM’16 [Bak18]. FDRA [QMG^{+}24]. FeatherNet [MHS^{+}19]. Features [SM25]. Feel [AB20]. FEM [BGSL17]. FFT [LLL^{+}23]. FGPU [MHT^{+}21]. Fiber [LZZ^{+}24]. FiberFlex [LZZ^{+}24]. Fibonacci [SCY^{+}23]. Field [AC14, BCW21, CAPA^{+}09, DPHT19, HNM^{+}22, SCC10]. Field-Programmable [AC14, DPHT19, HNM^{+}22, SCC10]. FIFOs [GWPK20]. Filter [BPCC09]. Filtering [LP15]. Filters [CNE^{+}15, SWT^{+}22]. Financial [TB10]. Find [FZ25]. Fine [KD19, NZS^{+}23, RBR16, YHK^{+}21, ZNA^{+}18]. Fine-Grained [NZS^{+}23, RBR16, ZNA^{+}18, KD19, YHK^{+}21]. FinFET [SY24, TMLS21]. Fingerprint [XJD^{+}16]. Finite [NJLW14, SLH^{+}10, GDHG11]. Finite-Difference [NJLW14]. FINN [BPF^{+}18]. FINN-R [BPF^{+}18]. FIR [LP15]. First [LAA^{+}17, Che22]. Fixed [LLL^{+}23, RGCL16, SFNP23, WL10, WMG^{+}10]. Fixed- [WL10]. Fixed-point [LLL^{+}23]. FlexCNN [BSW^{+}23]. flexibility [LW08]. Flexible [DS15, LBRS16, LZ19, MCD^{+}18, TLC^{+}24, US23, dBVW^{+}24]. FlexSaaS [CNZ^{+}18]. Flight [QRDC^{+}15]. Floating [FL16, HU10, OBD13, RC10, USY17, WL10, WS16, WWC^{+}22, dDELVP13]. Floating-Point [FL16, HU10, OBD13, USY17, WL10, WS16, dDELVP13, WWC^{+}22]. Floorplan [DLZ^{+}24, KSCC10]. Floorplan-Aware [DLZ^{+}24]. Floorplanning [MSSM10]. Flow [BNW^{+}10, BMR16, BHB14, CJWC23, GKM^{+}12, HNM^{+}22, KA17, RLY^{+}15, SCC10, ZG16, ZMH^{+}23]. Fluid [SFT^{+}23]. FNNG [LLZ^{+}24]. Footprint [CW09]. Forest [TTY25]. Forget [AI22]. format [SLL^{+}20]. Formulation [FZ25]. FOS [VPPK20]. Fourier [SP20]. FPGA [BYB23, AZM^{+}19, AKM^{+}25, AHSS^{+}21, AVCP20, APR^{+}22, ABCC09, AIH^{+}24, AGM^{+}22, AB20, BCE^{+}10, BAG15, BHXJ24, BHF^{+}25, BSW^{+}23, BPFD11, BFBN^{+}20, BDGH15, BE19, BMC^{+}22, BYB18, BBO^{+}25, BF23, CHG22, CA11, CZD^{+}25, Che11, CW09, CCF^{+}18, CSS^{+}23, CSK17, CZ09, CLL^{+}22, CAG^{+}22, DFB^{+}22, DW13, DVK15, DHL^{+}18, DNL19, DE22, DL09, ES22, EBYB20, EAAAA19, FB24, FRS^{+}15, FLM^{+}17, GKA^{+}24, GP13, GWPK20, GFBF12, GMBC17, GSJC13, GER19, GRG08, GHWS22, GHO17, GZY^{+}18, GMZ^{+}23, HBXA23, VDG^{+}24, HF14, HGLS11, HCOB13, HLW^{+}21, HTG25, Ien23, IF23, IPC14, JCG^{+}12, JRHK15, JCCM09, JM14, KLD16, KLC11, KZB23, KVV24, KM10, Kap16, KBM09, KVK^{+}11, KMK^{+}10, KY18, KAL14, KA17, KGS15, KBT09, KD10, KS20, LA17, LUX^{+}21, LCS14, LW08, LZF^{+}10, LGD^{+}14, Leo22, LAL13, LLL^{+}23, LML^{+}23, LLS^{+}24, LZZ^{+}24, LXS^{+}24]. FPGA [LDJ^{+}17, LFN^{+}18, LLYC24, LLZ^{+}24, LCL^{+}24, LT09, LBL23, LAF24, LKJ^{+}11, MLPK22, MLFS22, MCD^{+}18, MKP23, MAK^{+}12, MCN12, MZLS20, MPZ^{+}20, MHS09, NNY12, NGM^{+}23, NZI22, NI24, PWP^{+}16, PDH11, PMGL22, PABI09, PMKM11, PB18, PCC^{+}25, PBPLA17, PCFM23, RDC^{+}21, RVN24, RC10, RRLW22, SUK24, SLH^{+}10, SB15, SC08, SV09, Sha22, SLL^{+}20, SZZW23, SML^{+}24, SCL24, SM25, SMN^{+}23, SXZ^{+}25a, SXZ^{+}25b, TL11, TLC^{+}24, TWL^{+}23, Tho15, TB10, TTY25, TKB^{+}24, TES24, USY17, US23, UNBR14, VPPK20, VDdSN23, WTS19, WYZ16, WBCW24, WHQ^{+}08, WZL24, WWC^{+}22, WZY^{+}24, WGGR16, WGGR17, XPN^{+}24, XCG^{+}09, XRC23, YXC^{+}11, YB18, YOY17, YGH^{+}18, ZDS^{+}22, ZBR12, ZZJB13, ZQ19, ZCK22, ZBC^{+}09, ZNA^{+}18, ZMH^{+}23, ZVS20, ZQBL25, ZBB^{+}16]. FPGA-Accelerated [GKA^{+}24, BE19, SM25]. FPGA-Array [SLH^{+}10]. FPGA-Aware [LCS14]. FPGA-Based [TLC^{+}24, UNBR14, ZZJB13, AIH^{+}24, CZ09, GHO17, JCG^{+}12, Kap16, KBT09, LCL^{+}24, LT09, NNY12, RC10, SB15, SXZ^{+}25a, USY17, WGGR16, YOY17, ZNA^{+}18, ZBB^{+}16, CZD^{+}25, GZY^{+}18, KZB23, LZZ^{+}24, LLYC24, LBL23, MLFS22, NGM^{+}23, SUK24, SML^{+}24, TWL^{+}23, TTY25, TKB^{+}24, WTS19, WWC^{+}22, YXC^{+}11, ZBR12, ZQ19, ZQBL25]. FPGA-enabled [TES24]. FPGA-optimized [ZCK22]. FPGA-Orchestrated [WZL24]. FPGA-SoCs [GHWS22]. FPGADefender [LMG^{+}20]. FPGAs [AB14, AKA09, AHAM^{+}19, AJYH18, ABB^{+}23, AI22, AB23, BKT14, BAMR10, BNW^{+}10, BPCC09, BHB14, CMT^{+}24, CAPA^{+}09, CBFM14, CPW18, CCT^{+}22, CXG^{+}12, CPFM21, CPN^{+}09, CFBS15, DRF24, DH08, DCS24, DLCJ20, DDH^{+}11, DD15, DGP^{+}15, DGP10, DLZ^{+}24, DB15, ENPR22a, ENPR22b, FL16, FAB22, GRNW22, GTS23, GCL^{+}23, HZS^{+}25, HU10, HBA^{+}15, HDN^{+}24, IAG23, JS24, KG17, KA21, KW22, KTB^{+}24, KGT19, KD21, LMG^{+}20, LLO^{+}14, LMD^{+}25, LN23, LSP^{+}23, LLS^{+}23, LOM10, LFS22, LGW^{+}14, MA24, MHK^{+}08, MKSB22, MMMT09, MPK22, MVGB15, MDL^{+}23, MSSM10, MHS^{+}19, OKA19, PFL22, PANBI11, PVA^{+}09, PVB13, RVHP16, RLM^{+}17, RDB^{+}18, RHLK18, SY24, SGM09, SWT^{+}22, SKH^{+}22, SSF^{+}13, SPS12, SB08, Ste10, SDM^{+}18, SSC16, SKB^{+}22, SMOP15, TMLS21, TWG^{+}20, TYB18, TYL^{+}23, VVZL25, VMV15, WSC09, WAT15, ZZB^{+}20, ZML^{+}22]. FPL [BGSL17, YFW^{+}17, CDM15, CS17, GKK24, KS24, LAA^{+}17, MST22]. FPT [KZ23, Lee23]. FPT’12 [AC14]. FPT’20 [SLD23]. FPT’21 [Che22]. FPU [KGC^{+}24]. Framework [ASGY12, AHSS^{+}21, BSW^{+}23, BPF^{+}18, CCT^{+}22, CNK^{+}24, CCZ^{+}24, CKG^{+}10, DFB^{+}22, FHL^{+}24, GCL^{+}23, JCG^{+}12, JRHK15, KHK24, KD21, LZ19, QMG^{+}24, RGGW10, SGW20, SGC21, SLH^{+}25, TDH^{+}22, TYL^{+}23, UAS16, VTN09, WPSI18, WGGR16, ZDS^{+}22, HLL08, SSF^{+}13, SPS12]. Free [AB20, GWPK20, SM25]. Frequency [WSDH23]. Frequent [PBPLA17, ZZJB13]. FRoC [AZM^{+}19]. Fruit [YSC^{+}23]. Fruit-80 [YSC^{+}23]. fSEAD [LBL23]. FSM [GDHG11]. FT [WTS19]. Full [CPN^{+}09, DFB^{+}22]. Full-stack [DFB^{+}22]. Full-System [CPN^{+}09]. Fully [KAL14]. Function [LGD^{+}14, LML^{+}23, MWBL21, SFNP23, ZSP^{+}21]. Functional [RUC11]. Functionality [LMD^{+}25]. Functions [NCJ^{+}15, SAD10, SB25]. Future [BMC^{+}22, LUX^{+}21, LCL^{+}24].

Galois [LYC^{+}24, SCY^{+}23]. Game [MCL^{+}13]. Gap [MWL^{+}15, TOS17]. Gaps [BYB18]. Gate [BCW21, DPHT19, HNM^{+}22, SCC10]. Gaussian [SBC10, TL08, Tho15]. GCN [HLW^{+}21, TWL^{+}23]. General [AJYH18, GFBF12, ZDS^{+}22]. General-Purpose [AJYH18, ZDS^{+}22]. Generalized [ZWM19]. Generated [HLC^{+}15, LP15, GRNW22]. Generating [BMR16, GNM^{+}15, HZS^{+}25]. Generation [BS15, LSP^{+}23, LGW^{+}14, MWK^{+}12, PRV21, SP20, SCC10, TL08, GL08]. Generator [GHO17, LYC^{+}24, SBC10, SSC16, Tho15, XJXQ24]. Generators [RVHP16]. Generic [KA21, KD21, SZKR22]. Genetics [CA23]. Genome [XRC23]. Genomes [AVCP20]. GIB [SZZW23]. Global [GFL^{+}15, JSC14, KVV24]. GPP [TB10]. GPU [RNTW22, TB10]. GPUs [AJYH18, BNW^{+}10, CFBS15]. Gradient [RC10]. Grain [IZO^{+}10, YSC^{+}23]. Grained [NZS^{+}23, RBR16, VL11, XJD^{+}16, ZNA^{+}18, KD19, WCK21, YHK^{+}21]. Graph [CM14, CCT^{+}22, FRS^{+}15, GWXW21, JS24, LLZ^{+}24, MVGB15, MS25, TLC^{+}24, TWL^{+}23, ZG16, ZFCZ25, TLC^{+}24]. Graph-Based [MVGB15, GWXW21]. Graph-OPU [TLC^{+}24]. Graphic [BCW21]. graphics [BG08]. Graphs [DRF24, SLL^{+}20]. GraphScale [DRF24]. Grid [SLL^{+}20]. Grid-format [SLL^{+}20]. GRNG [Tho15]. GROK [GNM^{+}15]. GROK-LAB [GNM^{+}15]. GS [TKB^{+}24]. GS/s [TKB^{+}24]. Guest [AN09, CDM15, DH08, GSCB15, WBAM10, SJT09].

Hadamard [Tho15]. Handling [SKW^{+}21]. Hard [AB14, ZCK22]. Hardened [LFS22]. Hardware [ACFm^{+}25, ADSH18, AV13, AIH^{+}24, BHF^{+}25, BCW21, BPFD11, BS15, CBC^{+}12, CKRG^{+}24, CZZ23, CBR^{+}14, CZ09, DD18, DFB^{+}22, DLCJ20, DS15, FHL^{+}24, FZ25, GKLLA23, GPP08, HCGG24, HBXA23, HNM^{+}22, HHSC10, HLC^{+}15, HLN^{+}10, ILP24, IBH^{+}15, JSLO24, KVV24, KHK24, KBT09, LLYC24, LVC24, MOG^{+}13, MCC10, NG25, PD15, PSM^{+}14, PCC^{+}25, RNTW22, SBC10, SKW^{+}21, SP20, TL08, TOS17, WL10, YSC^{+}23, YLF^{+}24, YBS16, ZG16, ZHL^{+}21, BG08, HH13, SC11, SMN^{+}23]. Hardware-Accelerated [MCC10, GKLLA23]. Hardware-Based [HLN^{+}10]. Hardware/Software [HHSC10, KHK24, HH13, SC11]. Hash [IABV15]. Hashes [LVC24]. HBM [DRF24, JS24, LXS^{+}24, LLS^{+}23, SKH^{+}22, TYL^{+}23]. HBM-based [LLS^{+}23, TYL^{+}23]. HBM-Enabled [JS24]. HBM-enhanced [LXS^{+}24]. HBM2 [BYB23]. HBM2-enabled [BYB23]. HDL [RDC^{+}21]. Healing [BHI15]. healthier [ZH12]. Heap [BAG15]. Heap-Based [BAG15]. Hessian [CMT^{+}24]. Hessian-aware [CMT^{+}24]. heterogeneity [LKJ^{+}11]. Heterogeneous [ASGY12, AHL^{+}14, BPCC09, CNE^{+}15, CCF^{+}18, GFL^{+}15, KSCC10, KP14, LLZ^{+}24, OVI^{+}12, RVN24, TZWZ15, TDH^{+}22, UAS16, WSDH23, YB18, ZSP^{+}21, ZLY^{+}24, PMKM11, SPS12]. Hiding [MMMT09, THK12]. Hierarchical [CCZ^{+}24]. Hierarchies [YFW^{+}17]. HierCGRA [CCZ^{+}24]. High [AIH^{+}24, BHF^{+}25, BGSL17, BS15, CKRG^{+}24, CZZ23, CJWC23, CH10, CSS^{+}23, CKG^{+}10, DHL^{+}18, DLZ^{+}24, EAGEG09, FYZ^{+}25, FKM^{+}25, GFC^{+}25, GWXW21, HZS^{+}25, HNM^{+}22, HDMN24, HDN^{+}24, HNS^{+}10, HLC^{+}15, IPC14, JZL^{+}24, JSG^{+}22, JYJC23, LN23, LXS^{+}24, LLS^{+}23, LYH^{+}23, MH15, MCL^{+}23, MPZ^{+}20, NBS13, OROS^{+}19, PMGL22, PBBP18, RC10, SPM^{+}10, SGM09, SSK^{+}23, SFNP23, SLH^{+}25, SFT^{+}23, SSC16, TB10, USY17, WBC16, WBR18, WWC^{+}22, XPN^{+}24, ZZB^{+}20, ZBC^{+}09, MAK^{+}12, PANBI11]. High-Accuracy [DHL^{+}18]. High-Bandwidth [HNM^{+}22, SFT^{+}23, ZZB^{+}20]. High-Efficiency [PBBP18, HDN^{+}24, LYH^{+}23]. High-Level [CKG^{+}10, DLZ^{+}24, FKM^{+}25, GFC^{+}25, HLC^{+}15, IPC14, JYJC23, LN23, NBS13, OROS^{+}19, SLH^{+}25, WBC16, XPN^{+}24, CJWC23, CSS^{+}23, GWXW21, HDMN24, SFNP23]. High-Order [BGSL17]. High-Performance [BHF^{+}25, CH10, EAGEG09, HZS^{+}25, HNM^{+}22, HNS^{+}10, JZL^{+}24, JSG^{+}22, LLS^{+}23, MH15, PMGL22, SPM^{+}10, SSC16, TB10, USY17, WBR18, LXS^{+}24, MCL^{+}23, MPZ^{+}20, WWC^{+}22, PANBI11]. High-Speed [BS15, ZBC^{+}09]. High-Throughput [FYZ^{+}25, SSK^{+}23, MAK^{+}12]. High-Volume [CKRG^{+}24]. Highly [DLBM18, ES22, IGM^{+}20, RDC^{+}21, TLC^{+}24]. Hipernetch [PMGL22]. HLPerf [ZFCZ25]. HLS [BBO^{+}25, CCT^{+}22, CLL^{+}22, DRHM22, GRNW22, GCL^{+}23, JCGW20, KTB^{+}24, WZL24, ZFCZ25]. HLS-based [ZFCZ25]. HLS-generated [GRNW22]. HLS-specific [BBO^{+}25]. HMAC [MAK^{+}12]. Holistic [FZ25]. Homogeneous [LAL13]. Hoplite [KG17]. HopliteBuf [GWPK20]. HopliteML [MLPK22]. HPC [MPK22, MKP23, PCC^{+}25, US23]. HTTP [MA24]. HW [MCL^{+}23]. HyBNN [YLF^{+}24]. Hybrid [DS15, FAB22, GHWS22, MFOM23, RGCL16, SGW20, TYL^{+}23]. Hybrid-TPM [GHWS22]. HyperTransport [SGNB08].

I/O [MHS09, RGCL16]. ICAP [AEM^{+}24]. ICFPT [AN09]. iDEA [CBFM14]. Identification [DVH^{+}15, GHO17]. Idle [NCJ^{+}15]. IEEE [KGC^{+}24]. IEEE-Compliant [KGC^{+}24]. II [ENPR22b, SMOP15]. III [SMOP15]. ILP [FZ25]. Image [BAG15, CZ09, SDM^{+}18]. Images [TZWZ15]. Impact [HBA^{+}15, KLD16, KW22, SY24]. Implementation [AV13, BAG15, BHXJ24, BHF^{+}25, BCW21, DNL19, GRG08, GMZ^{+}23, HBXA23, HF14, KVV24, LGD^{+}14, LLL^{+}23, LML^{+}23, LVC24, MKP09, OBD13, RC10, SSK^{+}23, SV09, SAD10, WZY^{+}24, CA11, SSF^{+}13]. Implementations [BDGH15, FLM^{+}17, ILP24, MDL^{+}23]. Implemented [PVB13]. Implementing [BKT14, BNW^{+}10, SG15]. Imprecise [SBC15]. Improve [BYB18, LZF^{+}10, SDG12]. Improved [GHO17, JCCM09]. Improving [DRHM22, LZ19, NZS^{+}23, WBCW24, YKBS10]. In-Circuit [KS20]. In-Depth [CCF^{+}18]. In-memory [LMD^{+}25]. In-Network [NG25]. In-the-Cloud [BDX^{+}19]. Increasing [SRR23]. Incremental [ERO25, GGR^{+}18, XPN^{+}24, GL08]. Independent [PMC^{+}14]. Index [BAG15]. Index-Aware [BAG15]. Inducing [CAG^{+}22]. Industrial [SML^{+}24]. Inference [APR^{+}22, BYB18, CJH^{+}24, CZD^{+}25, DWN^{+}22, EBYB20, GZY^{+}18, MCD^{+}18, NGM^{+}23, OKA19, RHLK18, SFNP23, WML^{+}24]. Infinite [SWT^{+}22]. Information [GSJC13]. Infrastructure [HBA^{+}15, MKSB22, ZZB^{+}20, HH13]. Input [CAPA^{+}09, FK08]. Insensitive [GMZ^{+}23]. Insertion [LOM10]. Instance [RLM^{+}17]. Instance-Specific [RLM^{+}17]. Instruction [GB11, GWXW21, WBR18, YGH^{+}18]. Instruction-Set [GB11]. Instructions [LCS14]. Integer [MLFS22]. Integration [GS10, JRHK15, LRA13, YBS16]. Intellectual [HTG25]. Intelligent [LZZ^{+}24, MS25]. Intensive [BBO^{+}25, ZG16]. Inter [MKP23]. Inter-FPGA [MKP23]. Interactions [KD19]. Interconnect [FK08, NZI22, RBR16, TMLS21, SPS12]. Interconnects [KS20]. Interface [GKGS23, JB15, RUC11]. Internal [AEM^{+}24, HBA^{+}15]. International [AC14, DH08, VG14]. Interrupt [AB20]. Intersection [WHJ^{+}23]. Intra [GNM^{+}15, HF14]. Intra-cluster [GNM^{+}15]. Intra-Masking [HF14]. Intrinsic [MHK^{+}08]. Introduction [AC14, Bak18, Bec14, BE19, BL08, CS17, Che16, CPW18, Che11, CWBD09, DC16, Deh20, ENPR22a, ENPR22b, FB24, GC13, GKK24, Hüb12, Ien23, KZ23, KS24, Leo22, LH23, MST22, SJT09, Sha22, SML^{+}24, SLD23, VG14, AN09]. Introspection [GGR^{+}18]. Invariant [PD15]. IoT [SML^{+}24]. IP [IZO^{+}10]. IPs [EAAAA19]. IR [ZG16]. Irregular [WSDH23]. Isolated [MMMT09]. Issue [AC14, CWBD09, DC16, ENPR22a, ENPR22b, Hüb12, KZ23, KS24, SML^{+}24, VG14]. Itemset [ZZJB13]. Itemsets [PBPLA17]. Iterative [DCS24, LZ19, BC11].

JIT [BPFD11]. JITPR [SSF^{+}13]. Jitter [LSP^{+}23]. Jitter-based [LSP^{+}23]. Join [YOY17]. Journal [Che22, Lee23]. Journal-first [Che22]. Junction [TYB18]. Junction-Based [TYB18].

KAPow [DHL^{+}18]. Karatsuba [BHXJ24]. Kernel [FLM^{+}17, FL20, PWP^{+}16]. Kernels [JB15]. Key [ADSH18, GFBF12]. KLT [DB15]. KNN [LLZ^{+}24]. KNNv2 [LLS^{+}23]. Knowledge [GNM^{+}15]. Kyber [CKRG^{+}24].

L [LLZ^{+}24]. L-FNNG [LLZ^{+}24]. Lab [MCN12, GNM^{+}15]. LambdaRank [YXC^{+}11]. Language [CZD^{+}25, CKG^{+}10, SDM^{+}18]. Languages [SFT^{+}23]. Large [CCZ^{+}24, CZD^{+}25, CSK17, DRF24, KM10, KD21, LLZ^{+}24, MWL^{+}15, SWT^{+}22, SXZ^{+}25a, YCV^{+}21]. Large-Scale [CSK17, LLZ^{+}24, SXZ^{+}25a, YCV^{+}21, CCZ^{+}24, KD21]. Latencies [BAMR13]. Latency [GMZ^{+}23, THK12, ZW25]. Latency-Insensitive [GMZ^{+}23]. Latest [HDN^{+}24]. Lattice [LYC^{+}24]. Lattice-Based [LYC^{+}24]. Layer [IBH^{+}15, YGH^{+}18]. Layers [SY24]. Layout [KY18, SY24]. LDPC [CA11]. Leading [HH24]. Leakier [GER19]. Learnable [SB25]. Learned [WAS^{+}23]. Learning [AHAM^{+}19, AHSS^{+}21, ABB^{+}23, BPF^{+}18, BATM22, CPW18, EBYB20, ERO25, GKLLA23, LLS^{+}24, MHT^{+}21, MYST25, MCN12, NGM^{+}23, RRW^{+}22, RVN24, RHLK18, TDH^{+}22, XJXQ24, ZLY^{+}24, AGM^{+}22]. Least [FLM^{+}17, FL20, PWP^{+}16]. Length [LRA13, EA11]. Level [ASGY12, BBO^{+}25, CKG^{+}10, DD18, DS15, DLZ^{+}24, DL09, FKM^{+}25, GSJC13, GFC^{+}25, HLC^{+}15, IPC14, JYJC23, LN23, NBS13, NZI22, OROS^{+}19, PVB13, QMG^{+}24, SLH^{+}25, TTY25, WBC16, XPN^{+}24, ZW25, CA11, CJWC23, CSS^{+}23, GWXW21, HDMN24, KGT19, SFNP23]. Leveraging [ERO25, SC11]. libraries [KVK^{+}11]. Library [FL16, HZS^{+}25, LBL23, WL10, HCGG24]. Licensing [EAAAA19]. Lightening [DLBM18]. LightNNs [DLBM18]. Lightweight [BHXJ24, IF23, RD11, TWL^{+}23]. like [AB23]. Line [IABV15]. Line-Rate [IABV15]. Linear [FC24, HZS^{+}25, SB25, WGGR17, ZBR12]. Linux [MSF16, MKSB22]. Linux-based [MKSB22]. Listen [DSW^{+}24]. Listing [NI24]. Lists [WAT15]. Lithographic [CZ09]. LLVM [BBO^{+}25]. Load [DLBM18, THK12, RD11]. load-time [RD11]. Local [JSC14, MYST25]. Locality [TTY25]. Logarithm [GPP08]. Logarithmic [BPFD11]. Logarithmic-Time [BPFD11]. Logic [AZM^{+}19, DGP10, EBYB20, IZO^{+}10, MHS09, PABI09, SFNP23, TMLS21, WAS^{+}23, WBC16, KD19, MBJJ11, PMKM11]. Long [GER19, UHU09]. Loop [DCS24, DRHM22, DSB09, DPHT19, FZ23, SLH^{+}25]. Loops [PMC^{+}14, PFC15]. Low [CJH^{+}24, DHL^{+}18, DS15, DPHT19, FRS^{+}15, HBA^{+}15, KBM09, KCC^{+}14, KGS15, NZS^{+}23, VVZL25, VDdSN23, WWC^{+}22, ZBC^{+}09, ZH12]. Low-Complexity [FRS^{+}15]. Low-Cost [DPHT19, VVZL25, ZH12]. Low-Level [DS15]. Low-Overhead [DHL^{+}18, KGS15, NZS^{+}23]. Low-Power [KBM09, KCC^{+}14, ZBC^{+}09, CJH^{+}24]. Low-precision [WWC^{+}22]. Low-Speed [HBA^{+}15]. LSTM [IF23, RNTW22]. LUT [FK08, HF14, JCCM09, LZ19, NZI22, WAS^{+}23]. LUT-Based [LZ19, WAS^{+}23]. LUT-Level [NZI22]. LW [TWL^{+}23]. LW-GCN [TWL^{+}23]. LWE [BHXJ24, BHF^{+}25, HBXA23].

Machine [AHAM^{+}19, ERO25, KAL14, RRW^{+}22, TDH^{+}22, GDHG11]. Magnetic [TYB18]. Major [FL16]. Malicious [AEM^{+}24, LMG^{+}20]. Management [BBND10, GRNW22, KP14, LP15, HZW^{+}13]. Manager [BF23, OWMZ11]. Many [BYB23]. Many-core [BYB23]. Mapping [BBND10, CTH16, CNE^{+}15, CM14, CBR^{+}14, JSC14, JCCM09, LZ19, MCL^{+}13, MS25, SFNP23, KVK^{+}11]. Mapping-Scheduling [CBR^{+}14]. MARTE [GdLlG^{+}14]. Masking [HF14]. Massive [RNTW22]. Massively [JB15]. Matching [KVV24, SCL24, SXZ^{+}25b, TKB^{+}24, XJD^{+}16]. Matchings [BFBN^{+}20]. Mathematical [SUK24]. Matrices [RC10]. Matrix [DDB^{+}10, LCL^{+}24, UCR^{+}19, BC11]. Matrix-Vector [DDB^{+}10, BC11]. Maximizing [DSW^{+}24]. Maximum [BFBN^{+}20, GWXW21, RLY^{+}15]. Mean [FLM^{+}17, FL20]. Means [TK16]. Measure [BYB18]. Measurement [WSC09]. Mechanism [SLH^{+}10, SMOP15]. Mechanisms [SLL^{+}20]. Median [SWT^{+}22]. MEG [ZZB^{+}20]. Memories [AL16, BDGH15, DD15, IAG23, LLO^{+}14]. Memory [ABB^{+}23, AI22, BAMR10, BAMR13, CW09, HF14, HNM^{+}22, JB15, KLD16, LYZ^{+}18, LFS22, MSF16, PFC15, RGCL16, SDGL^{+}22, SFT^{+}23, THK12, TTY25, WBR16, YFW^{+}17, ZZB^{+}20, ASPP22, BC11, IAG23, LJS11, LMD^{+}25, LLS^{+}24]. Memory-Enabled [HNM^{+}22]. Memristor [dMdLC23]. Mercury [JLB^{+}08]. Merged [GRG08]. Merging [JCCM09, KD10, LCS14, TWG^{+}20]. Mesh [WZY^{+}24]. Meshes [BGSL17]. Message [Che19]. Metal [SY24]. Method [LLL^{+}23, NJLW14]. Methodology [DGP^{+}15, LBRS16]. Methods [AHAM^{+}19, KSCC10, SLH^{+}10, BC11]. Metrics [RGCL16]. MHz [WBR16]. Microarchitectural [LA17]. Microarchitecture [WBR16]. Microarchitectures [CCF^{+}18]. Microbenchmarking [LFS22]. MicroBlaze [MSF16]. Microcoded [GRG08, PWP^{+}16]. Microkernel [IBH^{+}15]. Microscopy [TZWZ15]. Middleware [KGS15]. Minifloat [ZQBL25]. Minimization [ZW25]. Minimum [KY18]. Minimum-Width [KY18]. Mining [PBPLA17, ZZJB13]. Minor [CM14]. Miss [AI22]. Miss-Optimized [AI22]. Mitigating [KGT19]. Mixed [AB14, WSZ^{+}25, YHK^{+}21]. Mixed-Precision [WSZ^{+}25]. mobile [NSS^{+}11]. MOdel [KGS^{+}12, AHL^{+}14, CZD^{+}25, CPFM21, DLCJ20, FZ25, FL20, LAL13, MOG^{+}13, MVGB15, SPM^{+}10, WZY^{+}24, YBS16, AGY^{+}11, DW13, HGLS11, PDH11]. MOdel-Based [KGS^{+}12, FZ25, DLCJ20]. Modeling [CCZ^{+}24, ERO25, GdLlG^{+}14, MZLS20, PVA^{+}09, SC08, SLH^{+}25]. Modelling [HCGG24, MPZ^{+}20, YB18]. Models [DLZ^{+}24, FHL^{+}24, JTLC09, KY18, RVN24]. Modern [AHAM^{+}19, CCF^{+}18, GCL^{+}23, KTB^{+}24, LFS22, LAF24, ZMH^{+}23]. MODES [KGS^{+}12]. Modifications [SDG12]. Modular [AL16, GKGS23, IPC14, NBS13, OWMZ11, VPPK20]. Module [DHL^{+}18, KD10, SGM09, TWG^{+}20, ZNA^{+}18]. Module-Based [KD10, ZNA^{+}18]. Modules [ZMH^{+}23]. Modulo [BAMR10, OROS^{+}19]. Molecular [CH10]. Monitoring [BCE^{+}10, GGR^{+}18]. Monte [TB10]. Montgomery [NEPV24]. Motion [MFOM23]. MPI [MKP23, SPM^{+}10]. MPSoC [BHI15, PSM^{+}14]. MPSoCs [DSK15, KP14, OVI^{+}12]. MRAM [ZBC^{+}09]. Multi [BYB23, AKM^{+}25, CAPA^{+}09, DLZ^{+}24, ES22, GWS23, GMBC17, GFC^{+}25, HGLS11, JSC14, KTB^{+}24, KGT19, LLO^{+}14, LN23, LYZ^{+}18, LYH^{+}23, MKSB22, PCC^{+}25, QMG^{+}24, QNF^{+}23, RDC^{+}21, WMG^{+}10, ZW25, MKP23]. Multi-Agent [GMBC17]. Multi-Application [JSC14]. Multi-bit [LYH^{+}23]. Multi-Chip [ZW25]. Multi-Core [WMG^{+}10, QNF^{+}23]. Multi-Die [BYB23, DLZ^{+}24, KTB^{+}24, ZW25]. Multi-FPGA [ES22, HGLS11, PCC^{+}25, RDC^{+}21, MKP23]. Multi-Input [CAPA^{+}09]. Multi-Level [QMG^{+}24]. Multi-Object [GWS23]. Multi-Pattern [LYZ^{+}18]. Multi-Ported [LLO^{+}14]. Multi-Technique [LN23]. Multi-Tenant [AKM^{+}25, KGT19, MKSB22]. Multi-Threaded [GFC^{+}25, QNF^{+}23]. Multicontext [VL11]. Multicore [MSF16, SDG12].

Multicores [AHL^{+}14, CFBS15]. Multidimensional [SB15]. MultiFactor [KLC11]. Multigigabit [PP10]. multilevel [AGY^{+}11, HGLS11]. multilevel-PGAS [AGY^{+}11]. Multimedia [DSK15]. Multiple [DCS24, MHK^{+}08, RVN24, SY24, SKW^{+}21]. Multiple-Metal [SY24]. Multiplication [DDB^{+}10, FAB22, LCL^{+}24, NEPV24, UCR^{+}19, BC11]. Multiplier [HCOB13, VVZL25]. Multipliers [HU10]. Multiported [AL16]. Multiprocessor [CPN^{+}09]. Multiprocessors [PPR^{+}10]. Multivariate [SBC10, TL08]. mutate [GL08]. MXP [LA17].

NAPOLY [KB23]. NASCENT2 [SZKR22]. Native [TBS^{+}25, DFB^{+}22]. NC [HCGG24]. NC-Library [HCGG24]. NCBI [MH15]. NDP [TBS^{+}25]. Near [ASPP22, DD15, IAG23, SZKR22, SDGL^{+}22, ZZB^{+}20]. Near-Associative [DD15]. Near-data [ZZB^{+}20]. Near-Memory [SDGL^{+}22, ASPP22, IAG23]. Near-Storage [SZKR22]. Nearest [LLS^{+}23]. Neighbors [LLS^{+}23]. Nested [HCGG24]. Net [EA11, LRA13]. Net-Length [LRA13]. Net-length-based [EA11]. Netlist [HTG25]. Network [BYB18, CTH16, CJH^{+}24, GWPK20, GMBC17, GZY^{+}18, JSC14, KAL14, LLYC24, MHS^{+}19, NG25, PMGL22, RNTW22, SFNP23, TWL^{+}23, US23, ZSP^{+}21, ZQBL25]. Network-on-Chip [CTH16, JSC14]. Networks [AB14, BPF^{+}18, CSK17, DWN^{+}22, KD10, LDJ^{+}17, LL12, MKP23, MVGB15, MYST25, PFL22, PVA^{+}09, PBBP18, QNF^{+}23, RKV23, SP20, SB25, TLC^{+}24, TKH^{+}19, VDdSN23, WAS^{+}23, YLF^{+}24, ZDS^{+}22, ZFCZ25, CMT^{+}24, HZW^{+}13, LW08]. Networks-on-Chip [AB14, CSK17]. NEURAghe [MCD^{+}18]. Neural [BPF^{+}18, BYB18, CJH^{+}24, DWN^{+}22, GZY^{+}18, KAL14, LDJ^{+}17, LLYC24, MHS^{+}19, PFL22, PBBP18, QNF^{+}23, RKV23, RNTW22, SFNP23, SB25, TLC^{+}24, TKH^{+}19, VDdSN23, WAS^{+}23, YLF^{+}24, ZDS^{+}22, ZFCZ25, CMT^{+}24]. NeuroHSMD [MFOM23]. Neuromorphic [MFOM23]. Next [LGW^{+}14]. Nine [NW11]. Nine-Context [NW11]. Ninth [VDG^{+}24]. nm [TMLS21]. NN [FC24]. NoC [KG17, KP14]. NoC-Based [KP14]. NoCs [GWPK20, MLPK22]. Non [CAG^{+}22, FC24, KB23, SB25]. Non-deterministic [KB23]. Non-Linear [FC24, SB25]. Non-uniform [CAG^{+}22]. Normalised [FLM^{+}17, FL20]. Note [Che22]. Novel [AHAM^{+}19, CCZ^{+}24, DNL19, EWL15, IGM^{+}20, RNTW22, VL11, SPS12]. Novelty [FYZ^{+}25]. NPN [ZWM19]. Number [VDG^{+}24, LSP^{+}23, PRV21, RVHP16, SBC10, TL08, Tho15]. Numerical [SLH^{+}10, USY17]. NVMe [WZL24]. NX [LNGP22].

O [MHS09, RGCL16]. Object [AIH^{+}24, GWS23]. Octavo [LA17]. ODoST [YBS16]. OFDM [BATM22, SAD10]. off [LW08]. Offs [SAD10]. On-Chip [LL12, GNM^{+}15]. On-cloud [EAAAA19]. Online [DHL^{+}18, GGR^{+}18]. Only [BDX^{+}19]. onto [SFNP23, SSF^{+}13]. Open [FL16, KHK24, TDH^{+}22, ZML^{+}22, SGNB08]. Open-Source [FL16, KHK24, ZML^{+}22, SGNB08]. OpenCL [MZLS20, TK16, WTS19].

OpenCL-Based [MZLS20]. OpenGL [BCW21]. Operate [US23]. Operating [AHL^{+}14, IBH^{+}15, VPPK20]. Operation [NBS13]. Operations [CKRG^{+}24, PSM^{+}14]. Operators [OBD13, SUK24]. Opportunities [CLL^{+}22, DVH^{+}15, LCL^{+}24]. Optical [BNW^{+}10, NW11]. Optimal [DSB09]. Optimization [BPCC09, CXG^{+}12, DSK15, DDH^{+}11, DLZ^{+}24, FKM^{+}25, KSCC10, LZ19, LP15, LT09, WYZ16, YB18, GCL^{+}23, MBJJ11, SMN^{+}23]. Optimizations [HLC^{+}15, IAG23, YSC^{+}23]. Optimized [AI22, CNK^{+}24, FZ23, GS10, LDJ^{+}17, NEPV24, RDC^{+}21, SBC10, SZZW23, SXZ^{+}25a, TYL^{+}23, VDdSN23, YFW^{+}17, ZCK22]. Optimizing [BAMR13, BC11, Kap16, LFN^{+}18, MZLS20, UCR^{+}19, WGGR16, WGGR17, YLF^{+}24]. Option [JTLC09]. Options [FT17]. OPU [TLC^{+}24]. Orchestrated [WZL24]. Order [BGSL17, WBR16, WBR18]. Oriented [JSLO24, TL11, VL11, WHQ^{+}08]. Oscillator [LYH^{+}23, YKBS10, LMG^{+}20, ZH12]. Oscillators [PRV21]. Out-of-Order [WBR16, WBR18]. Over-Clocking [DB15]. Overclocking [SBC15]. Overhead [DHL^{+}18, KGS15, NZS^{+}23]. Overlay [BYB23, IF23, TLC^{+}24, KB23]. Overlays [JCGW20, LA17]. own [RD11].

Packet [MS23]. Packing [AKA09]. Papers [Lee23, LAA^{+}17]. Parallel [AV13, BAG15, ES22, GFC^{+}25, GMZ^{+}23, IGM^{+}20, JB15, KTB^{+}24, MWBL21, SB15, SDM^{+}18, SSC16, TZWZ15, YOY17, GCL^{+}23]. Parallelising [CJWC23]. Parallelism [BBO^{+}25, GWXW21, INF^{+}14, KLD16, PVB13, QMG^{+}24, RNTW22, TYL^{+}23, CA11]. Parallelization [DRHM22, ZVS20]. Parallelized [LZ19]. Parallelizing [WAT15]. parameters [DW13]. Parametric [SC08]. Parser [LBRS16]. Parser-Based [LBRS16]. Part [ENPR22b]. Partial [BF23, EAGEG09, GFBF12, GGR^{+}18, GMZ^{+}23, RDB^{+}18, TWG^{+}20, XPN^{+}24, ZMH^{+}23, PDH11]. Partial-Reconfiguration [GGR^{+}18]. Partially [HHSC10, KMK^{+}10, HH13]. Particle [BG08, CNE^{+}15]. Partition [BS15]. Partitioned [SCL24]. Partitioning [APR^{+}22, LYZ^{+}18, TL11, TWG^{+}20, ZW25]. Passes [BBO^{+}25]. PASTA [KTB^{+}24]. Path [SLH^{+}25]. Pattern [LYZ^{+}18, YHK^{+}21]. Patterns [NI24]. Pay [EAAAA19]. Pay-per-use [EAAAA19]. Pedestrian [LZZ^{+}24]. Pentium(R) [LYS^{+}08]. Per-Module [DHL^{+}18]. Perfecto [HLL08]. Performance [APR^{+}22, BHF^{+}25, CH10, CKG^{+}10, EAGEG09, HZS^{+}25, HNM^{+}22, HNG09, HNS^{+}10, JZL^{+}24, JSG^{+}22, LLS^{+}23, LP15, MH15, MZLS20, PDH11, PMGL22, RKV23, SPM^{+}10, SDG12, SSC16, TL11, Tak17, TB10, TOS17, USY17, UNBR14, WPSI18, WBR18, WGGR17, YCV^{+}21, ZFCZ25, BC11, GP13, HGLS11, LXS^{+}24, MCL^{+}23, MPZ^{+}20, PANBI11, WWC^{+}22]. Performance-Oriented [TL11]. Perl [LT09]. Persistent [MHT^{+}21]. Perturb [GL08]. PEs [GRG08]. PGAS [AGY^{+}11]. Phylogenetics [CA23]. Phylogeny [ASPP22]. Physical [GCL^{+}23, INF^{+}14, JSLO24, MVGB15, SMOP15]. PIMap [LZ19]. Pinch [DGP10]. PipeArch [KA21]. Pipeline [SLH^{+}25]. Pipelined [KAL14, SV09, YOY17]. Pipelining [FZ23]. pixel [Oli12]. Placement [AHSS^{+}21, FRS^{+}15, GSJC13, HHSC10, JSG^{+}22, MVGB15, MSSM10, NZI22, Ste10, ZCK22, GL08]. Platform [BDX^{+}19, DSW^{+}24, JZL^{+}24, KSG11, KD19, LLZ^{+}24, NNY12, TJCP25, ZSP^{+}21]. Platform-aware [KSG11]. Platforms [CCF^{+}18, CBR^{+}14, GFL^{+}15, GKM^{+}12, LAF24, RMSK16, SAD10, YCV^{+}21]. Point [FL16, HU10, KD10, OBD13, RC10, SM25, USY17, WL10, WS16, dDELVP13, LLL^{+}23, WWC^{+}22]. Point-to-Point [KD10]. PointNet [SM25]. Polar [DLCJ20]. Policy [SDG12]. Population [CA23]. Port [AEM^{+}24, PVA^{+}09]. Portability [KGS15]. Portable [WS16, ZBR12]. Ported [LLO^{+}14]. Positive [AVCP20]. Post [HBXA23, MCL^{+}23, WBCW24]. Post-quantum [HBXA23, MCL^{+}23]. Post-Radiation [WBCW24]. Potential [ACFm^{+}25, AEM^{+}24, CZD^{+}25, MS25]. POWer [KGS^{+}12, CXG^{+}12, DHL^{+}18, KBM09, KCC^{+}14, KP14, LAL13, MMMT09, MDL^{+}23, NZS^{+}23, SLH^{+}10, TJCP25, UNBR14, WGGR17, ZBC^{+}09, CJH^{+}24, EA11, LW08, KGS^{+}12]. power-aware [EA11]. Power-Efficient [SLH^{+}10]. POWer-EmulatoR- [KGS^{+}12]. POWER-MODES [KGS^{+}12]. Powered [JZL^{+}24]. PQA [ACFm^{+}25]. PQC [BHXJ24, BHF^{+}25]. Practical [CPFM21, OROS^{+}19]. Precision [FL16, KGC^{+}24, LGD^{+}14, WL10, WSZ^{+}25, YHK^{+}21, Oli12, WWC^{+}22]. Predict [AHSS^{+}21]. Predicting [MOG^{+}13]. Prediction [HNG09, KZB23, SDGL^{+}22, ZQBL25, AHAM^{+}19, HGLS11]. Preemption [RDB^{+}18]. Preserving [PVA^{+}09, RHLK18]. Pricing [FT17, JTLC09, KLC11]. Primitives [HLN^{+}10]. Priority [BAG15, KVK^{+}11]. PRISA [MS25]. Privacy [AKM^{+}25, RHLK18]. Privacy-Preserving [RHLK18]. Private [WHJ^{+}23]. Probabilistic [CSS^{+}23]. Problem [GB11, GPP08]. Problems [KM10]. Process [DB15, RDB^{+}18, SB08, SRR23, TMLS21, LKJ^{+}11, SC11]. Processing [ASPP22, BCW21, BDX^{+}19, BHB14, CCT^{+}22, CNK^{+}24, CA23, DRF24, IABV15, Kap16, KA21, LP15, LAF24, MVGB15, SLL^{+}20, SKH^{+}22, SDM^{+}18, SSC16, WAT15, YEC^{+}09, ZZB^{+}20, ZBB^{+}16]. Processor [CBFM14, KB23, KCC^{+}14, KD19, LA17, MWK^{+}12, PWP^{+}16, Tak17, TLC^{+}24, WBR16, YEC^{+}09, Tak12]. Processor-logic [KD19]. Processors [FLM^{+}17, GFBF12, KGC^{+}24, SFNP23, VTN09, WPSI18, WBR18, IYY^{+}11, LJS11, MLFS22]. Product [ACFm^{+}25, VDdSN23]. Production [UHU09]. Productively [HZS^{+}25]. Productivity [KGS15]. Profiling [EWL15]. Profiling-Based [EWL15]. Program [PD15]. Program-Invariant [PD15]. ProgramGalois [LYC^{+}24]. Programmability [GKM^{+}12]. Programmable [AC14, BCW21, CAPA^{+}09, DPHT19, GS10, HH24, HNM^{+}22, LYC^{+}24, OWMZ11, SCC10, WSZ^{+}25, dBVW^{+}24]. Programmers [LFS22]. Programming [GCL^{+}23, KTB^{+}24, LUX^{+}21, MWK^{+}12, SUK24, SPM^{+}10, WGGR17, AGY^{+}11]. Programs [KTB^{+}24]. Property [HTG25]. Prospects [LUX^{+}21]. Protected [BDGH15, SG15]. Protection [EAAAA19]. Protein [JLB^{+}08]. Protocol [ADSH18, SSK^{+}23]. ProtoFlex [CPN^{+}09]. Prototype [PCC^{+}25]. Provably [GWPK20]. Pruning [LLYC24]. PTME [SXZ^{+}25b]. PUF [GHO17, SKB^{+}22]. PUFs [MKP09]. Pulsar [WTS19]. Pulses [PEM^{+}09]. Purpose [AJYH18, GFBF12, GPP08, LGD^{+}14, ZDS^{+}22]. Purposes [BHI15]. Puzzle [GPL24].

QiCells [GKGS23]. Quantifying [YLF^{+}24]. Quantization [ACFm^{+}25, YHK^{+}21]. Quantized [BPF^{+}18, CMT^{+}24]. Quantum [GKGS23, HBXA23, MCL^{+}23]. Quasi [TB10]. Quasi-Monte [TB10]. Query [LAF24, ZBB^{+}16]. Queue [BAG15]. Quick [MLFS22]. Quick-Div [MLFS22]. Quipu [MOG^{+}13].

R [BPF^{+}18, PP10, dBVW^{+}24]. R-Blocks [dBVW^{+}24]. R3TOS [IBH^{+}15]. Rack [PCC^{+}25]. Radiation [KW22, WBCW24]. Radiations [SGM09]. Radio [PEM^{+}09]. Radix [LYC^{+}24]. Radix-4 [LYC^{+}24]. Rail [HF14]. RAiSD [AVCP20]. RAiSD-X [AVCP20]. RAMs [TYB18]. Random [LSP^{+}23, LOM10, PRV21, RVHP16, SBC10, TL08, Tho15]. Randomized [DL09]. Range [MS23]. Range-enhanced [MS23]. RankBoost [XCG^{+}09]. Ranking [KY18]. Ransomware [GKA^{+}24]. Rapid [HNG09, RGGW10, WPSI18, WCK21]. RapidLayout [ZCK22]. RapidStream [GMZ^{+}23]. RAT [HNG09]. Rate [IABV15]. RAW [GC13, RVHP16]. RC [HNG09]. RD [GKA^{+}24]. RD-FAXID [GKA^{+}24]. Reactor [JSLO24]. Reactor-Oriented [JSLO24]. Real [ABCC09, BATM22, BHB14, CNK^{+}24, CJH^{+}24, GKLLA23, GWS23, GNM^{+}15, HHSC10, INF^{+}14, IBH^{+}15, LLL^{+}23, LZZ^{+}24, RDB^{+}18, RMSK16, RHLK18, SLL^{+}20, TKB^{+}24]. Real-Time [ABCC09, BATM22, BHB14, CNK^{+}24, HHSC10, INF^{+}14, IBH^{+}15, RDB^{+}18, RMSK16, RHLK18, CJH^{+}24, GKLLA23, GWS23, LLL^{+}23, LZZ^{+}24, TKB^{+}24]. Real-world [SLL^{+}20]. Realizable [RGCL16]. Realizable-Utilization [RGCL16]. Receiver [BATM22]. Recipes [DGP10]. Recoding [ZCL15]. Recognition [DDH^{+}11, LZZ^{+}24]. reconfigurability [SC11]. Reconfigurable [ASGY12, ADSH18, AV13, ATJZ16, BBND10, Bec14, BHI15, BDX^{+}19, BHB14, CBC^{+}12, CNZ^{+}18, CZZ23, CTH16, CNE^{+}15, CH10, CBR^{+}14, CKG^{+}10, DC16, DGP^{+}15, DSB09, DDB^{+}10, EAGEG09, ERO25, FT17, FHL^{+}24, FKS^{+}12, GFL^{+}15, GKM^{+}12, GWS23, GC13, GdLlG^{+}14, HCOB13, HHSC10, HNS^{+}10, HLN^{+}10, IZO^{+}10, IGM^{+}20, IBH^{+}15, JCG^{+}12, JTLC09, KMK^{+}10, KCC^{+}14, LP22, LYS^{+}08, MS23, MH15, MKP09, MWK^{+}12, MSSM10, NNY12, NBS13, NJLW14, Oli12, PP10, PD15, PFC15, QMG^{+}24, QNF^{+}23, RGGW10, RGCL16, RMSK16, RUC11, SGW20, SGC21, SPM^{+}10, SJT09, SAD10, SDGL^{+}22, TL11, TWG^{+}20, THK12, TL08, UAS16, UCR^{+}19, UHU09, VL11, VTN09, VG14, WL10, WCK21, WMG^{+}10, YCV^{+}21, ZBB^{+}16, dDELVP13, AGY^{+}11, BG08, GDHG11, HLL08, HH13, IYY^{+}11, KSG11, ZH12, HCGG24].

Reconfiguration [BDX^{+}19, BF23, DS15, EAGEG09, GFBF12, GGR^{+}18, GMZ^{+}23, HNS^{+}10, JSC14, KD10, KS20, LCS14, LZF^{+}10, NW11, NCJ^{+}15, PPR^{+}10, RLY^{+}15, RDB^{+}18, VVZL25, VMV15, XPN^{+}24, ZBC^{+}09, ZMH^{+}23, NSS^{+}11, PDH11]. ReconROS [LP22]. Reconstruction [ASPP22, TZWZ15]. ReCoSoC [Hüb12]. ReCoSoC’12 [VG14]. Recovery [DSW^{+}24, ZNA^{+}18]. Rectangular [SWT^{+}22]. Recurrent [CJH^{+}24, QNF^{+}23]. Recursive [PWP^{+}16]. ReDCrypt [RHLK18]. Reduce [PSM^{+}14]. Reducing [BAMR10, TOS17]. Reduction [CW09, SLH^{+}10]. References [BAMR13]. Region [MS25]. Region-based [MS25]. Registration [SM25]. Regular [LT09, SXZ^{+}25b, YHK^{+}21]. Regulator [AV13]. Regulators [MLPK22]. Related [OBD13]. relating [DW13]. Reliability [DSK15, GHO17]. Reliable [IBH^{+}15, JCG^{+}12]. Relocatable [HHSC10]. Remarn [QNF^{+}23]. Remote [BCE^{+}10, MDL^{+}23, MCN12, SSK^{+}23, VMV15]. Repair [GGR^{+}18]. Replacing [HBA^{+}15]. Reprogrammable [FC24]. Request [AI22]. ReRAM [FC24]. ReShape [NBS13]. Reshaping [BHI15]. Resilience [SGW20]. Resilient [INF^{+}14, SGC21]. Resistant [HF14]. Resolution [ABCC09, CZZ23, SGM09]. Resource [CCT^{+}22, JMGI23, MHS^{+}19, SSK^{+}23, SMN^{+}23, WML^{+}24, HZW^{+}13]. Resource-constrained [MHS^{+}19, SMN^{+}23]. Resource-Efficient [CCT^{+}22, SSK^{+}23, WML^{+}24]. Resources [MOG^{+}13]. Restricted [KAL14]. resynthesis [MBJJ11]. Retargetable [UAS16]. Rethinking [MLFS22, RRW^{+}22]. Reusable [JRHK15]. Reuse [LYZ^{+}18, dMdLC23]. Review [ILP24]. RF [GKLLA23]. RFSoC [GKGS23]. RFSoC-based [GKGS23]. RIFFA [JRHK15]. Ring [BHXJ24, BHF^{+}25, HBXA23, PRV21, YKBS10, LYH^{+}23, ZH12]. Ring-Binary-LWE-Based [BHXJ24, BHF^{+}25, HBXA23]. RIPL [SDM^{+}18]. RISCV [ZZB^{+}20]. RISCV-based [ZZB^{+}20]. RIVER [BHB14]. RNA [MCC10]. Road [UHU09]. Robotics [LP22, NSS^{+}11]. Robust [ABCC09, GKLLA23]. Robustness [LZF^{+}10, SRR23, YKBS10]. RoCEv2 [SSK^{+}23]. Rotation [ZCL15]. Routability [AHAM^{+}19, AHSS^{+}21, JCCM09, LRA13, PB18, DW13, EA11]. Routability-Driven [LRA13, EA11]. Routability-prediction [AHAM^{+}19]. Routed [KG17]. Router [LL12, ZML^{+}22]. Routers [MLPK22, GP13]. Routing [CW09, FRS^{+}15, IZO^{+}10, KA17, SZZW23, SB08, WYZ16, ZVS20, GL08, LKJ^{+}11, RD11]. RTL [DVH^{+}15]. RTR [ZBC^{+}09]. Run [DNL19]. Run-Time [DNL19]. Runtime [BF23, EAGEG09, FRS^{+}15, LCS14, NCJ^{+}15, PPR^{+}10, ZBC^{+}09]. RWRoute [ZML^{+}22].

s [TKB^{+}24]. Safe [AB20, BHI15]. Sampling [PRV21]. Samsung [KCC^{+}14]. SARFUM [BCE^{+}10]. SASA [TYL^{+}23]. SAT [KM10]. Satisfiability [FZ23]. Satisfiability-based [FZ23]. SATTA [DGP^{+}15]. SCA [HF14]. SCA-Resistant [HF14]. Scala [SP20]. ScalaBFS2 [LXS^{+}24]. Scalability [RDC^{+}21, Tak17]. Scalable [BYB23, ASPP22, CPN^{+}09, DRF24, GCL^{+}23, ILP24, JYJC23, KTB^{+}24, LML^{+}23, MBJJ11, MYST25, NEPV24, OWMZ11, RKV23, SLH^{+}10, TYL^{+}23, ZSP^{+}21, ZBR12]. Scalar [TOS17]. Scale [CSK17, LLZ^{+}24, SXZ^{+}25a, YCV^{+}21, CCZ^{+}24, KD21]. Scaling [AZM^{+}19, APR^{+}22, DCS24, MPK22, MKP23, NNY12, LKJ^{+}11]. Scanning [LMG^{+}20]. Scavenger [YFW^{+}17]. SCF [ASGY12]. Scheduling [BAMR10, CBR^{+}14, HHSC10, HNS^{+}10, OROS^{+}19, TWG^{+}20, WBR18, CJWC23]. Scheme [EAAAA19, ZHL^{+}21]. Schemes [OBD13, SV09]. Science [UHU09]. Scientific [RUC11, SKJ22]. Score [MYST25]. SDK [TK16]. SDM [LL12]. Search [CNZ^{+}18, MS25, WTS19, XCG^{+}09, YXC^{+}11]. Searching [PEM^{+}09]. Secondary [MCC10]. Secondary-Structure [MCC10]. Secret [WZY^{+}24]. Section [Bak18, Bec14, BE19, CS17, CPW18, Deh20, FB24, GKK24, Ien23, Leo22, LH23, MST22, Sha22, SLD23, Che11, GC13]. Secure [GFBF12, KGT19, MKP09, VMV15]. Security [AKM^{+}25, BCE^{+}10, BE19, GHWS22, HLN^{+}10, KGS^{+}12, SJT09, SMOP15]. Segmentation [LFN^{+}18, SGC21]. Selection [AVCP20, CNZ^{+}18, ZMH^{+}23]. Self [AHL^{+}14, AV13, Bec14, BKT14, BHI15, GGR^{+}18, HCOB13, LMG^{+}20, LZF^{+}10, NJLW14, OBD13, PMC^{+}14, WSC09, DGP^{+}15]. Self-Adaption [BHI15]. Self-Adaptive [PMC^{+}14, DGP^{+}15]. Self-Alignment [OBD13]. Self-Aware [BKT14, NJLW14]. Self-Awareness [AHL^{+}14, Bec14]. Self-Healing [BHI15]. Self-Measurement [WSC09]. Self-Monitoring [GGR^{+}18]. Self-oscillator [LMG^{+}20]. Self-Reconfigurable [HCOB13]. Self-Reconfiguration [LZF^{+}10]. Self-Tuning [AV13]. Semantic [SGC21]. Semi [KVV24]. Semi-Global [KVV24]. Semisynthetic [GL08]. Senju [DCS24]. sensing [ZH12]. Sensor [LZZ^{+}24, MDL^{+}23]. Sentential [CSS^{+}23]. Separable [LLYC24, LP15]. Separation [WBC16]. Sequence [JLB^{+}08]. Sequences [PBPLA17]. Sequential [ES22]. Serial [UCR^{+}19]. Series [KZB23, ZQBL25]. Serve [AI22]. Server [MA24]. Set [GB11, WHJ^{+}23]. SHA [MAK^{+}12]. SHA-256 [MAK^{+}12]. Shared [MSF16]. Sharing [JMGI23, WZY^{+}24]. Shifter [WAT15]. Shifting [DSB09]. SHMEM [AGY^{+}11]. Short [CAG^{+}22]. Shrinkage [WAS^{+}23]. Side [DSW^{+}24, GER19, GTS23, SG15, TJCP25]. Side-Channel [DSW^{+}24, GTS23, SG15, TJCP25, GER19]. Signal [BHB14, DDH^{+}11, SFNP23, SSC16]. Signature [MCL^{+}23, ZHL^{+}21]. signatures [LJS11]. Significant [LAA^{+}17]. Silicon [UHU09, WHQ^{+}08]. SILVIA [BBO^{+}25]. Similarity [KZB23, WZY^{+}24]. Simple [US23]. Simple-to-Operate [US23]. Simulation [AB23, CZ09, MVGB15, RGGW10, TB10]. Simulations [CH10, CPN^{+}09, MPK22, MHS09, SLH^{+}10, SC08]. Single [CSK17, DCS24, LKJ^{+}11]. single-driver [LKJ^{+}11]. Sizing [HDMN24, JSG^{+}22]. Skew [SB08]. Skip [WML^{+}24, DCL^{+}22]. Slices [AGM^{+}22]. Sliding [CFBS15, GS23, SSC16]. Sliding-Window [CFBS15, SSC16]. Smaller [YSC^{+}23]. SmartSSD [SZKR22]. SmartSSDs [TES24]. SMP [MSF16]. SoC [LBRS16]. SoCs [BF23, GHWS22, IF23, MCD^{+}18, WBCW24, WGGR16]. Soft [AB14, AJYH18, CBFM14, DVK15, Kap16, KGC^{+}24, LA17, LFS22, MLFS22, PD15, TOS17, WPSI18, WBR16, WBR18, YEC^{+}09, LJS11]. Soft-Error [DVK15, PD15]. Soft-Processor [LA17]. Soft-processors [MLFS22]. Software [AB23, HHSC10, JSLO24, KHK24, LUX^{+}21, LVC24, LFS22, ZHL^{+}21, HH13, SC11]. Software-defined [LUX^{+}21]. Software-like [AB23]. Software/Hardware [ZHL^{+}21]. Solver [GPL24, ZBR12]. Solving [GFL^{+}15, GPP08, KM10]. Sort [SZKR22]. Sorting [SP20, SXZ^{+}25a]. Source [DC16, FL16, KHK24, SGNB08, ZML^{+}22]. Space [CCZ^{+}24, DCS24, JCG^{+}12, JYJC23, LZF^{+}10, LT09, MZLS20, SGW20, SGC21, SCY^{+}23, HLL08]. Sparkle [XJXQ24]. Sparse [DDB^{+}10, LCL^{+}24]. Sparsification [WAS^{+}23]. Sparsity [YHK^{+}21]. SPARTA [GFC^{+}25]. Spatial [CZD^{+}25, MS25, SGM09, TYL^{+}23, ZG16]. Special [Bak18, Bec14, BE19, CS17, CPW18, CWBD09, DH08, DC16, Deh20, ENPR22a, ENPR22b, FB24, GKK24, GPP08, Hüb12, Ien23, KZ23, KS24, LGD^{+}14, Leo22, LH23, MST22, Sha22, SML^{+}24, SLD23, VG14, Che11, GC13, AC14]. Special-Purpose [GPP08, LGD^{+}14]. Specialization [DVH^{+}15]. Specializing [MHT^{+}21]. Specific [DDH^{+}11, PMKM11, RLM^{+}17, WYZ16, BBO^{+}25, LJS11, SLL^{+}20, SFT^{+}23]. Spectral [LML^{+}23]. Speculation [CTH16, SXZ^{+}25b, THK12]. Speculative [SLH^{+}25]. Speed [BS15, HBA^{+}15, NW11, ZBC^{+}09]. SPHINCS [LVC24]. Spiking [MFOM23, PFL22]. SpMM [XJXQ24]. SQL [LAF24, ZBB^{+}16]. SQL2FPGA [LAF24]. Squares [FLM^{+}17, FL20, PWP^{+}16]. SRAM [AL16, LZF^{+}10, Ste10]. SRAM-Based [AL16, LZF^{+}10, Ste10]. SRCS’12 [Bec14]. SRP [KCC^{+}14]. stack [DFB^{+}22]. stacked [IAG23]. staged [KVK^{+}11]. Stall [GWPK20]. Stall-Free [GWPK20]. Standard [KA17]. Standing [NGM^{+}23]. State [LCL^{+}24, ZG16, GDHG11]. State-of-the-Art [LCL^{+}24]. States [BAMR13]. Static [CW09, DRHM22, LAL13]. Statistical [CXG^{+}12, MOG^{+}13, SB08, WSDH23]. Status [LUX^{+}21]. Stealing [RLM^{+}17]. Stencil [DCS24, JB15, RDC^{+}21, TYL^{+}23]. Step [BPFD11]. Stereo [JM14, KVV24]. Stopping [AB20]. Storage [DLBM18, KHK24, SZKR22, TBS^{+}25, WZL24]. Storage- [DLBM18]. STR [FYZ^{+}25]. Strategies [MCL^{+}13, PCFM23]. Strategy [DCL^{+}22, KMK^{+}10]. Stratix [SMOP15, KVV24, LNGP22, SMOP15]. Stream [CNK^{+}24, GS23, PBPLA17, SCY^{+}23, YSC^{+}23]. Streaming [DD18, IF23, LBL23, PVB13, RMSK16]. Streams [SKW^{+}21, USY17]. Strega [MA24]. Strong [MPK22]. Strongly [ATJZ16]. Structure [LGD^{+}14, MYST25, MCC10]. Structures [DL09]. Study [BNW^{+}10, NSS^{+}11]. Submitted [Lee23]. Substream [BFBN^{+}20]. Substream-Centric [BFBN^{+}20]. Successes [CLL^{+}22]. Super [ABCC09, CZZ23]. Super-Resolution [ABCC09]. Supercomputer [DDB^{+}10]. Supercomputing [VDG^{+}24, UHU09, AGY^{+}11]. Superconducting [GKGS23]. SuperDragon [TZWZ15]. Superscalar [WBR18]. Superword [BBO^{+}25]. Superword-Level [BBO^{+}25]. Support [GdLlG^{+}14, KTB^{+}24, MSF16, PSM^{+}14, PBPLA17]. Supporting [DNL19, QMG^{+}24, SSF^{+}13]. Supports [KGC^{+}24]. Suppression [MHK^{+}08]. Survey [AKM^{+}25, CA23, GB11, GZY^{+}18, NG25, PCFM23, PDH11]. SW [MCL^{+}23]. SW/HW [MCL^{+}23]. Switch [BMR16, KA21, NI24, PMGL22, SY24, SKW^{+}21]. Switch-based [SKW^{+}21]. Switch-Blocks [NI24]. Switched [AL16, LL12, MKP23]. Switching [AB20, US23]. Symbol [BDX^{+}19]. Symbol-Only [BDX^{+}19]. Symmetric [GFBF12, WZY^{+}24]. Symmetries [ZWM19]. Symposium [DH08]. Synchronous [GKM^{+}12, PVA^{+}09]. SyncNN [PFL22]. Synergies [MCD^{+}18]. Synthesis [BAMR10, BAMR13, BPCC09, CJWC23, CSS^{+}23, DD18, DLZ^{+}24, FKM^{+}25, GFC^{+}25, GWXW21, GdLlG^{+}14, HH24, HDMN24, HLC^{+}15, JYJC23, LUX^{+}21, LN23, OROS^{+}19, RBR16, SFNP23, SLH^{+}25, WBC16, PANBI11]. Synthesis-Generated [HLC^{+}15]. Synthesizable [KA17, WHQ^{+}08]. System [AVCP20, BCW21, CPN^{+}09, ES22, GSJC13, GS10, HH24, IBH^{+}15, JM14, JB15, LZZ^{+}24, LGW^{+}14, MSF16, TZWZ15, VPPK20, WBR16, ZZB^{+}20, ZBR12]. System-Level [GSJC13]. System-on-Chip [GS10]. Systematic [ILP24]. SystemC [HCGG24, HLL08]. SystemC-based [HLL08]. Systems [ASGY12, AI22, Bec14, BKT14, BHI15, CNE^{+}15, CH10, CA23, ERO25, GMBC17, GdLlG^{+}14, HHSC10, HLN^{+}10, IGM^{+}20, INF^{+}14, JSLO24, Kap16, KMK^{+}10, KBT09, LP22, LFS22, MH15, MCN12, NBS13, NJLW14, PMC^{+}14, PVA^{+}09, RGGW10, SGW20, SJT09, SML^{+}24, VG14, ZQ19, ZNA^{+}18, ZW25, HGLS11, HH13, PDH11, ZH12]. Systems-on-Chip [GdLlG^{+}14, VG14]. Systolic [LML^{+}23, NEPV24, ZCL16, ZCK22].

Table [IABV15, Tho15]. Table-Hadamard [Tho15]. Tailor [WML^{+}24]. TAPA [GCL^{+}23]. Targetable [KA17]. Targeting [DDH^{+}11, SFNP23, TL08]. TAS [ZBC^{+}09]. TAS-MRAM-Based [ZBC^{+}09]. Task [ASGY12, AB20, CTH16, GCL^{+}23, HNS^{+}10, KTB^{+}24, PVB13]. Task- [PVB13]. Task-Level [ASGY12]. Task-Parallel [KTB^{+}24, GCL^{+}23]. Tasks [HHSC10]. TATAA [WSZ^{+}25]. TCAMs [dMdLC23]. TDC [FYZ^{+}25]. TDF [DGP^{+}15]. TDM [LL12]. TDM-Based [LL12]. Technique [LN23]. Techniques [AKA09, KBT09, MKP09, OVI^{+}12]. Technology [AC14, BCW21, JCCM09, KVV24, LZ19, PWP^{+}16, KVK^{+}11]. Telescope [PEM^{+}09]. Temperature [DGP^{+}15, DB15]. Temperature-Based [DGP^{+}15]. Temporal [TYL^{+}23]. Tenant [AKM^{+}25, KGT19, MKSB22]. Tensor [AGM^{+}22]. Termination [TTY25]. Ternary [PBBP18, TKH^{+}19]. TERO [SRR23]. TERO-TRNGs [SRR23]. Terrestrial [KW22]. Test [HNG09, IYY^{+}11]. Testing [AZM^{+}19]. Them [FZ25]. Thermal [KP14]. Thousands [AVCP20]. Threaded [GFC^{+}25, QNF^{+}23]. Throughput [AIH^{+}24, CNK^{+}24, FYZ^{+}25, LDJ^{+}17, RC10, SSK^{+}23, MAK^{+}12]. Throughput-Optimized [CNK^{+}24, LDJ^{+}17]. ThunderGP [CCT^{+}22]. Tile [SY24]. Tile-Based [SY24]. TILT [TOS17]. Time [ABCC09, BPFD11, BATM22, BHB14, CNK^{+}24, DCS24, DNL19, DSW^{+}24, HHSC10, INF^{+}14, IBH^{+}15, KZB23, PPR^{+}10, RDB^{+}18, RMSK16, RHLK18, ZQBL25, CJH^{+}24, GKLLA23, GWS23, LLL^{+}23, LZZ^{+}24, RD11, TKB^{+}24]. Time-to-Digital [DNL19, DSW^{+}24]. Timed [PVA^{+}09]. Timing [CXG^{+}12, GGR^{+}18, GNM^{+}15, LRA13, MWL^{+}15, Ste10, WYZ16, ZML^{+}22]. Timing- [LRA13]. Timing-Driven [MWL^{+}15, ZML^{+}22]. Titan [MWL^{+}15, PP10]. Titan-R [PP10]. TMR [ZNA^{+}18]. Today [CLL^{+}22]. Tolerance [DVK15, JCG^{+}12, WBCW24]. Tolerant [BKT14, RLY^{+}15]. Tool [BF23]. Tools [BKT14, LKJ^{+}11]. Toolset [KMK^{+}10]. Topgun [WHJ^{+}23]. Topology [RLY^{+}15]. Torus [KG17]. TPM [GHWS22]. TR [GDHG11]. TR-FSM [GDHG11]. Trace [DSK15]. Trace-Driven [DSK15]. Track [Lee23, Che22]. Tracking [GWS23]. Trade [SAD10, LW08]. trade-off [LW08]. Trade-Offs [SAD10]. Tradeoff [CFBS15]. Tradeoffs [UNBR14]. Traffic [OWMZ11]. Training [ZQBL25]. Transaction [AB23]. Transaction-based [AB23]. transactional [LJS11]. Transfer [DD18]. Transformable [WSZ^{+}25]. Transformation [LYC^{+}24]. Transformer [WSZ^{+}25]. Transforms [SP20]. Transient [PEM^{+}09]. Transistor [KY18]. Transition [GDHG11]. Transition-Based [GDHG11]. Traversal [FRS^{+}15]. Tree [BAG15, DCL^{+}22, JTLC09, OKA19, PANBI11]. Tree-Based [JTLC09]. Trees [CAPA^{+}09, HDN^{+}24, PBBP18]. TRETS [Bec14, Che22, DH08, Lee23]. TRIP [GGR^{+}18]. TRNG [FYZ^{+}25, LYH^{+}23]. TRNGs [SRR23, YKBS10]. True [LSP^{+}23, PRV21]. Trust [AKM^{+}25, DL09]. Trust-Based [DL09]. TrustZone [GHWS22]. Tunable [SKB^{+}22]. Tune [DSW^{+}24]. Tuning [AV13, NJLW14]. Tunnel [TYB18]. Turn [DSW^{+}24]. Turnaround [JCGW20]. Two [DL09]. Two-Level [DL09].

ULP [KCC^{+}14]. ULP-SRP [KCC^{+}14]. Ultra [KCC^{+}14]. UltraScale [KVV24, LMG^{+}20]. UML [GdLlG^{+}14]. UML/MARTE [GdLlG^{+}14]. Unary [FAB22, MWBL21]. Understanding [CZD^{+}25]. Unfriendly [AI22]. Unified [BHF^{+}25, WS16, ZDS^{+}22]. uniform [CAG^{+}22]. UNILOGIC [IGM^{+}20]. Unit [BCW21, PP10, RUC11]. Units [VDdSN23, dDELVP13]. Unpredictable [BAMR13]. Unrolling [DSB09, DPHT19, TKH^{+}19]. Unstructured [BGSL17]. UNTANGLED [MCL^{+}13]. Update [BCE^{+}10]. Usage [GHWS22]. use [BC11, EAAAA19]. Using [BAG15, BCW21, CSK17, CPN^{+}09, CAG^{+}22, DL09, FC24, FAB22, FK08, FRS^{+}15, GNM^{+}15, VDG^{+}24, HNM^{+}22, JCGW20, LP15, NW11, PWP^{+}16, RLY^{+}15, RLM^{+}17, RHLK18, SY24, SDGL^{+}22, TK16, WTS19, ZZB^{+}20, ZCK22, ZWM19, dMdLC23, JSC14, KSCC10, MHK^{+}08, PD15, PMKM11, SUK24, SB25, VVZL25]. Utilization [CZZ23, RGCL16, SXZ^{+}25a]. Utilizing [SFNP23].

V [KVV24]. Validation [IPC14]. Value [THK12, ZG16]. Variability [TMLS21]. Variable [FL16, IZO^{+}10, WL10, Oli12]. Variable-Grain [IZO^{+}10]. Variable-Precision [FL16]. Variant [XRC23]. Variation [DB15, MHK^{+}08, SB08, SRR23]. Variation-Aware [SB08]. Variations [SC08]. VBSME [Oli12]. VCSN [US23]. Vector [DDB^{+}10, Kap16, YEC^{+}09, BC11]. Verilog [KA17]. Verilog-to-Routing [KA17]. Versal [JZL^{+}24, ZLY^{+}24]. Versatile [PBPLA17]. Very [CJH^{+}24]. VFloat [WL10]. via [AIH^{+}24, AB23, BBO^{+}25, CBC^{+}12, GHWS22, LYZ^{+}18, LZ19, MKP23, RDC^{+}21, SKJ22, TTY25]. Video [ABCC09, LP15]. Virtex [AKA09]. Virtex-5 [AKA09]. Virtual [HZW^{+}13, US23, GP13]. Virtualizable [HH13]. Virtualization [OVI^{+}12, ZDS^{+}22, ZSP^{+}21]. Vision [FHL^{+}24, JM14, NSS^{+}11]. VLIW [LGD^{+}14]. VM [GTS23]. Voltage [AZM^{+}19, DB15, MDL^{+}23, NNY12]. Volume [CKRG^{+}24]. VPR [LKJ^{+}11]. vs [BYB18, TB10]. VTR [LGW^{+}14, MPZ^{+}20].

Wait [BAMR13]. Wave [SV09]. Wave-Pipelined [SV09]. Waveform [TKB^{+}24]. WDDL [MMMT09]. Weather [SDGL^{+}22]. Web [CNZ^{+}18, XCG^{+}09, YXC^{+}11]. Weight [PBBP18]. Where [FZ25, NGM^{+}23]. Width [KY18]. WiMax [SAD10]. Window [CFBS15, SSC16, WYZ16]. Windows [SWT^{+}22, GS23]. Winograd [YHK^{+}21]. Winograd-based [YHK^{+}21]. Wireless [CNK^{+}24, GKLLA23]. WireMap [JCCM09]. Wires [GER19, SZZW23]. Within [SC08, MKSB22]. Within-Die [SC08]. without [NI24, PB18]. Work [RLM^{+}17]. Workflow [HDMN24]. Workloads [ERO25, VPPK20]. Workshop [GC13, VG14, Bec14]. world [SLL^{+}20]. Wotan [PB18].

X [AVCP20]. xDNN [DWN^{+}22]. XGBoost [GKA^{+}24]. Xilinx [AEM^{+}24, LMG^{+}20]. XVDPU [JZL^{+}24].

Years [LAA^{+}17]. Yield [SC08].

Zynq [KVV24, KD19, MCD^{+}18]. ZyPR [BF23].

References

Abdelfattah:2014:NCF

[AB14] Mohamed S. Abdelfattah and Vaughn Betz. Networks-on-chip for FPGAs: Hard, soft or mixed? ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):20:1–20:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Attia:2020:FFI

[AB20] Sameh Attia and Vaughn Betz. Feel free to interrupt: Safe task stopping to enable FPGA checkpointing and context switching. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(1):3:1–3:27, February 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3372491.

Attia:2023:TSL

[AB23] Sameh Attia and Vaughn Betz. Toward software-like debugging for FPGAs via checkpointing and transaction-based co-simulation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):31:1–31:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3552521.

Arora:2023:CDC

[ABB^{+}23] Aman Arora, Atharva Bhamburkar, Aatman Borda, Tanmay Anand, Rishabh Sehgal, Bagus Hanindhito, Pierre-Emmanuel Gaillardon, Jaydeep Kulkarni, and Lizy K. John. CoMeFa: Deploying compute-in-memory on FPGAs for deep learning acceleration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):50:1–50:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3603504.

Angelopoulou:2009:RRT

[ABCC09] Maria E. Angelopoulou, Christos-Savvas Bouganis, Peter Y. K. Cheung, and George A. Constantinides. Robust real-time superresolution on FPGA and an application to video enhancement. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(4):22:1–22:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Anderson:2014:ISI

[AC14] Jason Anderson and Kiyoung Choi. Introduction to the Special Issue on the 11th International Conference on Field-Programmable Technology (FPT’12). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):18:1–18:??, August 2014. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Abouelhamayed:2025:PEP

[ACFm^{+}25] Ahmed Abouelhamayed, Angela Cui, Javier Fernandez-marques, Nicholas Lane, and Mohamed Abdelfattah. PQA: Exploring the potential of product quantization in DNN hardware acceleration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):6:1–6:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3656643.

Anandakumar:2018:RHA

[ADSH18] N. Nalla Anandakumar, M. Prem Laxman Das, Somitra K. Sanadhya, and Mohammad S. Hashmi. Reconfigurable hardware architecture for authenticated key agreement protocol over binary Edwards curve. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(2):12:1–12:??, November 2018. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Albartus:2024:MPX

[AEM^{+}24] Nils Albartus, Maik Ender, Jan-Niklas Möller, Marc Fyrbiak, Christof Paar, and Russell Tessier. On the malicious potential of Xilinx’s internal configuration access port (ICAP). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (2):26:1–26:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3633204.

Arora:2022:TSF

[AGM^{+}22] Aman Arora, Moinak Ghosh, Samidh Mehta, Vaughn Betz, and Lizy K. John. Tensor slices: FPGA building blocks for the Deep Learning era. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):46:1–46:??, December 2022. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3529650.

Aggarwal:2011:SMP

[AGY^{+}11] Vikas Aggarwal, Alan D. George, Changil Yoon, Kishore Yalamanchili, and Herman Lam. SHMEM+: a multilevel-PGAS programming model for reconfigurable supercomputing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3): 26:1–26:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Al-Hyari:2019:NCE

[AHAM^{+}19] Abeer Al-Hyari, Ziad Abuowaimer, Timothy Martin, Gary Gréwal, Shawki Areibi, and Anthony Vannelli. Novel congestionestimation and routability-prediction methods based on machine learning for modern FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(3):16:1–16:??, September 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3337930.

Agne:2014:SAM

[AHL^{+}14] Andreas Agne, Markus Happe, Achim Lösch, Christian Plessl, and Marco Platzner. Self-awareness as a model for designing and operating heterogeneous multicores. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):13:1–13:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Al-Hyari:2021:DLF

[AHSS^{+}21] Abeer Al-Hyari, Hannah Szentimrey, Ahmed Shamli, Timothy Martin, Gary Gréwal, and Shawki Areibi. A deep learning framework to predict routability for FPGA circuit placement. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(3):16:1–16:28, September 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3465373.

Asiatici:2022:RCS

[AI22] Mikhail Asiatici and Paolo Ienne. Request, coalesce, serve, and forget: Miss-optimized memory systems for bandwidthbound cache-unfriendly applications on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (2):13:1–13:33, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3466823.

Anupreetham:2024:HTF

[AIH^{+}24] Anupreetham Anupreetham, Mohamed Ibrahim, Mathew Hall, Andrew Boutros, Ajay Kuzhively, Abinash Mohanty, Eriko Nurvitadhi, Vaughn Betz, Yu Cao, and Jae-Sun Seo. High throughput FPGA-based object detection via algorithm-hardware co-design. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):1:1–1:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3634919.

AlKadi:2018:GPC

[AJYH18] Muhammed Al Kadi, Benedikt Janssen, Jones Yudi, and Michael Huebner. General-purpose computing with soft GPUs on FP-GAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):5:1–5:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ahmed:2009:PTV

[AKA09] Taneem Ahmed, Paul D. Kundarewich, and Jason H. Anderson. Packing techniques for Virtex-5 FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(3):18:1– 18:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ahmed:2025:MTC

[AKM^{+}25] Muhammed Kawser Ahmed, Maximillian Panoff Kealoha, Joel Mandebi Mbongue, Sujan Kumar Saha, Erman Nghonda Tchinda, Peter Esenju Mbua, and Christophe Bobda. Multitenant cloud FPGA: a survey on security, trust, and privacy. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):23:1–23:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Abdelhadi:2016:MSM

[AL16] Ameer M. S. Abdelhadi and Guy G. F. Lemieux. Modular switched multiported SRAM-based memories. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(3):22:1– 22:??, July 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Amano:2009:GEI

[AN09] Hideharu Amano and Tadao Nakamura. Guest editors’ introduction: ICFPT 2007. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2):7:1–7:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Alonso:2022:EDS

[APR^{+}22] Tobias Alonso, Lucian Petrica, Mario Ruiz, Jakoba Petri-Koenig, Yaman Umuroglu, Ioannis Stamelos, Elias Koromilas, Michaela Blott, and Kees Vissers. Elastic-DF: Scaling performance of DNN inference in FPGA clouds through automatic partitioning. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):15:1–15:34, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3470567.

Aggarwal:2012:SFT

[ASGY12] Vikas Aggarwal, Greg Stitt, Alan George, and Changil Yoon. SCF: a framework for task-level coordination in reconfigurable, heterogeneous systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(2):7:1–7:??, June 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Alachiotis:2022:SPR

[ASPP22] Nikolaos Alachiotis, Panagiotis Skrimponis, Manolis Pissadakis, and Dionisios Pnevmatikatos. Scalable phylogeny reconstruction with disaggregated near-memory processing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (3):25:1–25:32, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3484983.

Attia:2016:RAD

[ATJZ16] Osama G. Attia, Kevin R. Townsend, Phillip H. Jones, and Joseph Zambreno. A reconfigurable architecture for the detection of strongly connected components. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(2):16:1–16:??, February 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ananthan:2013:RPH

[AV13] T. Ananthan and M. V. Vaidyan. A reconfigurable parallel hardware implementation of the self-tuning regulator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6 (4):17:1–17:??, December 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Alachiotis:2020:RXF

[AVCP20] Nikolaos Alachiotis, Charalampos Vatsolakis, Grigorios Chrysos, and Dionisios Pnevmatikatos. RAiSD-X: a fast and accurate FPGA system for the detection of positive selection in thousands of genomes. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(1):2:1–2:30, February 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3364225.

Ahmed:2019:FAB

[AZM^{+}19] Ibrahim Ahmed, Shuze Zhao, James Meijers, Olivier Trescases, and Vaughn Betz. FRoC 2.0: Automatic BRAM and logic testing to enable dynamic voltage scaling for FPGA applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(4):20:1–20:??, October 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3354188.

Bai:2015:ATF

[BAG15] Yuhui Bai, Syed Zahid Ahmed, and Bertrand Granado. ARC 2014: Towards a fast FPGA implementation of a heap-based priority queue for image coding using a parallel index-aware tree. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1):8:1–8:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bakos:2018:ISS

[Bak18] Jason D. Bakos. Introduction to the special section on FCCM’16. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):1:1–1:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ben-Asher:2010:RMC

[BAMR10] Yosi Ben-Asher, Danny Meisler, and Nadav Rotem. Reducing memory constraints in modulo scheduling synthesis for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(3):15:1–15:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ben-Asher:2013:OWS

[BAMR13] Yosi Ben-Asher, Ron Meldiner, and Nadav Rotem. Optimizing wait states in the synthesis of memory references with unpredictable latencies. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(4):19:1–19:??, December 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Brennsteiner:2022:RTD

[BATM22] Stefan Brennsteiner, Tughrul Arslan, John Thompson, and Andrew McCormick. A real-time deep learning OFDM receiver. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):26:1–26:25, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3494049.

Banerjee:2010:BMA

[BBND10] Sudarshan Banerjee, Elaheh Bozorgzadeh, Juanjo Noguera, and Nikil Dutt. Bandwidth management in application mapping for dynamically reconfigurable architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(3):18:1– 18:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Brignone:2025:SAS

[BBO^{+}25] Giovanni Brignone, Roberto Bosio, Fabrizio Ottati, Claudio Sansoè, and Luciano Lavagno. SILVIA: Automated superwordlevel parallelism exploitation via HLS-specific LLVM passes for compute-intensive FPGA accelerators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):18:1–18:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Boland:2011:OMB

[BC11] David Boland and George A. Constantinides. Optimizing memory bandwidth use and performance for matrix-vector multiplication in iterative methods. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):22:1–22:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Badrignans:2010:SSA

[BCE^{+}10] Benoît Badrignans, David Champagne, Reouven Elbaz, Catherine Gebotys, and Lionel Torres. SARFUM: Security architecture for remote FPGA update and monitoring. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(2):8:1–8:??, May 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Beasley:2021:OCH

[BCW21] Alexander E. Beasley, C. T. Clarke, and R. J. Watson. An OpenGL compliant hardware implementation of a graphic processing unit using field programmable gate array-system on chip technology. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(1):2:1–2:24, January 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3410357.

Bhasin:2015:EFB

[BDGH15] Shivam Bhasin, Jean-Luc Danger, Sylvain Guilley, and Wei He. Exploiting FPGA block memories for protected cryptographic implementations. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(3):16:1–16:??, May 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bo:2019:APR

[BDX^{+}19] Chunkun Bo, Vinh Dang, Ted Xie, Jack Wadden, Mircea Stan, and Kevin Skadron. Automata processing in reconfigurable architectures: In-the-cloud deployment, cross-platform evaluation, and fast symbol-only reconfiguration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(2):9:1–9:??, June 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3314576.

Bobda:2019:ISS

[BE19] Chistophe Bobda and Ken Eguro. Introduction to the special section on security in FPGA-accelerated cloud and datacenters. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(3):11:1–11:??, September 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3352060.

Becker:2014:ITS

[Bec14] Tobias Becker. Introduction to the TRETS special section on the Workshop on Self-Awareness in Reconfigurable Computing Systems (SRCS’12). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):11:1–11:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bucknall:2023:ZEE

[BF23] Alex R. Bucknall and Suhaib A. Fahmy. ZyPR: End-to-end build tool and runtime manager for partial reconfiguration of FPGA SoCs at the edge. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):34:1–34:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3585521.

Besta:2020:SCM

[BFBN^{+}20] Maciej Besta, Marc Fischer, Tal Ben-Nun, Dimitri Stanojevic, Johannes De Fine Licht, and Torsten Hoefler. Substream-centric maximum matchings on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(2):8:1–8:33, June 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3377871.

Beeckler:2008:PGR

[BG08] John S. Beeckler and Warren J. Gross. Particle graphics on reconfigurable hardware. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(3):15:1–15:??, September 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Burovskiy:2017:EAH

[BGSL17] Pavel Burovskiy, Paul Grigoras, Spencer Sherwin, and Wayne Luk. Efficient assembly for high-order unstructured FEM meshes (FPL 2015). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(2):12:1–12:??, April 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Brugger:2014:RRF

[BHB14] Christian Brugger, Dominic Hillenbrand, and Matthias Balzer. RIVER: Reconfigurable flow and fabric for real-time signal processing on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):24:1–24:??, August 2014. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bao:2025:CCH

[BHF^{+}25] Tianyou Bao, Pengzhou He, Daisuke Fujimoto, Yuichi Hayashi, and Jiafeng Xie. CHIRP: Compact and high-performance FPGA implementation of unified hardware accelerators for ring-binary-LWE-based PQC. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):19:1–19:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Biedermann:2015:SDR

[BHI15] Alexander Biedermann, Sorin A. Huss, and Adeel Israr. Safe dynamic reshaping of reconfigurable MPSoC embedded systems for self-healing and self-adaption purposes. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(4):26:1–26:??, October 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bao:2024:AFI

[BHXJ24] Tianyou Bao, Pengzhou He, Jiafeng Xie, and H. S. Jacinto. AEKA: FPGA implementation of area-efficient Karatsuba accelerator for ring-binary-LWE-based lightweight PQC. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (2):29:1–29:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3637215.

Beckhoff:2014:DTI

[BKT14] Christian Beckhoff, Dirk Koch, and Jim Torresen. Design tools for implementing self-aware and fault-tolerant systems on FP-GAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):14:1–14:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Buell:2008:I

[BL08] Duncan Buell and Wayne Luk. Introduction. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(1): 1:1–1:??, March 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bobda:2022:FFA

[BMC^{+}22] Christophe Bobda, Joel Mandebi Mbongue, Paul Chow, Mohammad Ewais, Naif Tarafdar, Juan Camilo Vega, Ken Eguro, Dirk Koch, Suranga Handagala, Miriam Leeser, Martin Herbordt, Hafsah Shahzad, Peter Hofste, Burkhard Ringlein, Jakub Szefer, Ahmed Sanaullah, and Russell Tessier. The future of FPGA acceleration in datacenters and the cloud. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):34:1–34:42, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3506713.

Bourge:2016:GEC

[BMR16] Alban Bourge, Olivier Muller, and Frédéric Rousseau. Generating efficient context-switch capable circuits through autonomous design flow. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):9:1–9:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bodily:2010:CSI

[BNW^{+}10] John Bodily, Brent Nelson, Zhaoyi Wei, Dah-Jye Lee, and Jeff Chase. A comparison study on implementing optical flow and digital communications on FPGAs and GPUs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(2): 6:1–6:??, May 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bouganis:2009:SOF

[BPCC09] Christos-S. Bouganis, Sung-Boem Park, George A. Constantinides, and Peter Y. K. Cheung. Synthesis and optimization of 2D filter designs for heterogeneous FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4): 24:1–24:??, January 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Blott:2018:FRE

[BPF^{+}18] Michaela Blott, Thomas B. Preußer, Nicholas J. Fraser, Giulio Gambardella, Kenneth O’brien, Yaman Umuroglu, Miriam Leeser, and Kees Vissers. FINN-R: an end-to-end deep-learning framework for fast exploration of quantized neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(3):16:1–16:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Bergeron:2011:LTF

[BPFD11] Etienne Bergeron, Louis-David Perron, Marc Feeley, and Jean Pierre David. Logarithmic-time FPGA bitstream analysis: a step towards JIT hardware compilation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):12:1– 12:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Butler:2015:HSH

[BS15] Jon T. Butler and Tsutomu Sasao. High-speed hardware partition generation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):1:1–1:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Basalama:2023:FEE

[BSW^{+}23] Suhail Basalama, Atefeh Sohrabizadeh, Jie Wang, Licheng Guo, and Jason Cong. FlexCNN: an end-to-end framework for composing CNN accelerators on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):23:1–23:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3570928.

Boutros:2018:YCI

[BYB18] Andrew Boutros, Sadegh Yazdanshenas, and Vaughn Betz. You cannot improve what you do not measure: FPGA vs. ASIC efficiency gaps for convolutional neural network inference. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(3):20:1–20:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Abdelhamid:2023:SMC

[BYB23] Riadh Ben Abdelhamid, Yoshiki Yamaguchi, and Taisuke Boku. A scalable many-core overlay architecture on an HBM2-enabled multi-die FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):15:1–15:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3547657.

Chen:2011:EDL

[CA11] Xiaoheng Chen and Venkatesh Akella. Exploiting data-level parallelism for energy-efficient implementation of LDPC decoders and DCT on an FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):37:1–37:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Corts:2023:SPS

[CA23] Reinout Corts and Nikolaos Alachiotis. A survey of processing systems for phylogenetics and population genetics. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (3):35:1–35:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3588033.

Cook:2022:INU

[CAG^{+}22] Hayden Cook, Jacob Arscott, Brent George, Tanner Gaskin, Jeffrey Goeders, and Brad Hutchings. Inducing non-uniform FPGA aging using configuration-based short circuits. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (4):41:1–41:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3517042.

Cevrero:2009:FPC

[CAPA^{+}09] Alessandro Cevrero, Panagiotis Athanasopoulos, Hadi Parandeh-Afshar, Ajay K. Verma, Hosein Seyed Attarzadeh Niaki, Chrysostomos Nicopoulos, Frank K. Gurkaynak, Philip Brisk, Yusuf Leblebici, and Paolo Ienne. Field programmable compressor trees: Acceleration of multi-input addition on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2): 13:1–13:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Cancare:2012:EHC

[CBC^{+}12] Fabio Cancare, Davide B. Bartolini, Matteo Carminati, Donatella Sciuto, and Marco D. Santambrogio. On the evolution of hardware circuits via reconfigurable architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(4):22:1– 22:??, December 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Cheah:2014:IDB

[CBFM14] Hui Yan Cheah, Fredrik Brosser, Suhaib A. Fahmy, and Douglas L. Maskell. The iDEA DSP block-based soft processor for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):19:1–19:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Clemente:2014:MSA

[CBR^{+}14] Juan Antonio Clemente, Ivan Beretta, Vincenzo Rana, David Atienza, and Donatella Sciuto. A mapping-scheduling algorithm for hardware acceleration on reconfigurable platforms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):9:1–9:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Choi:2018:DAM

[CCF^{+}18] Young-Kyu Choi, Jason Cong, Zhenman Fang, Yuchen Hao, Glenn Reinman, and Peng Wei. In-depth analysis on microarchitectures of modern heterogeneous CPU–FPGA platforms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12 (1):4:1–4:??, April 2018. CODEN ???? ISSN 1936-7406 (print),

1936-7414 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3294054.

Chen:2022:TRE

[CCT^{+}22] Xinyu Chen, Feng Cheng, Hongshi Tan, Yao Chen, Bingsheng He, Weng-Fai Wong, and Deming Chen. ThunderGP: Resource-efficient graph processing framework on FPGAs with HLS. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):44:1–44:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3517141.

Chen:2024:HNF

[CCZ^{+}24] Sichao Chen, Chang Cai, Su Zheng, Jiangnan Li, Guowei Zhu, Jingyuan Li, Yazhou Yan, Yuan Dai, Wenbo Yin, and Lingli Wang. HierCGRA: a novel framework for large-scale CGRA with hierarchical modeling and automated design space exploration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):35:1–35:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3656176.

Cardoso:2015:GEF

[CDM15] João M. P. Cardoso, Pedro C. Diniz, and Katherine (Compton) Morrow. Guest editorial: FPL 2013. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):8:1–8:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Cooke:2015:TAF

[CFBS15] Patrick Cooke, Jeremy Fowers, Greg Brown, and Greg Stitt. A tradeoff analysis of FPGAs, GPUs, and multicores for slidingwindow applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(1):2:1–2:??, February 2015. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chiu:2010:MDS

[CH10] Matt Chiu and Martin C. Herbordt. Molecular dynamics simulations on high-performance reconfigurable computing systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):23:1–23:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Cheung:2011:ISS

[Che11] Peter Y. K. Cheung. Introduction to special section FPGA 2009. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):31:1–31:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chen:2016:I

[Che16] Deming Chen. Introduction. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):28:1–28:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chen:2019:EMN

[Che19] Deming Chen. Editorial: a message from the new Editor-in-Chief. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(2):6:1–6:??, June 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3326451.

Chen:2022:NTE

[Che22] Deming Chen. Note from the TRETS EiC about the new journalfirst track in FPT’21. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):7e:1, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3501280.

Cahill:2022:AFD

[CHG22] Eli Cahill, Brad Hutchings, and Jeffrey Goeders. Approaches for FPGA design assurance. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):28:1–28:29, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3491233.

Chen:2024:EVL

[CJH^{+}24] Jeffrey Chen, Sang-Woo Jun, Sehwan Hong, Warrick He, and Jinyeong Moon. Eciton: Very low-power recurrent neural network accelerator for real-time inference at the edge. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (1):16:1–16:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3629979.

Cheng:2023:PCF

[CJWC23] Jianyi Cheng, Lana Josipović, John Wickerson, and George A. Constantinides. Parallelising control flow in dynamic-scheduling high-level synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):55:1–55:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3599973.

Curreri:2010:PAF

[CKG^{+}10] John Curreri, Seth Koehler, Alan D. George, Brian Holland, and Rafael Garcia. Performance analysis framework for high-level language applications in reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(1): 5:1–5:??, January 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Carril:2024:HAH

[CKRG^{+}24] Xavier Carril, Charalampos Kardaris, Jordi Ribes-GonzáLez, Oriol Farràs, Carles Hernandez, Vatistas Kostalabros, Joel Ulises González-Jiménez, and Miquel Moretó. Hardware acceleration for high-volume operations of CRYSTALS-Kyber and CRYSTALS-Dilithium. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):41:1–41:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3675172.

Cong:2022:FHT

[CLL^{+}22] Jason Cong, Jason Lau, Gai Liu, Stephen Neuendorffer, Peichen Pan, Kees Vissers, and Zhiru Zhang. FPGA HLS today: Successes, challenges, and opportunities. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):51:1–51:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3530775.

Chen:2014:GMA

[CM14] Liang Chen and Tulika Mitra. Graph minor approach for application mapping on CGRAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):21:1–21:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Campos:2024:EEC

[CMT^{+}24] Javier Campos, Jovan Mitrevski, Nhan Tran, Zhen Dong, Amir Gholaminejad, Michael W. Mahoney, and Javier Duarte. Endto-end codesign of Hessian-aware quantized neural networks for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):36:1–36:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3662000.

Chau:2015:MAP

[CNE^{+}15] Thomas C. P. Chau, Xinyu Niu, Alison Eele, Jan Maciejowski, Peter Y. K. Cheung, and Wayne Luk. Mapping adaptive particle filters to heterogeneous reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4): 9:1–9:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chen:2024:CTO

[CNK^{+}24] Kuan-Yu Chen, Thomas Mason Nelson, Alireza Khadem, Morteza Fayazi, Sanjay Sri Vallabh Singapuram, Ronald Dreslinski, Nishil Talati, Hun-Seok Kim, and David Blaauw. Canalis: a throughputoptimized framework for real-time stream processing of wireless communication. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):61:1–61:??, December 2024. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3695880.

Cao:2018:FRA

[CNZ^{+}18] Shijie Cao, Lanshun Nie, Dechen Zhan, Wenqiang Wang, Ningyi Xu, Ramashis Das, Ming Wu, Lintao Zhang, and Derek Chiou. FlexSaaS: a reconfigurable accelerator for Web search selection. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(1):5:1–5:??, April 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3301409.

Cho:2021:PMC

[CPFM21] Shenghsun Cho, Mrunal Patel, Michael Ferdman, and Peter Milder. Practical model checking on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(2):8:1– 8:18, July 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3448272.

Chung:2009:PTS

[CPN^{+}09] Eric S. Chung, Michael K. Papamichael, Eriko Nurvitadhi, James C. Hoe, Ken Mai, and Babak Falsafi. ProtoFlex: Towards scalable, full-system multiprocessor simulations using FP-GAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2):15:1–15:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chen:2018:ISS

[CPW18] Deming Chen, Andrew Putnam, and Steve Wilton. Introduction to the special section on deep learning in FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11 (3):14:1–14:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Cardoso:2017:ISS

[CS17] João M. P. Cardoso and Cristina Silvano. Introduction to the special section on FPL 2015. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(2):10:1–10:??, April 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chu:2017:FCA

[CSK17] Thiem Van Chu, Shimpei Sato, and Kenji Kise. Fast and cycleaccurate emulation of large-scale networks-on-chip using a single FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(4):27:1–27:??, December 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Choi:2023:FAP

[CSS^{+}23] Young-Kyu Choi, Carlos Santillana, Yujia Shen, Adnan Darwiche, and Jason Cong. FPGA acceleration of probabilistic sentential decision diagrams with high-level synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (2):18:1–18:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3561514.

Chao:2016:DTM

[CTH16] Hung-Lin Chao, Sheng-Ya Tung, and Pao-Ann Hsiung. Dynamic task mapping with congestion speculation for reconfigurable network-on-chip. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):3:1–3:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chin:2009:SDM

[CW09] Scott Y. L. Chin and Steven J. E. Wilton. Static and dynamic memory footprint reduction for FPGA routing algorithms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4):18:1–18:??, January 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Compton:2009:ISI

[CWBD09] Katherine Compton, Roger Woods, Christos Bouganis, and Pedro Diniz. Introduction to the special issue ARC’08. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2 (4):20:1–20:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Cheng:2012:STP

[CXG^{+}12] Lerong Cheng, Wenyao Xu, Fang Gong, Yan Lin, Ho-Yan Wong, and Lei He. Statistical timing and power optimization of architecture and device for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(2):9:1–9:??, June 2012. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Cong:2009:FBH

[CZ09] Jason Cong and Yi Zou. FPGA-based hardware acceleration of lithographic aerial image simulation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(3):17:1–17:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chen:2025:UPF

[CZD^{+}25] Hongzheng Chen, Jiahao Zhang, Yixiao Du, Shaojie Xiang, Zichao Yue, Niansong Zhang, Yaohui Cai, and Zhiru Zhang. Understanding the potential of FPGA-based spatial acceleration for large language model inference. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):5:1–5:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3656177.

Chang:2023:AHC

[CZZ23] Liang Chang, Xin Zhao, and Jun Zhou. ADAS: a high computational utilization dynamic reconfigurable hardware accelerator for super resolution. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):47:1–47:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3570927.

Duarte:2015:ACK

[DB15] Rui Policarpo Duarte and Christos-Savvas Bouganis. ARC 2014 over-clocking KLT designs on FPGAs under process, voltage, and temperature variation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1):7:1–7:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

deBruin:2024:RBE

[dBVW^{+}24] Barry de Bruin, Kanishkan Vadivel, Mark Wijtvliet, Pekka Jääskeläinen, and Henk Corporaal. R-Blocks: an energy-efficient, flexible, and programmable CGRA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):34:1–34:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3656642.

Dehon:2016:ISI

[DC16] André Dehon and Derek Chiou. Introduction to special issue on reconfigurable components with source code. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(3):19:1– 19:??, July 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Du:2022:BAB

[DCL^{+}22] Gaoming Du, Bangyi Chen, Zhenmin Li, Zhenxing Tu, Junjie Zhou, Shenya Wang, Qinghao Zhao, Yongsheng Yin, and Xiaolei Wang. A BNN accelerator based on edge-skip-calculation strategy and consolidation compressed tree. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):30:1–30:20, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3494569.

DelSozzo:2024:ATS

[DCS24] Emanuele Del Sozzo, Davide Conficconi, and Kentaro Sano. Across time and space: Senju’s approach for scaling iterative stencil loop accelerators on single and multiple FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (2):28:1–28:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3634920.

Dhawan:2015:AEN

[DD15] Udit Dhawan and André Dehon. Area-efficient near-associative memories on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):3:1–3:??, January 2015. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Daigneault:2018:ASS

[DD18] Marc-Andre Daigneault and Jean Pierre David. Automated synthesis of streaming transfer level hardware designs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11 (2):13:1–13:??, November 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Dubois:2010:SMV

[DDB^{+}10] David Dubois, Andrew Dubois, Thomas Boorman, Carolyn Connor, and Steve Poole. Sparse matrix-vector multiplication on a reconfigurable supercomputer with application. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(1): 2:1–2:??, January 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

deDinechin:2013:FPE

[dDELVP13] Florent de Dinechin, Pedro Echeverría, Marisa López-Vallejo, and Bogdan Pasca. Floating-point exponentiation units for reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(1):4:1–4:??, May 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Demertzi:2011:DSO

[DDH^{+}11] Melina Demertzi, Pedro C. Diniz, Mary W. Hall, Anna C. Gilbert, and Yi Wang. Domain-specific optimization of signal recognition targeting FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):17:1–17:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Dogan:2022:CBB

[DE22] Atakan Dogan and Kemal Ebcioglu. Cloud building block chip for creating FPGA and ASIC clouds. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):14:1–14:35, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3466822.

Dehon:2020:ISS

[Deh20] André Dehon. Introduction to special section on FCCM 2019. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(4):17:1–17:2, October 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3410373.

Damiani:2022:BFS

[DFB^{+}22] Andrea Damiani, Giorgia Fiscaletti, Marco Bacis, Rolando Brondolin, and Marco D. Santambrogio. BlastFunction: a full-stack framework bringing FPGA hardware acceleration to cloud-native applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):17:1–17:27, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3472958.

Drimer:2010:DBP

[DGP10] Saar Drimer, Tim Güneysu, and Christof Paar. DSPs, BRAMs, and a pinch of logic: Extended recipes for AES on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(1):3:1–3:??, January 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

DiCarlo:2015:SSA

[DGP^{+}15] Stefano Di Carlo, Giulio Gambardella, Paolo Prinetto, Daniele Rolfo, and Pascal Trotta. SATTA: a Self-Adaptive Temperature-Based TDF Awareness methodology for dynamically reconfigurable FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(1):1:1–1:??, February 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

DeHon:2008:GET

[DH08] André DeHon and Mike Hutton. Guest editorial: TRETS special edition on the 15th International Symposium on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(1):2:1–2:??, March 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Davis:2018:KHA

[DHL^{+}18] James J. Davis, Eddie Hung, Joshua M. Levine, Edward A. Stott, Peter Y. K. Cheung, and George A. Constantinides. KAPow: High-accuracy, low-overhead online per-module power estimation for FPGA designs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):2:1–2:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Dutt:2009:TBD

[DL09] Shantanu Dutt and Li Li. Trust-based design and check of FPGA circuits using two-level randomized ECC structures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2 (1):6:1–6:??, March 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ding:2018:LLH

[DLBM18] Ruizhou Ding, Zeye Liu, R. D. (Shawn) Blanton, and Diana Marculescu. Lightening the load with highly accurate storage- and energy-efficient LightNNs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(3):17:1–17:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Delomier:2020:MBD

[DLCJ20] Yann Delomier, Bertrand Le Gal, Jerémie Crenne, and Christophe Jego. Model-based design of hardware SC polar decoders for FP-GAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(2):10:1–10:27, June 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3391431.

Du:2024:FFA

[DLZ^{+}24] Linfeng Du, Tingyuan Liang, Xiaofeng Zhou, Jinming Ge, Shangkun Li, Sharad Sinha, Jieru Zhao, Zhiyao Xie, and Wei Zhang. FADO: Floorplan-aware directive optimization based on synthesis and analytical models for high-level synthesis designs on multi-die FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):47:1–47:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3653458.

deMoura:2023:DCR

[dMdLC23] Rafael Fão de Moura, Joao Paulo Cardoso de Lima, and Luigi Carro. Data and computation reuse in CNNs using memristor TCAMs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):14:1–14:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3549536.

Dinh:2019:NFI

[DNL19] Van Luan Dinh, Xuan Truong Nguyen, and Hyuk-Jae Lee. A novel FPGA implementation of a time-to-digital converter supporting run-time estimation and compensation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(2):

10:1–10:??, June 2019. CODEN ???? ISSN 1936-7406 (print),

1936-7414 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3322482.

Dumpala:2019:LUE

[DPHT19] Naveen Kumar Dumpala, Shivukumar B. Patil, Daniel Holcomb, and Russell Tessier. Loop unrolling for energy efficiency in lowcost field-programmable gate arrays. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(4):26:1–26:??, January 2019. CODEN ???? ISSN 1936-7406 (print), 1936-

7414 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3289186.

Dann:2024:GSP

[DRF24] Jonas Dann, Daniel Ritter, and Holger Fröning. GraphScale: Scalable processing on FPGAs for HBM and large graphs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):22:1–22:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3616497.

Dewald:2022:ILP

[DRHM22] Florian Dewald, Johanna Rohde, Christian Hochberger, and Heiko Mantel. Improving loop parallelization by a combination of static and dynamic analyses in HLS. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):31:1–31:31, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3501801.

Dobai:2015:LLF

[DS15] Roland Dobai and Lukas Sekanina. Low-level flexible architecture with hybrid reconfiguration for evolvable hardware. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8 (3):20:1–20:??, May 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Dragomir:2009:OLU

[DSB09] Ozana Silvia Dragomir, Todor Stefanov, and Koen Bertels. Optimal loop unrolling and shifting for reconfigurable architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(4):25:1–25:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Das:2015:ETD

[DSK15] Anup Das, Amit Kumar Singh, and Akash Kumar. Execution trace-driven energy-reliability optimization for multimedia MP-SoCs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(3):18:1–18:??, May 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Drewes:2024:TTL

[DSW^{+}24] Colin Drewes, Tyler Sheaves, Olivia Weng, Keegan Ryan, Bill Hunter, Christopher McCarty, Ryan Kastner, and Dustin Richmond. Turn on, tune in, and listen up: Maximizing side-channel recovery in cross-platform time-to-digital converters. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (3):49:1–49:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3666092.

Davidson:2015:IDC

[DVH^{+}15] Tom Davidson, Elias Vansteenkiste, Karel Heyse, Karel Bruneel, and Dirk Stroobandt. Identification of dynamic circuit specialization opportunities in RTL code. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(1):4:1–4:??, February 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Das:2015:ASE

[DVK15] Anup Das, Shyamsundar Venkataraman, and Akash Kumar. Autonomous soft-error tolerance of FPGA configuration bits. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):12:1–12:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Das:2013:TDA

[DW13] Joydip Das and Steven J. E. Wilton. Towards development of an analytical model relating FPGA architecture parameters to routability. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2):10:1–10:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

DAlberto:2022:XID

[DWN^{+}22] Paolo D’Alberto, Victor Wu, Aaron Ng, Rahul Nimaiyar, Elliott Delaye, and Ashish Sirasao. xDNN: Inference for deep convolutional neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):18:1–18:29, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3473334.

Easwaran:2011:NLB

[EA11] Lakshmi Easwaran and Ali Akoglu. Net-length-based routabilitydriven power-aware clustering. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):38:1–38:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Elrabaa:2019:PPP

[EAAAA19] Muhammad E. S. Elrabaa, Mohamed A. Al-Asli, and Marwan H. Abu-Amara. A protection and pay-per-use licensing scheme for on-cloud FPGA circuit IPs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(3):13:1–13:??, September 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3329861.

El-Araby:2009:EPR

[EAGEG09] Esam El-Araby, Ivan Gonzalez, and Tarek El-Ghazawi. Exploiting partial runtime reconfiguration for high-performance reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4):21:1–21:??, January 2009. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Eldafrawy:2020:FLB

[EBYB20] Mohamed Eldafrawy, Andrew Boutros, Sadegh Yazdanshenas, and Vaughn Betz. FPGA logic block architectures for efficient deep learning inference. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(3):12:1–12:34, September 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3393668.

Eguro:2022:ISIa

[ENPR22a] Ken Eguro, Stephen Neuendorffer, Viktor Prasanna, and Hongbo Rong. Introduction to special issue on FPGAs in data centers. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):11:1–11:2, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3493607.

Eguro:2022:ISIb

[ENPR22b] Ken Eguro, Stephen Neuendorffer, Viktor Prasanna, and Hongbo Rong. Introduction to special issue on FPGAs in data centers, Part II. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):22:1–22:2, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3495231.

Encinas:2025:LIM

[ERO25] Juan Encinas, Alfonso Rodríguez, and Andrés Otero. Leveraging incremental machine learning for reconfigurable systems modeling under dynamic workloads. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):15:1–15:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3715154.

Ebcioglu:2022:HPM

[ES22] Kemal Ebcioglu and Ismail San. Highly parallel multi-FPGA system compilation from sequential C/C++ code in the AWS cloud. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):47:1–47:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3507698.

Eusse:2015:CNP

[EWL15] Juan Fernando Eusse, Christopher Williams, and Rainer Leupers. CoEx: a novel profiling-based algorithm/architecture coexploration for ASIP design. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(3):17:1–17:??, May 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Faraji:2022:ACC

[FAB22] S. Rasoul Faraji, Pierre Abillama, and Kia Bazargan. Approximate constant-coefficient multiplication using hybrid binaryunary computing for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):29:1–29:25, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3494570.

Fahmy:2024:ISS

[FB24] Suhaib A. Fahmy and Jason D. Bakos. Introduction to the special section on FPGA 2023. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):43:1–43:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3695841.

FaoDeMoura:2024:RNL

[FC24] Rafael Fão De Moura and Luigi Carro. Reprogrammable nonlinear circuits using ReRAM for NN accelerators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (1):7:1–7:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3617894.

Fan:2024:HDF

[FHL^{+}24] Zimeng Fan, Wei Hu, Fang Liu, Dian Xu, Hong Guo, Yanxiang He, and Min Peng. A hardware design framework for computer vision models based on reconfigurable devices. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):2:1–2:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3635157.

Feng:2008:DEI

[FK08] Wenyi Feng and Sinan Kaptanoglu. Designing efficient input interconnect blocks for LUT clusters using counting and entropy. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(1):6:1–6:??, March 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ferikoglou:2025:CCA

[FKM^{+}25] Aggelos Ferikoglou, Andreas Kakolyris, Dimosthenis Masouros, Dimitrios Soudris, and Sotirios Xydis. CollectiveHLS: a collaborative approach to high-level synthesis design optimization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):11:1–11:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3702005.

Fekete:2012:DDR

[FKS^{+}12] Sándor P. Fekete, Tom Kamphans, Nils Schweer, Christopher Tessars, Jan C. van der Veen, Josef Angermeier, Dirk Koch, and Jürgen Teich. Dynamic defragmentation of reconfigurable devices. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(2):8:1–8:??, June 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Fang:2016:OSV

[FL16] Xin Fang and Miriam Leeser. Open-source variable-precision floating-point library for major commercial FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(3): 1–17, July 2016. ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/2851507.

Fraser:2020:KNL

[FL20] Nicholas J. Fraser and Philip H. W. Leong. Kernel normalised least mean squares with delayed model adaptation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13 (2):7:1–7:30, June 2020. CODEN ???? ISSN 1936-7406 (print),

1936-7414 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3376924.

Fraser:2017:FIK

[FLM^{+}17] Nicholas J. Fraser, Junkyu Lee, Duncan J. M. Moss, Julian Faraone, Stephen Tridgell, Craig T. Jin, and Philip H. W. Leong. FPGA implementations of kernel normalised least mean squares processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(4):26:1–26:??, December 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ferreira:2015:RFP

[FRS^{+}15] Ricardo Ferreira, Luciana Rocha, André G. Santos, José A. M. Nacif, Stephan Wong, and Luigi Carro. A runtime FPGA placement and routing using low-complexity graph traversal. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):9:1–9:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Fabry:2017:ERA

[FT17] Pieter Fabry and David Thomas. Efficient reconfigurable architecture for pricing exotic options. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(4):29:1–29:??, December 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Feng:2025:HTT

[FYZ^{+}25] Yongkang Feng, Liang Yao, Hongli Zhou, Minjie Wu, Shuai Xiang, Wanting Sun, Xiumin Xu, and Yingchun Lu. Highthroughput TRNG design with novelty adjustable TDC based on STR. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):29:1–29:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Fiege:2023:BBS

[FZ23] Nicolai Fiege and Peter Zipf. BLOOP: Boolean satisfiability-based optimized loop pipelining. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):49:1–49:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3599972.

Fiege:2025:FCW

[FZ25] Nicolai Fiege and Peter Zipf. Fantastic circuits and where to find them — a holistic ILP formulation for model-based hardware design. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):12:1–12:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3705325.

Galuzzi:2011:ISE

[GB11] Carlo Galuzzi and Koen Bertels. The instruction-set extension problem: a survey. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):18:1–18:28, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Goehringer:2013:ISS

[GC13] Diana Goehringer and René Cumplido. Introduction to the special section on 19th Reconfigurable Architectures Workshop (RAW 2012). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2):6:1–6:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Guo:2023:TST

[GCL^{+}23] Licheng Guo, Yuze Chi, Jason Lau, Linghao Song, Xingyu Tian, Moazin Khatti, Weikang Qiao, Jie Wang, Ecenur Ustun, Zhenman Fang, Zhiru Zhang, and Jason Cong. TAPA: a scalable taskparallel dataflow programming framework for modern FPGAs with co-optimization of HLS and physical design. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (4):63:1–63:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3609335.

Glaser:2011:TFT

[GDHG11] Johann Glaser, Markus Damm, Jan Haase, and Christoph Grimm. TR-FSM: Transition-based reconfigurable finite state machine. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):23:1–23:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Guillet:2014:EUM

[GdLlG^{+}14] Sébastien Guillet, Florent de Lamotte, Nicolas le Griguer, Éric Rutten, Guy Gogniat, and Jean-Philippe Diguet. Extending UML/MARTE to support discrete controller synthesis, application to reconfigurable systems-on-chip modeling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3): 27:1–27:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Giechaskiel:2019:LWE

[GER19] Ilias Giechaskiel, Ken Eguro, and Kasper B. Rasmussen. Leakier wires: Exploiting FPGA long wires for covert- and side-channel attacks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(3):11:1–11:??, September 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3322483.

Gaspar:2012:SEF

[GFBF12] Lubos Gaspar, Viktor Fischer, Lilian Bossuet, and Robert Fouquet. Secure extension of FPGA general purpose processors for symmetric key cryptography with partial reconfiguration capabilities. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3):16:1–16:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gozzi:2025:SHL

[GFC^{+}25] Giovanni Gozzi, Michele Fiorito, Serena Curzel, Claudio Barone, Vito Giovanni Castellana, Marco Minutoli, Antonino Tumeo, and Fabrizio Ferrandi. SPARTA: High-level synthesis of parallel multithreaded accelerators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):9:1–9:??, March 2025. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3677035.

Gan:2015:SGA

[GFL^{+}15] Lin Gan, Haohuan Fu, Wayne Luk, Chao Yang, Wei Xue, Xiaomeng Huang, Youhui Zhang, and Guangwen Yang. Solving the global atmospheric equations through heterogeneous reconfigurable platforms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):11:1–11:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Giesen:2018:COS

[GGR^{+}18] Hans Giesen, Benjamin Gojman, Raphael Rubin, Ji Kim, and André Dehon. Continuous online self-monitoring introspection circuitry for timing repair by incremental partial-reconfiguration (COSMIC TRIP). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):3:1–3:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gu:2017:IRF

[GHO17] Chongyan Gu, Neil Hanley, and Máire O’neill. Improved reliability of FPGA-based PUF identification generator design. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10 (3):20:1–20:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gross:2022:ESF

[GHWS22] Mathieu Gross, Konrad Hohentanner, Stefan Wiehler, and Georg Sigl. Enhancing the security of FPGA-SoCs via the usage of ARM TrustZone and a hybrid-TPM. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):5:1–5:26, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3472959.

Gajjar:2024:RFR

[GKA^{+}24] Archit Gajjar, Priyank Kashyap, Aydin Aysu, Paul Franzon, Yongjin Choi, Chris Cheng, Giacomo Pedretti, and Jim Ignowski. RD-FAXID: Ransomware detection with FPGA-accelerated XG-Boost. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):56:1–56:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3688396.

Gebauer:2023:QMR

[GKGS23] Richard Gebauer, Nick Karcher, Mehmed Güler, and Oliver Sander. QiCells: a modular RFSoC-based approach to interface superconducting quantum bits. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):32:1–32:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3571820.

Gohringer:2024:IFS

[GKK24] Diana Göhringer, Georgios Keramidas, and Akash Kumar. Introduction to the FPL 2021 special section. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (1):13:1–13:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3635115.

Ganewattha:2023:HAR

[GKLLA23] Chanaka Ganewattha, Zaheer Khan, Janne Lehtomäki, and Matti Latva-Aho. Hardware-accelerated real-time drift-awareness for robust deep learning on wireless RF data. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):19:1–19:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3563394.

Gantel:2012:ERP

[GKM^{+}12] Laurent Gantel, Amel Khiar, Benoit Miramond, Mohamed El Amine Benkhelifa, Lounis Kessal, Fabrice Lemonnier, and Jimmy Le Rhun. Enhancing reconfigurable platforms programmability for synchronous data-flow applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3):14:1– 14:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Grant:2008:PMS

[GL08] David Grant and Guy Lemieux. Perturb + mutate: Semisynthetic circuit generation for incremental placement and routing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(3):16:1–16:??, September 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gerlein:2017:NCA

[GMBC17] Eduardo A. Gerlein, T. M. Mcginnity, Ammar Belatreche, and Sonya Coleman. Network on chip architecture for multi-agent systems in FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(4):25:1–25:??, December 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Guo:2023:RAP

[GMZ^{+}23] Licheng Guo, Pongstorn Maidee, Yun Zhou, Chris Lavin, Eddie Hung, Wuxi Li, Jason Lau, Weikang Qiao, Yuze Chi, Linghao Song, Yuanlong Xiao, Alireza Kaviani, Zhiru Zhang, and Jason Cong. RapidStream 2.0: Automated parallel implementation of latency-insensitive FPGA designs through partial reconfiguration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):59:1–59:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3593025.

Gojman:2015:GLG

[GNM^{+}15] Benjamin Gojman, Sirisha Nalmela, Nikil Mehta, Nicholas Howarth, and André Dehon. GROK-LAB: Generating real onchip knowledge for intra-cluster delays using timing extraction. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):5:1–5:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ganegedara:2013:CPA

[GP13] Thilan Ganegedara and Viktor Prasanna. A comprehensive performance analysis of virtual routers on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2): 9:1–9:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gribok:2024:CCP

[GPL24] Sergey Gribok, Bogdan Pasca, and Martin Langhammer. CSAIL2019 crypto-puzzle solver architecture. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (3):44:1–44:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3639056.

Guneysu:2008:SPH

[GPP08] Tim Güneysu, Christof Paar, and Jan Pelzl. Special-purpose hardware for solving the elliptic curve discrete logarithm problem. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(2):8:1–8:??, June 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gorjiara:2008:MDC

[GRG08] Bita Gorjiara, Mehrdad Reshadi, and Daniel Gajski. Merged dictionary code compression for FPGA implementation of custom microcoded PEs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(2):11:1–11:??, June 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gibson:2022:ACM

[GRNW22] Kahlan Gibson, Esther Roorda, Daniel Holanda Noronha, and Steven J. E. Wilton. Adaptive clock management of HLSgenerated circuits on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):49:1–49:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3520140.

Guo:2010:OSC

[GS10] Xu Guo and Patrick Schaumont. Optimized system-on-chip integration of a programmable ECC coprocessor. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1): 6:1–6:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Geethakumari:2023:SAC

[GS23] Prajith Ramakrishnan Geethakumari and Ioannis Sourdis. Stream aggregation with compressed sliding Windows. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (3):37:1–37:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3590774.

Goehringer:2015:GEA

[GSCB15] Diana Goehringer, Marco D. Santambrogio, João M. P. Cardoso, and Koen Bertels. Guest editorial: ARC 2014. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1): 5:1–5:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gharibian:2013:ASL

[GSJC13] Farnaz Gharibian, Lesley Shannon, Peter Jamieson, and Kevin Chung. Analyzing system-level information’s correlation to FPGA placement. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(3):15:1–15:??, October 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Giechaskiel:2023:CVC

[GTS23] Ilias Giechaskiel, Shanquan Tian, and Jakub Szefer. Cross-VM covert- and side-channel attacks in cloud FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (1):6:1–6:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3534972.

Garg:2020:HNC

[GWPK20] Tushar Garg, Saud Wasly, Rodolfo Pellizzoni, and Nachiket Kapre. HopliteBuf: Network calculus-based design of FPGA NoCs with provably stall-free FIFOs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(2):6:1–6:35, June 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3375899.

Gao:2023:RAR

[GWS23] Yizhao Gao, Song Wang, and Hayden Kwok-Hay So. A reconfigurable architecture for real-time event-based multi-object tracking. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):58:1–58:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3593587.

Gu:2021:DGB

[GWXW21] Zhenghua Gu, Wenqing Wan, Jundong Xie, and Chang Wu. Dependency graph-based high-level synthesis for maximum instruction parallelism. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(4):20:1–20:15, December 2021. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3468875.

Guo:2018:DSF

[GZY^{+}18] Kaiyuan Guo, Shulin Zeng, Jincheng Yu, Yu Wang, and Huazhong Yang. [DL] a survey of FPGA-based neural network inference accelerators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(1):2:1–2:??, April 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3289185.

Heyse:2015:IRL

[HBA^{+}15] Karel Heyse, Jente Basteleus, Brahim Al Farisi, Dirk Stroobandt, Oliver Kadlcek, and Oliver Pell. On the impact of replacing lowspeed configuration buses on FPGAs with the chip’s internal configuration infrastructure. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1):6:1–6:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

He:2023:FIC

[HBXA23] Pengzhou He, Tianyou Bao, Jiafeng Xie, and Moeness Amin. FPGA implementation of compact hardware accelerators for ringbinary-LWE-based post-quantum cryptography. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (3):45:1–45:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3569457.

Haase:2024:NLE

[HCGG24] Julian Haase, Najdet Charaf, Alexander Groß, and Diana Göhringer. NC-Library: Expanding SystemC capabilities for nested reConfigurable hardware modelling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):37:1– 37:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3662001.

Hormigo:2013:SRC

[HCOB13] Javier Hormigo, Gabriel Caffarena, Juan P. Oliver, and Eduardo Boemo. Self-reconfigurable constant multiplier for FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(3):14:1–14:??, October 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Honorat:2024:ABS

[HDMN24] Alexandre Honorat, Mickaël Dardaillon, Hugo Miomandre, and Jean-François Nezan. Automated buffer sizing of dataflow applications in a high-level synthesis workflow. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):8:1–8:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3626103.

Hossfeld:2024:HEC

[HDN^{+}24] Konstantin Hoßfeld, Hans Jakob Damsgaard, Jar Nurmi, Michaela Blott, and Thomas B. Preußer. High-efficiency compressor trees for latest AMD FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):30:1–30:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3645097.

Hoang:2014:IMD

[HF14] Anh-Tuan Hoang and Takeshi Fujino. Intra-masking dual-rail memory on LUT implementation for SCA-resistant AES on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):10:1–10:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Holland:2011:AMM

[HGLS11] Brian Holland, Alan D. George, Herman Lam, and Melissa C. Smith. An analytical model for multilevel performance prediction of multi-FPGA systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):27:1–27:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Huang:2013:VHS

[HH13] Chun-Hsian Huang and Pao-Ann Hsiung. Virtualizable hardware/ software design infrastructure for dynamically partially reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2):11:1–11:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hasler:2024:PAS

[HH24] Jennifer Hasler and Cong Hao. Programmable analog system benchmarks leading to efficient analog computation synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):12:1–12:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3625298.

Hsiung:2010:SPH

[HHSC10] Pao-Ann Hsiung, Chun-Hsian Huang, Jih-Sheng Shen, and Chen-Chi Chiang. Scheduling and placement of hardware/software real-time relocatable tasks in dynamically partially reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):9:1–9:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Huang:2015:ECO

[HLC^{+}15] Qijing Huang, Ruolong Lian, Andrew Canis, Jongsok Choi, Ryan Xi, Nazanin Calagar, Stephen Brown, and Jason Anderson. The effect of compiler optimizations on high-level synthesis-generated hardware. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(3):14:1–14:??, May 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hsiung:2008:PSB

[HLL08] Pao-Ann Hsiung, Chao-Sheng Lin, and Chih-Feng Liao. Perfecto: a SystemC-based design-space exploration framework for dynamically reconfigurable architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(3):17:1–17:??, September 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Huffmire:2010:SPR

[HLN^{+}10] Ted Huffmire, Timothy Levin, Thuy Nguyen, Cynthia Irvine, Brett Brotherton, Gang Wang, Timothy Sherwood, and Ryan Kastner. Security primitives for reconfigurable hardware-based systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(2):10:1–10:??, May 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hung:2021:AGF

[HLW^{+}21] José Romero Hung, Chao Li, Pengyu Wang, Chuanming Shao, Jinyang Guo, Jing Wang, and Guoyong Shi. ACE-GCN: a fast data-driven FPGA accelerator for GCN embedding. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14 (4):21:1–21:23, December 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3470536.

Holland:2009:RRA

[HNG09] Brian Holland, Karthik Nagarajan, and Alan D. George. RAT: RC amenability test for rapid performance prediction. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4): 22:1–22:??, January 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hogervorst:2022:HAH

[HNM^{+}22] Tom Hogervorst, Razvan Nane, Giacomo Marchiori, Tong Dong Qiu, Markus Blatt, and Alf Birger Rustad. Hardware acceleration of high-performance computational flow dynamics using high-bandwidth memory-enabled field-programmable gate arrays. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):20:1–20:35, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3476229.

Huang:2010:RCA

[HNS^{+}10] Miaoqing Huang, Vikram K. Narayana, Harald Simmler, Olivier Serres, and Tarek El-Ghazawi. Reconfiguration and communication-aware task scheduling for high-performance reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):20:1–20:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hutchings:2025:TFI

[HTG25] Daniel Hutchings, Adam Taylor, and Jeffrey Goeders. Toward FPGA intellectual property encryption from netlist to bitstream. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):1:1–1:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3656644.

Hemmert:2010:FEF

[HU10] K. Scott Hemmert and Keith D. Underwood. Fast, efficient floating-point adders and multipliers for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3 (3):11:1–11:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hubner:2012:ISI

[Hüb12] Michael Hübner. Introduction to the special issue on ReCoSoC 2011. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3):11:1–11:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hao:2025:PGH

[HZS^{+}25] Xiaochen Hao, Mingzhe Zhang, Ce Sun, Zhuofu Tao, Hongbo Rong, Yu Zhang, Lei He, Eric Petit, Wenguang Chen, and Yun Liang. Productively generating a high-performance linear algebra library on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):27:1–27:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Heisswolf:2013:VND

[HZW^{+}13] Jan Heisswolf, Aurang Zaib, Andreas Weichslgartner, Ralf König, Thomas Wild, Jürgen Teich, Andreas Herkersdorf, and Jürgen Becker. Virtual networks — distributed communication resource management. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2):8:1–8:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Istvan:2015:HTL

[IABV15] Zsolt István, Gustavo Alonso, Michaela Blott, and Kees Vissers. A hash table for line-rate data processing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):13:1– 13:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Iskandar:2023:NMC

[IAG23] Veronia Iskandar, Mohamed A. Abd El Ghany, and Diana Göhringer. Near-memory computing on FPGAs with 3Dstacked memories: Applications, architectures, and optimizations. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):16:1–16:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3547658.

Iturbe:2015:MAH

[IBH^{+}15] Xabier Iturbe, Khaled Benkrid, Chuan Hong, Ali Ebrahim, Raul Torrego, and Tughrul Arslan. Microkernel architecture and hardware abstraction layer of a reliable reconfigurable real-time operating system (R3TOS). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(1):5:1–5:??, February 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ienne:2023:ISS

[Ien23] Paolo Ienne. Introduction to the special section on FPGA 2022. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):56:1–56:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3618114.

Ioannou:2023:SOA

[IF23] Lenos Ioannou and Suhaib A. Fahmy. Streaming overlay architecture for lightweight LSTM computation on FPGA SoCs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):8:1–8:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3543069.

Ioannou:2020:UNA

[IGM^{+}20] Aggelos D. Ioannou, Konstantinos Georgopoulos, Pavlos Malakonakis, Dionisios N. Pnevmatikatos, Vassilis D. Papaefstathiou, Ioannis Papaefstathiou, and Iakovos Mavroidis. UNILOGIC: a novel architecture for highly parallel reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(4):21:1–21:32, October 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3409115.

Ifrim:2024:SRF

[ILP24] Rares Ifrim, Dumitrel Loghin, and Decebal Popescu. A systematic review of fast, scalable, and efficient hardware implementations of elliptic curve cryptography for blockchain. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (4):62:1–62:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3696422.

Itturiet:2014:APE

[INF^{+}14] Fábio Itturiet, Gabriel Nazar, Ronaldo Ferreira, Álvaro Moreira, and Luigi Carro. Adaptive parallelism exploitation under physical and real-time constraints for resilient systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):25:1– 25:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Iskander:2014:HLA

[IPC14] Yousef Iskander, Cameron Patterson, and Stephen Craven. Highlevel abstractions and modular debugging for FPGA design validation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(1):2:1–2:??, February 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Inoue:2011:TCD

[IYY^{+}11] Hiroaki Inoue, Junya Yamada, Hideyuki Yoneda, Katsumi Togawa, Masato Motomura, and Koichiro Furuta. Test compression for dynamically reconfigurable processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):40:1– 40:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Inoue:2010:VGL

[IZO^{+}10] Kazuki Inoue, Qian Zhao, Yasuhiro Okamoto, Hiroki Yosho, Motoki Amagasaki, Masahiro Iida, and Toshinori Sueyoshi. A variable-grain logic cell and routing architecture for a reconfigurable IP core. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):5:1–5:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jin:2015:MID

[JB15] Zheming Jin and Jason D. Bakos. Memory interface design for 3D stencil kernels on a massively parallel memory system. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(4):24:1–24:??, October 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jang:2009:WFT

[JCCM09] Stephen Jang, Billy Chan, Kevin Chung, and Alan Mishchenko. WireMap: FPGA technology mapping for improved routability and enhanced LUT merging. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2):14:1–14:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jacobs:2012:RFT

[JCG^{+}12] Adam Jacobs, Grzegorz Cieslewski, Alan D. George, Ann Gordon-Ross, and Herman Lam. Reconfigurable fault tolerance: a comprehensive framework for reliable and adaptive FPGA-based space computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(4):21:1–21:??, December 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jamal:2020:FTH

[JCGW20] Al-Shahna Jamal, Eli Cahill, Jeffrey Goeders, and Steven J. E. Wilton. Fast turnaround HLS debugging using dependency analysis and debug overlays. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(1):4:1–4:26, February 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3372490.

Jacob:2008:MBA

[JLB^{+}08] Arpith Jacob, Joseph Lancaster, Jeremy Buhler, Brandon Harris, and Roger D. Chamberlain. Mercury BLASTP: Accelerating protein sequence alignment. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(2):9:1–9:??, June 2008. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jin:2014:FAS

[JM14] Minxi Jin and Tsutomu Maruyama. Fast and accurate stereo vision system on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(1):3:1–3:??, February 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Josipovic:2023:RSD

[JMGI23] Lana Josipović, Axel Marmet, Andrea Guerrieri, and Paolo Ienne. Resource sharing in dataflow circuits. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):54:1–54:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3597614.

Jacobsen:2015:RRI

[JRHK15] Matthew Jacobsen, Dustin Richmond, Matthew Hogains, and Ryan Kastner. RIFFA 2.1: a reusable integration framework for FPGA accelerators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(4):22:1–22:??, October 2015. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jaiyeoba:2024:DAD

[JS24] Oluwole Jaiyeoba and Kevin Skadron. Dynamic-ACTS — a dynamic graph analytics accelerator for HBM-enabled FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):48:1–48:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3662002.

J:2014:MAN

[JSC14] Soumya J., Ashish Sharma, and Santanu Chattopadhyay. Multiapplication network-on-chip design using global mapping and local reconfiguration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):7:1–7:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Josipovic:2022:BPS

[JSG^{+}22] Lana Josipović, Shabnam Sheikhha, Andrea Guerrieri, Paolo Ienne, and Jordi Cortadella. Buffer placement and sizing for highperformance dataflow circuits. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):4:1–4:32, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3477053.

Jellum:2024:CRO

[JSLO24] Erling Rennemo Jellum, Martin Schoeberl, Edward Ashford Lee, and Milica Orlandic. Codesign of reactor-oriented hardware and software for cyber-physical systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):55:1–55:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3672083.

Jin:2009:ERA

[JTLC09] Qiwei Jin, David B. Thomas, Wayne Luk, and Benjamin Cope. Exploring reconfigurable architectures for tree-based option pricing models. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(4):21:1–21:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jun:2023:ASD

[JYJC23] Hyegang Jun, Hanchen Ye, Hyunmin Jeong, and Deming Chen. AutoScaleDSE: a scalable design space exploration engine for high-level synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):46:1–46:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3572959.

Jia:2024:XHP

[JZL^{+}24] Xijie Jia, Yu Zhang, Guangdong Liu, Xinlin Yang, Tianyu Zhang, Jia Zheng, Dongdong Xu, Zhuohuan Liu, Mengke Liu, Xiaoyang Yan, Hong Wang, Rongzhang Zheng, Li Wang, Dong Li, Satyaprakash Pareek, Jian Weng, Lu Tian, Dongliang Xie, Hong Luo, and Yi Shan. XVDPU: a high-performance CNN accelerator on the versal platform powered by the AI engine. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):20:1–20:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3617836.

Kim:2017:SSC

[KA17] Jin Hee Kim and Jason H. Anderson. Synthesizable standard cell FPGA fabrics targetable by the Verilog-to-routing CAD flow. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(2):11:1–11:??, April 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kara:2021:PGC

[KA21] Kaan Kara and Gustavo Alonso. PipeArch: Generic and contextswitch capable data processing on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(1):3:1–3:28, January 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3418465.

Kim:2014:FPF

[KAL14] Lok-Won Kim, Sameh Asaad, and Ralph Linsker. A fully pipelined FPGA architecture of a factored restricted Boltzmann machine artificial neural network. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(1):5:1–5:??, February 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kapre:2016:OSV

[Kap16] Nachiket Kapre. Optimizing soft vector processing in FPGA-based embedded systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(3):17:1–17:??, July 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Karakchi:2023:NND

[KB23] Rasha Karakchi and Jason D. Bakos. NAPOLY: a nondeterministic automata processor OverLaY. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (3):38:1–38:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3593586.

Keller:2009:ECC

[KBM09] Maurice Keller, Andrew Byrne, and William P. Marnane. Elliptic curve cryptography on FPGA for low-power applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(1):2:1–2:??, March 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Koch:2009:HDT

[KBT09] Dirk Koch, Christian Beckhoff, and Jürgen Teich. Hardware decompression techniques for FPGA-based embedded systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2):9:1–9:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kim:2014:USU

[KCC^{+}14] Changmoo Kim, Mookyoung Chung, Yeongon Cho, Mario Konijnenburg, Soojung Ryu, and Jeongwook Kim. ULP-SRP: Ultra low-power Samsung reconfigurable processor for biomedical applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):22:1–22:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Koh:2010:CMP

[KD10] Shannon Koh and Oliver Diessel. Configuration merging in point-to-point networks for module-based FPGA reconfiguration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(1):4:1–4:??, January 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kroh:2019:EFG

[KD19] Alexander Kroh and Oliver Diessel. Efficient fine-grained processor-logic interactions on the cache-coherent Zynq platform. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(4):25:1–25:??, January 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3277506.

Kyparissas:2021:LSC

[KD21] Nikolaos Kyparissas and Apostolos Dollas. Large-scale cellular automata on FPGAs: a new generic architecture and a framework. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(1):5:1–5:32, January 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3423185.

Kapre:2017:HDR

[KG17] Nachiket Kapre and Jan Gray. Hoplite: a deflection-routed directional torus NoC for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(2):14:1–14:??, April 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Keilbart:2024:DIC

[KGC^{+}24] Chris Keilbart, Yuhui Gao, Martin Chua, Eric Matthews, Steven J. E. Wilton, and Lesley Shannon. Designing an IEEE-compliant FPU that supports configurable precision for soft processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):33:1–33:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3650036.

Krieg:2012:PMP

[KGS^{+}12] Armin Krieg, Johannes Grinschgl, Christian Steger, Reinhold Weiss, Holger Bock, and Josef Haid. POWER-MODES: POWer-EmulatoR- and MOdel-Based DEpendability and Security Evaluations. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(4):19:1–19:??, December 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kirchgessner:2015:LOF

[KGS15] Robert Kirchgessner, Alan D. George, and Greg Stitt. Lowoverhead FPGA middleware for application portability and productivity. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(4):21:1–21:??, October 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Krautter:2019:MEL

[KGT19] Jonas Krautter, Dennis R. E. Gnad, and Mehdi B. Tahoori. Mitigating electrical-level attacks towards secure multi-tenant FPGAs in the cloud. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(3):12:1–12:??, September 2019. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3328222.

Khan:2024:OSD

[KHK24] Babar Khan, Carsten Heinz, and Andreas Koch. The open-source DeLiBA2 hardware/software framework for distributed storage accelerators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):23:1–23:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3624482.

Kaganov:2011:FAM

[KLC11] Alexander Kaganov, Asif Lakhany, and Paul Chow. FPGA acceleration of MultiFactor CDO pricing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):20:1–20:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kadric:2016:IPM

[KLD16] Edin Kadric, David Lakata, and André Dehon. Impact of parallelism and memory architecture on FPGA communication energy. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):30:1–30:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kanazawa:2010:ASL

[KM10] Kenji Kanazawa and Tsutomu Maruyama. An approach for solving large SAT problems on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):10:1–10:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kepa:2010:DAS

[KMK^{+}10] K. Kepa, F. Morgan, K. Kościuszkiewicz, L. Braun, M. Hübner, and J. Becker. Design assurance strategy and toolset for partially reconfigurable FPGA systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):4:1–4:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kornaros:2014:DPT

[KP14] George Kornaros and Dionisios Pnevmatikatos. Dynamic power and thermal management of NoC-Based heterogeneous MPSoCs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(1):1:1–1:??, February 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kourfali:2020:CDD

[KS20] Alexandra Kourfali and Dirk Stroobandt. In-circuit debugging with dynamic reconfiguration of FPGA interconnects.

ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(1):5:1–5:29, February 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3375459.

Koch:2024:ISI

[KS24] Andreas Koch and Kentaro Sano. Introduction to the special issue on FPL 2022. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):19:1–19:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3643474.

Kahoul:2010:EHA

[KSCC10] Asma Kahoul, Alastair M. Smith, George A. Constantinides, and Peter Y. K. Cheung. Efficient heterogeneous architecture floorplan optimization using analytical methods. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):3:1–3:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Koehler:2011:PAB

[KSG11] Seth Koehler, Greg Stitt, and Alan D. George. Platform-aware bottleneck detection for reconfigurable computing applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):30:1–30:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Khatti:2024:PPA

[KTB^{+}24] Moazin Khatti, Xingyu Tian, Ahmad Sedigh Baroughi, Akhil Raj Baranwal, Yuze Chi, Licheng Guo, Jason Cong, and Zhenman Fang. PASTA: Programming and automation support for scalable task-parallel HLS programs on modern multi-die FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):42:1–42:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3676849.

Kennings:2011:FTM

[KVK^{+}11] Andrew Kennings, Kristofer Vorwerk, Arun Kundu, Val Pevzner, and Andy Fox. FPGA technology mapping with encoded libraries and staged priority cuts. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):35:1–35:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Kalomiros:2024:HAS

[KVV24] John Kalomiros, John Vourvoulakis, and Stavros Vologiannidis. A hardware accelerator for the semi-global matching stereo algorithm: an efficient implementation for the Stratix V and Zynq UltraScale+ FPGA technology. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):5:1–5:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3615869.

Keller:2022:ITR

[KW22] Andrew M. Keller and Michael J. Wirthlin. The impact of terrestrial radiation on FPGAs in data centers. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (2):12:1–12:21, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3457198.

Khan:2018:EAM

[KY18] Farheen Fatima Khan and Andy Ye. An evaluation on the accuracy of the minimum-width transistor area models in ranking the layout area of FPGA architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):8:1–8:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Koch:2023:ISI

[KZ23] Andreas Koch and Wei Zhang. Introduction to the special issue on FPT 2021. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):30:1–30:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3603701.

Kalantar:2023:FBA

[KZB23] Amin Kalantar, Zachary Zimmerman, and Philip Brisk. FPGAbased acceleration of time series similarity prediction: From cloud to edge. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):12:1–12:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3555810.

Laforest:2017:MCM

[LA17] Charles Eric Laforest and Jason H. Anderson. Microarchitectural comparison of the MXP and Octavo soft-processor FPGA overlays. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3):19:1–19:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Leong:2017:FYF

[LAA^{+}17] Philip H. W. Leong, Hideharu Amano, Jason Anderson, Koen Bertels, João M. P. Cardoso, Oliver Diessel, Guy Gogniat, Mike Hutton, Junkyu Lee, Wayne Luk, Patrick Lysaght, Marco Platzner, Viktor K. Prasanna, Tero Rissa, Cristina Silvano, Hayden Kwok-Hay So, and Yu Wang. The first 25 years of the FPL conference: Significant papers. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(2):15:1–15:??, April 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Lu:2024:SAA

[LAF24] Alec Lu, Jahanvi Narendra Agrawal, and Zhenman Fang. SQL2FPGA: Automated acceleration of SQL query processing on modern CPU-FPGA platforms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):39:1–39:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3674843.

Leow:2013:AME

[LAL13] Yoon Kah Leow, Ali Akoglu, and Susan Lysecky. An analytical model for evaluating static power of homogeneous FPGA architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(4):18:1–18:??, December 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Lou:2023:FCF

[LBL23] Binglei Lou, David Boland, and Philip Leong. fSEAD: a composable FPGA-based streaming ensemble anomaly detection library. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):42:1–42:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3568992.

LeGal:2016:FSM

[LBRS16] Bertrand Le Gal, Yérom-David Bromberg, Laurent Réveillère, and Jigar Solanki. A flexible SoC and its methodology for parser-based applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):4:1–4:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Liu:2024:FBS

[LCL^{+}24] Yajing Liu, Ruiqi Chen, Shuyang Li, Jing Yang, Shun Li, and Bruno da Silva. FPGA-based sparse matrix multiplication accelerators: From state-of-the-art to future opportunities. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (4):59:1–59:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3687480.

Lam:2014:EFA

[LCS14] Siew-Kei Lam, Christopher T. Clarke, and Thambipillai Srikanthan. Exploiting FPGA-aware merging of custom instructions for runtime reconfiguration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):26:1–26:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Liu:2017:TOF

[LDJ^{+}17] Zhiqiang Liu, Yong Dou, Jingfei Jiang, Jinwei Xu, Shijie Li, Yongmei Zhou, and Yingnan Xu. Throughput-optimized FPGA accelerator for deep convolutional neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3):17:1– 17:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Leeser:2023:AEA

[Lee23] Miriam Leeser. Artifact evaluation for ACM TRETS papers submitted from the FPT journal track. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):40:1–40:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3596513.

Leong:2022:ISS

[Leo22] Philip H. W. Leong. Introduction to special section on FPGA 2021. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):42:1–42:??, December 2022. CODEN ????

ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3536335.

Liu:2018:OCB

[LFN^{+}18] Shuanglong Liu, Hongxiang Fan, Xinyu Niu, Ho cheung Ng, Yang Chu, and Wayne Luk. Optimizing CNN-based segmentation with deeply customized convolutional and deconvolutional architectures on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(3):19:1–19:??, December 2018. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Lu:2022:DSH

[LFS22] Alec Lu, Zhenman Fang, and Lesley Shannon. Demystifying the soft and hardened memory systems of modern FPGAs for software programmers through microbenchmarking. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (4):43:1–43:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3517131.

Lei:2014:FIS

[LGD^{+}14] Yuanwu Lei, Lei Guo, Yong Dou, Sheng Ma, and Jinbo Xu. FPGA implementation of a special-purpose VLIW structure for doubleprecision elementary function. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):8:1–8:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Luu:2014:VNG

[LGW^{+}14] Jason Luu, Jeffrey Goeders, Michael Wainberg, Andrew Somerville, Thien Yu, Konstantin Nasartschuk, Miad Nasr, Sen Wang, Tim Liu, Nooruddin Ahmed, Kenneth B. Kent, Jason Anderson, Jonathan Rose, and Vaughn Betz. VTR 7.0: Next generation architecture and CAD system for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):6:1–6:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Li:2023:ISS

[LH23] Jing Li and Martin Herbordt. Introduction to the special section on FCCM 2022. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):51:1–51:??, December 2023. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3632092.

Labrecque:2011:ASS

[LJS11] Martin Labrecque, Mark C. Jeffrey, and J. Gregory Steffan. Application-specific signatures for transactional memory in soft processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):21:1–21:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Luu:2011:VFC

[LKJ^{+}11] Jason Luu, Ian Kuon, Peter Jamieson, Ted Campbell, Andy Ye, Wei Mark Fang, Kenneth Kent, and Jonathan Rose. VPR 5.0: FPGA CAD and architecture exploration tools with single-driver routing, heterogeneity and process scaling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):32:1– 32:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Lusala:2012:STB

[LL12] Angelo Kuti Lusala and Jean-Didier Legat. A SDM–TDM-based circuit-switched router for on-chip networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3):15:1– 15:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Li:2023:FPF

[LLL^{+}23] Carol Jingyi Li, Xiangwei Li, Binglei Lou, Craig T. Jin, David Boland, and Philip H. W. Leong. Fixed-point FPGA implementation of the FFT accumulation method for real-time cyclostationary analysis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):41:1–41:??, September 2023. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3567429.

Laforest:2014:CMP

[LLO^{+}14] Charles Eric Laforest, Zimo Li, Tristan O’rourke, Ming G. Liu, and J. Gregory Steffan. Composing multi-ported memories on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):16:1–16:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Liu:2023:CKC

[LLS^{+}23] Kenneth Liu, Alec Lu, Kartik Samtani, Zhenman Fang, and Licheng Guo. CHIP-KNNv2: a configurable and highperformance K-nearest neighbors accelerator on HBM-based FP-GAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):62:1–62:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3616873.

Li:2024:ADC

[LLS^{+}24] Yonggen Li, Xin Li, Haibin Shen, Jicong Fan, Yanfeng Xu, and Kejie Huang. An all-digital compute-in-memory FPGA architecture for deep learning acceleration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):18:1–18:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3640469.

Liu:2024:EFB

[LLYC24] Zhengyan Liu, Qiang Liu, Shun Yan, and Ray C. C. Cheung. An efficient FPGA-based depthwise separable convolutional neural network accelerator with hardware pruning. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (1):15:1–15:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3615661.

Liu:2024:FAL

[LLZ^{+}24] Chaoqiang Liu, Xiaofei Liao, Long Zheng, Yu Huang, Haifeng Liu, Yi Zhang, Haiheng He, Haoyan Huang, Jingyi Zhou, and Hai Jin. L-FNNG: Accelerating large-scale KNN graph construction on CPU–FPGA heterogeneous platform. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):46:1–46:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3652609.

Leggett:2025:AMD

[LMD^{+}25] Jordan Leggett, John McGlone, Suleyman Demirsoy, Christian Faerber, and Vadim Pelyushenko. Accelerating in-memory database functionality with FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):13:1–13:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3706113.

La:2020:FMS

[LMG^{+}20] Tuan Minh La, Kaspar Matas, Nikola Grunchevski, Khoa Dang Pham, and Dirk Koch. FPGADefender: Malicious self-oscillator scanning for Xilinx UltraScale + FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(3):15:1– 15:31, September 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3402937.

Li:2023:SSA

[LML^{+}23] Xiangwei Li, Douglas L. Maskell, Carol Jingyi Li, Philip H. W. Leong, and David Boland. A scalable systolic accelerator for estimation of the spectral correlation density function and its FPGA implementation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):9:1–9:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3546181.

Leipnitz:2023:CAM

[LN23] Marcos T. Leipnitz and Gabriel L. Nazar. Constraint-aware multi-technique approximate high-level synthesis for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):61:1–61:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3624481.

Langhammer:2022:SNA

[LNGP22] Martin Langhammer, Eriko Nurvitadhi, Sergey Gribok, and Bogdan Pasca. Stratix 10 NX architecture. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):45:1–45:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3520197.

Lu:2010:ERD

[LOM10] Yingxi Lu, Maire O’Neill, and John McCanny. Evaluation of random delay insertion against DPA on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):11:1– 11:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Llamocca:2015:DEP

[LP15] Daniel Llamocca and Marios Pattichis. Dynamic energy, performance, and accuracy optimization and management using automatically generated constraints for separable 2D FIR filtering for digital video processing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):4:1–4:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Lienen:2022:DDR

[LP22] Christian Lienen and Marco Platzner. Design of distributed reconfigurable robotics systems with ReconROS. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3): 27:1–27:20, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3494571.

Liu:2013:INL

[LRA13] Hanyu Liu, Senthilkumar T. Rajavel, and Ali Akoglu. Integration of net-length factor with timing- and routability-driven clustering algorithms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(3):12:1–12:??, October 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Li:2023:JBA

[LSP^{+}23] Xiang Li, Peter Stanwicks, George Provelengios, Russell Tessier, and Daniel Holcomb. Jitter-based adaptive true random number generation circuits for FPGAs in the cloud. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):3:1–3:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3487554.

Lo:2009:SOC

[LT09] Chia-Tien Dan Lo and Yi-Gang Tai. Space optimization on counters for FPGA-based Perl compatible regular expressions. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(4):23:1–23:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Lai:2021:PSS

[LUX^{+}21] Yi-Hsiang Lai, Ecenur Ustun, Shaojie Xiang, Zhenman Fang, Hongbo Rong, and Zhiru Zhang. Programming and synthesis for software-defined FPGA acceleration: Status and future prospects. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(4):17:1–17:39, December 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3469660.

Lopez-Valdivieso:2024:DIH

[LVC24] Jonathan Lopez-Valdivieso and Rene Cumplido. Design and implementation of hardware–software architecture based on hashes for SPHINCS+. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):54:1–54:??, December 2024. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3653459.

Lamoureux:2008:TBP

[LW08] Julien Lamoureux and Steven J. E. Wilton. On the trade-off between power and flexibility of FPGA clock networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1 (3):13:1–13:??, September 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Li:2024:SHP

[LXS^{+}24] Kexin Li, Shaoxian Xu, Zhiyuan Shao, Ran Zheng, Xiaofei Liao, and Hai Jin. ScalaBFS2: a high-performance BFS accelerator on an HBM-enhanced FPGA chip. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):32:1–32:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3650037.

Li:2024:PPG

[LYC^{+}24] Guangyan Li, Zewen Ye, Donglong Chen, Wangchen Dai, Gaoyu Mao, Kejie Huang, and Ray C. C. Cheung. ProgramGalois: a programmable generator of radix-4 discrete Galois transformation architecture for lattice-based cryptography. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (4):53:1–53:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3689437.

Lu:2023:HET

[LYH^{+}23] Yingchun Lu, Yun Yang, Rong Hu, Huaguo Liang, Maoxiang Yi, Huang Zhengfeng, Yuanming Ma, Tian Chen, and Liang Yao. High-efficiency TRNG design based on multi-bit dual-ring oscillator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):64:1–64:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3624991.

Lu:2008:DCR

[LYS^{+}08] Shih-Lien L. Lu, Peter Yiannacouras, Taeweon Suh, Rolf Kassa, and Michael Konow. A desktop computer with a reconfigurable Pentium(R). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(1):5:1–5:??, March 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Li:2018:EMP

[LYZ^{+}18] Wensong Li, Fan Yang, Hengliang Zhu, Xuan Zeng, and Dian Zhou. An efficient memory partitioning approach for multi-pattern data access via data reuse. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(1):1:1–1:??, April 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3301296.

Liu:2019:PFF

[LZ19] Gai Liu and Zhiru Zhang. PIMap: a flexible framework for improving LUT-based technology mapping via parallelized iterative optimization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(4):23:1–23:??, January 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3268344.

Lanuzza:2010:ESR

[LZF^{+}10] M. Lanuzza, P. Zicari, F. Frustaci, S. Perri, and P. Corsonello. Exploiting self-reconfiguration capability to improve SRAM-based FPGA robustness in space and avionics applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1): 8:1–8:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Li:2024:FRT

[LZZ^{+}24] Yuqi Li, Kehao Zhao, Jieru Zhao, Qirui Wang, Shuda Zhong, Nageswara Lalam, Ruishu Wright, Peipei Zhou, and Kevin P. Chen. FiberFlex: Real-time FPGA-based intelligent and distributed fiber sensor system for pedestrian recognition. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):57:1–57:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3690389.

Maschi:2024:SHS

[MA24] Fabio Maschi and Gustavo Alonso. Strega: an HTTP server for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):3:1–3:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3611312.

Michail:2012:EHT

[MAK^{+}12] Harris E. Michail, George S. Athanasiou, Vasilis Kelefouras, George Theodoridis, and Costas E. Goutis. On the exploitation of a high-throughput SHA-256 FPGA design for HMAC. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(1):2:1–2:??, March 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Mishchenko:2011:SDC

[MBJJ11] Alan Mishchenko, Robert Brayton, Jie-Hong R. Jiang, and Stephen Jang. Scalable don’t-care-based logic optimization and resynthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):34:1–34:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Moscola:2010:HAR

[MCC10] James Moscola, Ron K. Cytron, and Young H. Cho. Hardwareaccelerated RNA secondary-structure alignment. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3 (3):14:1–14:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Meloni:2018:NEC

[MCD^{+}18] Paolo Meloni, Alessandro Capotondi, Gianfranco Deriu, Michele Brian, Francesco Conti, Davide Rossi, Luigi Raffo, and Luca Benini. NEURAghe: Exploiting CPU–FPGA synergies for efficient and flexible CNN inference acceleration on Zynq SoCs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(3):18:1–18:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Mehta:2013:UGE

[MCL^{+}13] Gayatri Mehta, Carson Crawford, Xiaozhong Luo, Natalie Parde, Krunalkumar Patel, Brandon Rodgers, Anil Kumar Sistla, Anil Yadav, and Marc Reisner. UNTANGLED: a game environment for discovery of creative mapping strategies. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(3):13:1– 13:??, October 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Mao:2023:HPC

[MCL^{+}23] Gaoyu Mao, Donglong Chen, Guangyan Li, Wangchen Dai, Abdurrashid Ibrahim Sanka, Çetin Kaya Koç, and Ray C. C. Cheung. High-performance and configurable SW/HW co-design of post-quantum signature CRYSTALS-Dilithium. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (3):44:1–44:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3569456.

Morgan:2012:RFL

[MCN12] Fearghal Morgan, Seamus Cawley, and David Newell. Remote FPGA lab for enhancing learning of digital systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3): 18:1–18:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Moini:2023:VSI

[MDL^{+}23] Shayan Moini, Aleksa Deric, Xiang Li, George Provelengios, Wayne Burleson, Russell Tessier, and Daniel Holcomb. Voltage sensor implementations for remote power attacks on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):11:1–11:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3555048.

Machado:2023:NNH

[MFOM23] Pedro Machado, João Filipe Ferreira, Andreas Oikonomou, and T. M. McGinnity. NeuroHSMD: Neuromorphic hybrid spiking motion detector. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):36:1–36:??, September 2023. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3588318.

Mahram:2015:NBH

[MH15] Atabak Mahram and Martin C. Herbordt. NCBI BLASTP on high-performance reconfigurable computing systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):

6:1–6:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Matsumoto:2008:SID

[MHK^{+}08] Yohei Matsumoto, Masakazu Hioki, Takashi Kawanami, Hanpei Koike, Toshiyuki Tsutsumi, Tadashi Nakagawa, and Toshihiro Sekigawa. Suppression of intrinsic delay variation in FPGAs using multiple configurations. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(1):3:1–3:??, March 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Murtaza:2009:CBB

[MHS09] S. Murtaza, A. G. Hoekstra, and P. M. A. Sloot. Compute bound and I/O bound cellular automata simulations on FPGA logic. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4):23:1–23:??, January 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Morcel:2019:FAC

[MHS^{+}19] Raghid Morcel, Hazem Hajj, Mazen A. R. Saghir, Haitham Akkary, Hassan Artail, Rahul Khanna, and Anil Keshavamurthy. FeatherNet: an accelerated convolutional neural network design for resource-constrained FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(2):6:1–6:??, June 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3306202.

Ma:2021:SFP

[MHT^{+}21] Rui Ma, Jia-Ching Hsu, Tian Tan, Eriko Nurvitadhi, David Sheffield, Rob Pelt, Martin Langhammer, Jaewoong Sim, Aravind Dasu, and Derek Chiou. Specializing FGPU for persistent deep learning. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(2):10:1–10:23, July 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3457886.

Majzoobi:2009:TDI

[MKP09] Mehrdad Majzoobi, Farinaz Koushanfar, and Miodrag Potkonjak. Techniques for design and implementation of secure reconfigurable PUFs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(1):5:1–5:??, March 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Meyer:2023:MFD

[MKP23] Marius Meyer, Tobias Kenter, and Christian Plessl. Multi-FPGA designs and scaling of HPC challenge benchmarks via MPI and circuit-switched inter-FPGA networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):24:1–24:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3576200.

Mbongue:2022:DMT

[MKSB22] Joel Mandebi Mbongue, Danielle Tchuinkou Kwadjo, Alex Shuping, and Christophe Bobda. Deploying multi-tenant FPGAs within Linux-based cloud infrastructure. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):19:1– 19:31, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-

7414 (electronic). URL https://dl.acm.org/doi/10.1145/
3474058.

Matthews:2022:QDR

[MLFS22] Eric Matthews, Alec Lu, Zhenman Fang, and Lesley Shannon. Quick-Div: Rethinking integer divider design for FPGA-based soft-processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):32:1–32:27, September 2022. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3502492.

Malik:2022:HEA

[MLPK22] Gurshaant Malik, Ian Elmore Lang, Rodolfo Pellizzoni, and Nachiket Kapre. HopliteML: Evolving application customized FPGA NoCs with adaptable routers and regulators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (4):40:1–40:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3507699.

McEvoy:2009:IWH

[MMMT09] Robert P. McEvoy, Colin C. Murphy, William P. Marnane, and Michael Tunstall. Isolated WDDL: a hiding countermeasure for differential power analysis on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(1):3:1–3:??, March 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Meeuws:2013:QSM

[MOG^{+}13] Roel Meeuws, S. Arash Ostadzadeh, Carlo Galuzzi, Vlad Mihai Sima, Razvan Nane, and Koen Bertels. Quipu: a statistical model for predicting hardware resources. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(1):3:1–3:??, May 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Menzel:2022:SSA

[MPK22] Johannes Menzel, Christian Plessl, and Tobias Kenter. The strong scaling advantage of FPGAs in HPC for N-body simulations. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):10:1–10:30, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3491235.

Murray:2020:VHP

[MPZ^{+}20] Kevin E. Murray, Oleg Petelin, Sheng Zhong, Jia Min Wang, Mohamed Eldafrawy, Jean-Philippe Legault, Eugene Sha, Aaron G. Graham, Jean Wu, Matthew J. P. Walker, Hanqing Zeng, Panagiotis Patros, Jason Luu, Kenneth B. Kent, and Vaughn Betz. VTR 8: High-performance CAD and customizable FPGA architecture modelling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(2):9:1–9:55, June 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3388617.

M:2023:DAR

[MS23] Dhayalakumar M. and Noor Mahammad Sk. Deterministic approach for range-enhanced reconfigurable packet classification engine. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):29:1–29:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3586577.

Mohtavipour:2025:PPR

[MS25] Seyed Mehdi Mohtavipour and Hadi Shahriar Shahhoseini. PRISA: a potential region-based intelligent search algorithm for dataflow graph mapping in spatial CGRAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):28:1– 28:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Matthews:2016:SMM

[MSF16] Eric Matthews, Lesley Shannon, and Alexandra Fedorova. Shared memory multicore MicroBlaze system with SMP Linux support. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):26:1–26:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Montone:2010:PFD

[MSSM10] Alessio Montone, Marco D. Santambrogio, Donatella Sciuto, and Seda Ogrenci Memik. Placement and floorplanning in dynamically reconfigurable FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):24:1–24:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Mentens:2022:ISS

[MST22] Nele Mentens, Lionel Sousa, and Pedro Trancoso. Introduction to the special section on FPL 2020. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):35:1–35:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3536336.

Miller:2015:GBA

[MVGB15] Bailey Miller, Frank Vahid, Tony Givargis, and Philip Brisk. Graph-based approaches to placement of processing element networks on FPGAs for physical model simulation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4): 10:1–10:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Mohajer:2021:PUC

[MWBL21] Soheil Mohajer, Zhiheng Wang, Kia Bazargan, and Yuyang Li. Parallel unary computing based on function derivatives. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(1):4:1–4:25, January 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3418464.

Martin:2012:CPA

[MWK^{+}12] Kevin Martin, Christophe Wolinski, Krzysztof Kuchcinski, Antoine Floch, and François Charot. Constraint programming approach to reconfigurable processor extension generation and application compilation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(2):10:1–10:??, June 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Murray:2015:TDT

[MWL^{+}15] Kevin E. Murray, Scott Whitty, Suya Liu, Jason Luu, and Vaughn Betz. Timing-driven Titan: Enabling large benchmarks and exploring the gap between academic and commercial CAD. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):10:1–10:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Miyagi:2025:SAL

[MYST25] Ryota Miyagi, Ryota Yasudo, Kentaro Sano, and Hideki Takase. A scalable accelerator for local score computation of structure learning in Bayesian networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):8:1–8:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3674842.

Mu:2020:OOB

[MZLS20] Jiandong Mu, Wei Zhang, Hao Liang, and Sharad Sinha. Optimizing OpenCL-based CNN design on FPGA with comprehensive design space exploration and collaborative performance modeling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(3):13:1–13:28, September 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3397514.

Neely:2013:RTH

[NBS13] Christopher E. Neely, Gordon Brebner, and Weijia Shang. Re-Shape: Towards a high-level approach to design and operation of modular reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(1):5:1–5:??, May 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Niu:2015:AEI

[NCJ^{+}15] Xinyu Niu, Thomas C. P. Chau, Qiwei Jin, Wayne Luk, Qiang Liu, and Oliver Pell. Automating elimination of idle functions by runtime reconfiguration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(3):15:1–15:??, May 2015. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Noyez:2024:MMS

[NEPV24] Louis Noyez, Nadia El Mrabet, Olivier Potin, and Pascal Veron. Montgomery multiplication scalable systolic designs optimized for DSP48E2. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):9:1–9:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3624571.

Nickel:2025:SAH

[NG25] Matthias Nickel and Diana Göhringer. A survey on architectures, hardware acceleration and challenges for in-network computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):10:1–10:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3699514.

Nechi:2023:FBD

[NGM^{+}23] Anouar Nechi, Lukas Groth, Saleh Mulhem, Farhad Merchant, Rainer Buchty, and Mladen Berekovic. FPGA-based deep learning inference accelerators: Where are we standing? ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (4):60:1–60:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3613963.

Nikolic:2024:EFS

[NI24] Stefan Nikolić and Paolo Ienne. Exploring FPGA switch-blocks without explicitly listing connectivity patterns. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (1):14:1–14:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3597417.

Niu:2014:SAT

[NJLW14] Xinyu Niu, Qiwei Jin, Wayne Luk, and Stephen Weston. A selfaware tuning and self-aware evaluation method for finite-difference applications in reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2):15:1–15:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Nabina:2012:AVS

[NNY12] Atukem Nabina and Jose Luis Nunez-Yanez. Adaptive voltage scaling in a dynamically reconfigurable FPGA-based platform. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(4):20:1–20:??, December 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Nava:2011:ADR

[NSS^{+}11] Federico Nava, Donatella Sciuto, Marco Domenico Santambrogio, Stefan Herbrechtsmeier, Mario Porrmann, Ulf Witkowski, and Ulrich Rueckert. Applying dynamic reconfiguration in the mobile robotics domain: a case study on computer vision algorithms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):29:1–29:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Nakajima:2011:FOR

[NW11] Mao Nakajima and Minoru Watanabe. Fast optical reconfiguration of a nine-context DORGA using a speed adjustment control. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):15:1–15:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Nikolic:2022:DPD

[NZI22] Stefan Nikolić, Grace Zgheib, and Paolo Ienne. Detailed placement for dedicated LUT-level FPGA interconnect. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (4):37:1–37:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3501802.

Nayak:2023:IEE

[NZS^{+}23] Ankita Nayak, Keyi Zhang, Rajsekhar Setaluri, Alex Carsello, Makai Mann, Christopher Torng, Stephen Richardson, Rick Bahr, Pat Hanrahan, Mark Horowitz, and Priyanka Raina. Improving energy efficiency of CGRAs with low-overhead fine-grained power domains. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):26:1–26:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3558394.

Ould-Bachir:2013:SAS

[OBD13] Tarek Ould-Bachir and Jean Pierre David. Self-alignment schemes for the implementation of addition-related floating-point operators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(1):1:1–1:??, May 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Owaida:2019:DID

[OKA19] Muhsen Owaida, Amit Kulkarni, and Gustavo Alonso. Distributed inference over decision tree ensembles on clusters of FP-GAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(4):17:1–17:??, October 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3340263.

Olivares:2012:RAV

[Oli12] Joaquín Olivares. Reconfigurable architecture for VBSME with variable pixel precision. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(1):3:1–3:??, March 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Oppermann:2019:EPM

[OROS^{+}19] Julian Oppermann, Melanie Reuter-Oppermann, Lukas Sommer, Andreas Koch, and Oliver Sinnen. Exact and practical modulo scheduling for high-level synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(2):8:1–8:??, June 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3317670.

Ost:2012:EAT

[OVI^{+}12] Luciano Ost, Sameer Varyani, Leandro Soares Indrusiak, Marcelo Mandelli, Gabriel Marchesan Almeida, Eduardo Wachter, Fernando Moraes, and Gilles Sassatelli. Enabling adaptive techniques in heterogeneous MPSoCs based on virtualization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3): 17:1–17:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

ONeill:2011:SPM

[OWMZ11] Shane O’Neill, Roger Francis Woods, Alan James Marshall, and Qi Zhang. A scalable and programmable modular traffic manager architecture. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):14:1–14:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Parandeh-Afshar:2009:FLC

[PABI09] Hadi Parandeh-Afshar, Philip Brisk, and Paolo Ienne. An FPGA logic cell and carry chain configurable as a 6:2 or 7:2 compressor. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(3):19:1–19:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Parandeh-Afshar:2011:CTS

[PANBI11] Hadi Parandeh-Afshar, Arkosnato Neogy, Philip Brisk, and Paolo Ienne. Compressor tree synthesis on commercial high-performance FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):39:1–39:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Petelin:2018:WEF

[PB18] Oleg Petelin and Vaughn Betz. Wotan: Evaluating FPGA architecture routability without benchmarks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(2):11:1– 11:??, November 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Prost-Boucle:2018:HEC

[PBBP18] Adrien Prost-Boucle, Alban Bourge, and Frédéric Pétrot. Highefficiency convolutional ternary neural networks with custom adder trees and weight compression. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(3):15:1–15:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Prost-Boucle:2017:EVF

[PBPLA17] Adrien Prost-Boucle, Frédéric Pétrot, Vincent Leroy, and Hande Alemdar. Efficient and versatile FPGA acceleration of support counting for stream mining of sequences and frequent itemsets. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3):21:1–21:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ploumidis:2025:EPE

[PCC^{+}25] Manolis Ploumidis, Fabien Chaix, Nikolaos Chrysos, Marios Assiminakis, Nikolaos Kallimanis, Nikolaos Kossifidis, Michael Nikoloudakis, Nikolaos Dimou, Michalis Gianioudis, George Ieronymakis, Aggelos Ioannou, George Kalokerinos, Pantelis Xirouchakis, Astrinos Damianakis, Michael Ligerakis, Theocharis Vavouris, Manolis Katevenis, Vassilis Papaefstathiou, Manolis Marazakis, and Iakovos Mavroidis. The ExaNeSt prototype: Evaluation of efficient HPC communication hardware in an ARMbased multi-FPGA rack. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):24:1–24:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Proulx:2023:SFC

[PCFM23] Alexandre Proulx, Jean-Yves Chouinard, Paul Fortier, and Amine Miled. A survey on FPGA cybersecurity design strategies. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):20:1–20:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3561515.

Park:2015:PIC

[PD15] Joonseok Park and Pedro C. Diniz. Program-invariant checking for soft-error detection using reconfigurable hardware. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1): 1:1–1:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Papadimitriou:2011:PPR

[PDH11] Kyprianos Papadimitriou, Apostolos Dollas, and Scott Hauck. Performance of partial reconfiguration in FPGA systems: a survey and a cost model. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(4):36:1–36:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Patterson:2009:STP

[PEM^{+}09] C. D. Patterson, S. W. Ellingson, B. S. Martin, K. Deshpande, J. H. Simonetti, M. Kavic, and S. E. Cutchin. Searching for transient pulses with the ETA radio telescope. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4): 20:1–20:??, January 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Paulino:2015:RAB

[PFC15] Nuno Paulino, João Canas Ferreira, and João M. P. Cardoso. A reconfigurable architecture for binary acceleration of loops with memory accesses. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):2:1–2:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Panchapakesan:2022:SEA

[PFL22] Sathish Panchapakesan, Zhenman Fang, and Jian Li. SyncNN: Evaluating and accelerating spiking neural networks on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):48:1–48:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3514253.

Panerati:2014:CIL

[PMC^{+}14] Jacopo Panerati, Martina Maggio, Matteo Carminati, Filippo Sironi, Marco Triverio, and Marco D. Santambrogio. Coordination of independent loops in self-adaptive systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(2): 12:1–12:??, June 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Papaphilippou:2022:HHP

[PMGL22] Philippos Papaphilippou, Jiuxi Meng, Nadeen Gebara, and Wayne Luk. Hipernetch: High-performance FPGA network switch. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):3:1–3:31, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3477054.

Parvez:2011:ASF

[PMKM11] Husain Parvez, Zied Marrakchi, Alp Kilic, and Habib Mehrez. Application-specific FPGA using heterogeneous logic blocks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):24:1–24:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Papadopoulos:2010:TRM

[PP10] Konstantinos Papadopoulos and Ioannis Papaefstathiou. Titan-R: a multigigabit reconfigurable combined compression/decompression unit. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(2):7:1–7:??, May 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Purnaprajna:2010:RRM

[PPR^{+}10] Madhura Purnaprajna, Mario Porrmann, Ulrich Rueckert, Michael Hussmann, Michael Thies, and Uwe Kastens. Runtime reconfiguration of multiprocessors based on compile-time analysis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(3):17:1–17:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Peetermans:2021:DAC

[PRV21] Adriaan Peetermans, Vladimir Rozić, and Ingrid Verbauwhede. Design and analysis of configurable ring oscillators for true random number generation based on coherent sampling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(2):7:1–7:20, July 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3433166.

Peng:2014:BAH

[PSM^{+}14] Yuanxi Peng, Manuel Saldaña, Christopher A. Madill, Xiaofeng Zou, and Paul Chow. Benefits of adding hardware support for broadcast and reduce operations in MPSoC applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):17:1–17:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Pellauer:2009:PNP

[PVA^{+}09] Michael Pellauer, Muralidaran Vijayaraghavan, Michael Adler, Arvind, and Joel Emer. A-port networks: Preserving the timed behavior of synchronous systems for modeling on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(3):16:1–16:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Plavec:2013:ETD

[PVB13] Franjo Plavec, Zvonko Vranesic, and Stephen Brown. Exploiting task- and data-level parallelism in streaming applications implemented in FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(4):16:1–16:??, December 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Pang:2016:MKR

[PWP^{+}16] Yeyong Pang, Shaojun Wang, Yu Peng, Xiyuan Peng, Nicholas J. Fraser, and Philip H. W. Leong. A microcoded kernel recursive least squares processor using FPGA technology. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1): 5:1–5:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Qiu:2024:FFD

[QMG^{+}24] Yunhui Qiu, Yiqing Mao, Xuchen Gao, Sichao Chen, Jiangnan Li, Wenbo Yin, and Lingli Wang. FDRA: a framework for a dynamically reconfigurable accelerator supporting multi-level parallelism. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):4:1–4:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3614224.

Que:2023:RRM

[QNF^{+}23] Zhiqiang Que, Hiroki Nakahara, Hongxiang Fan, He Li, Jiuxi Meng, Kuen Hung Tsoi, Xinyu Niu, Eriko Nurvitadhi, and Wayne Luk. Remarn: a reconfigurable multi-threaded multi-core accelerator for recurrent neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):4:1–4:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3534969.

Quinn:2015:CFE

[QRDC^{+}15] Heather Quinn, Diane Roussel-Dupre, Mike Caffrey, Paul Graham, Michael Wirthlin, Keith Morgan, Anthony Salazar, Tony Nelson, Will Howes, Eric Johnson, Jon Johnson, Brian Pratt, Nathan Rollins, and Jim Krone. The Cibola Flight Experiment. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(1):3:1–3:??, February 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Rodionov:2016:FGI

[RBR16] Alex Rodionov, David Biancolin, and Jonathan Rose. Finegrained interconnect synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):31:1–31:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Roldao:2010:HTF

[RC10] Antonio Roldao and George A. Constantinides. A high throughput FPGA-based floating point conjugate gradient implementation for dense matrices. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(1):1:1–1:??, January 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Rubin:2011:CYO

[RD11] Raphael Rubin and André Dehon. Choose-your-own-adventure routing: Lightweight load-time defect avoidance. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4 (4):33:1–33:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Rossi:2018:PPR

[RDB^{+}18] Enrico Rossi, Marvin Damschen, Lars Bauer, Giorgio Buttazzo, and Jörg Henkel. Preemption of the partial reconfiguration process to enable real-time computing with FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(2):10:1– 10:??, November 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Reggiani:2021:ESM

[RDC^{+}21] Enrico Reggiani, Emanuele Del Sozzo, Davide Conficconi, Giuseppe Natale, Carlo Moroni, and Marco D. Santambrogio. Enhancing the scalability of multi-FPGA stencil computations via highly optimized HDL components. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(3):15:1–15:33, September 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3461478.

Richardson:2016:AFR

[RGCL16] Justin Richardson, Alan George, Kevin Cheng, and Herman Lam. Analysis of fixed, reconfigurable, and hybrid devices with computational, memory, I/O, & realizable-utilization metrics. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):2:1–2:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Reardon:2010:SFR

[RGGW10] Casey Reardon, Eric Grobelny, Alan D. George, and Gongyu Wang. A simulation framework for rapid analysis of reconfigurable computing systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):25:1–25:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Rouhani:2018:RRT

[RHLK18] Bita Darvish Rouhani, Siam Umar Hussain, Kristin Lauter, and Farinaz Koushanfar. ReDCrypt: Real-time privacy-preserving deep learning inference in clouds using FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11 (3):21:1–21:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Raut:2023:EAE

[RKV23] Gopal Raut, Saurabh Karkun, and Santosh Kumar Vishvakarma. An empirical approach to enhance performance for scalable CORDIC-based deep neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):39:1–39:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3596220.

Riebler:2017:EBB

[RLM^{+}17] Heinrich Riebler, Michael Lass, Robert Mittendorf, Thomas Löcke, and Christian Plessl. Efficient branch and bound on FPGAs using work stealing and instance-specific designs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3): 24:1–24:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ren:2015:EFT

[RLY^{+}15] Yu Ren, Leibo Liu, Shouyi Yin, Jie Han, and Shaojun Wei. Efficient fault-tolerant topology reconfiguration using a maximum flow algorithm. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(3):19:1–19:??, May 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Rouhani:2016:ART

[RMSK16] Bita Darvish Rouhani, Azalia Mirhoseini, Ebrahim M. Songhori, and Farinaz Koushanfar. Automated real-time analysis of streaming big and dense data on reconfigurable platforms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1): 8:1–8:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Rybalkin:2022:WMG

[RNTW22] Vladimir Rybalkin, Jonas Ney, Menbere Kina Tekleyohannes, and Norbert Wehn. When massive GPU parallelism ain’t enough: a novel hardware architecture of 2D-LSTM neural network. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):2:1–2:35, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3469661.

Roorda:2022:FAE

[RRLW22] Esther Roorda, Seyedramin Rasoulinezhad, Philip H. W. Leong, and Steven J. E. Wilton. FPGA architecture exploration for DNN acceleration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):33:1–33:37, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3503465.

Rasoulinezhad:2022:REB

[RRW^{+}22] Seyedramin Rasoulinezhad, Esther Roorda, Steve Wilton, Philip H. W. Leong, and David Boland. Rethinking embedded blocks for machine learning applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):9:1–9:30, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3491234.

Rupnow:2011:SAD

[RUC11] Kyle Rupnow, Keith D. Underwood, and Katherine Compton. Scientific application demands on a reconfigurable functional unit interface. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):19:1–19:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Raitza:2016:RRN

[RVHP16] Michael Raitza, Markus Vogt, Christian Hochberger, and Thilo Pionteck. RAW 2014: Random number generators on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(2):15:1–15:??, February 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Reis:2024:DDL

[RVN24] Miguel Reis, Mário Véstias, and Horácio Neto. Designing deep learning models on FPGA with multiple heterogeneous engines.

ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):6:1–6:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3615870.

Sghaier:2010:IAT

[SAD10] Ahmad Sghaier, Shawki Areibi, and Robert Dony. Implementation approaches trade-offs for WiMax OFDM functions on reconfigurable platforms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(3):12:1–12:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sivaswamy:2008:SAP

[SB08] Satish Sivaswamy and Kia Bazargan. Statistical analysis and process variation-aware routing and skew assignment for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(1):4:1–4:??, March 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Scicluna:2015:AMF

[SB15] Neil Scicluna and Christos-Savvas Bouganis. ARC 2014: a multidimensional FPGA-based parallel DBSCAN architecture. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1):2:1–2:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Singh:2025:CNN

[SB25] Gaurav Singh and Kia Bazargan. Compressing neural networks using learnable 1D non-linear functions. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):21:1–21:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Saiprasert:2010:OHA

[SBC10] Chalermpol Saiprasert, Christos-S. Bouganis, and George A. Constantinides. An optimized hardware architecture of a multivariate Gaussian random number generator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):2:1–2:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Shi:2015:IDD

[SBC15] Kan Shi, David Boland, and George A. Constantinides. Imprecise datapath design: an overclocking approach. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):6:1–6:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sedcole:2008:PYM

[SC08] Pete Sedcole and Peter Y. K. Cheung. Parametric yield modeling and simulations of FPGA circuits considering within-die delay variations. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(2):10:1–10:??, June 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Shannon:2011:LRH

[SC11] Lesley Shannon and Paul Chow. Leveraging reconfigurability in the hardware/software codesign process. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):28:1– 28:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Smith:2010:AFA

[SCC10] Alastair M. Smith, George A. Constantinides, and Peter Y. K. Cheung. An automated flow for arithmetic component generation in field-programmable gate arrays. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(3):13:1–13:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Soleimani:2024:PCA

[SCL24] Parastoo Soleimani, David W. Capson, and Kin Fun Li. A partitioned CAM architecture with FPGA acceleration for binary descriptor matching. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):10:1–10:??, March 2024. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3624749.

Shi:2023:DSE

[SCY^{+}23] Zhengyuan Shi, Cheng Chen, Gangqiang Yang, Hailiang Xiong, Fudong Li, Honggang Hu, and Zhiguo Wan. Design space exploration of Galois and Fibonacci configuration based on Espresso stream cipher. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):43:1–43:??, September 2023. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3567428.

Shield:2012:ACC

[SDG12] John Shield, Jean-Philippe Diguet, and Guy Gogniat. Asymmetric cache coherency: Policy modifications to improve multicore performance. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3):12:1–12:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Singh:2022:AWP

[SDGL^{+}22] Gagandeep Singh, Dionysios Diamantopoulos, Juan Gómez-Luna, Christoph Hagleitner, Sander Stuijk, Henk Corporaal, and Onur Mutlu. Accelerating weather prediction using near-memory reconfigurable fabric. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):39:1–39:??, December 2022. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3501804.

Stewart:2018:RPI

[SDM^{+}18] Robert Stewart, Kirsty Duncan, Greg Michaelson, Paulo Garcia, Deepayan Bhowmik, and Andrew Wallace. RIPL: a parallel image processing language for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):7:1–7:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Shahsavani:2023:ECM

[SFNP23] Soheil Nazar Shahsavani, Arash Fayyazi, Mahdi Nazemi, and Massoud Pedram. Efficient compilation and mapping of fixed function combinational logic onto digital signal processors targeting neural network inference and utilizing high-level synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):17:1–17:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3559543.

Soldavini:2023:ACH

[SFT^{+}23] Stephanie Soldavini, Karl Friebel, Mattia Tibaldi, Gerald Hempel, Jeronimo Castrillon, and Christian Pilato. Automatic creation of high-bandwidth memory architectures from domain-specific languages: The case of computational fluid dynamics. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (2):21:1–21:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3563553.

Sasdrich:2015:ICS

[SG15] Pascal Sasdrich and Tim Güneysu. Implementing Curve25519 for side-channel–protected elliptic curve cryptography. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1): 3:1–3:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sabogal:2021:RFR

[SGC21] Sebastian Sabogal, Alan George, and Gary Crum. Reconfigurable framework for resilient semantic segmentation for space applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(4):22:1–22:32, December 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3472770.

Sauvage:2009:ERF

[SGM09] Laurent Sauvage, Sylvain Guilley, and Yves Mathieu. Electromagnetic radiations of FPGAs: High spatial resolution cartography and attack on a cryptographic module. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(1):4:1–4:??, March 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Slogsnat:2008:OSH

[SGNB08] David Slogsnat, Alexander Giese, Mondrian Nüssle, and Ulrich Brüning. An open-source HyperTransport core. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1 (3):14:1–14:??, September 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sabogal:2020:RFE

[SGW20] Sebastian Sabogal, Alan George, and Christopher Wilson. Reconfigurable framework for environmentally adaptive resilience in hybrid space systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(3):14:1–14:32, September 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3398380.

Shannon:2022:ISS

[Sha22] Lesley Shannon. Introduction to special section on FPGA 2020. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):1:1–1:2, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3485586.

Schaumont:2009:GEI

[SJT09] Patrick R. Schaumont, Alex K. Jones, and Steve Trimberger. Guest Editors’ introduction to security in reconfigurable systems design. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(1):1:1–1:??, March 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Streit:2022:DET

[SKB^{+}22] Franz-Josef Streit, Paul Krüger, Andreas Becher, Stefan Wildermann, and Jürgen Teich. Design and evaluation of a tunable PUF architecture for FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):7:1–7:27, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3491237.

Shi:2022:EHF

[SKH^{+}22] Runbin Shi, Kaan Kara, Christoph Hagleitner, Dionysios Diamantopoulos, Dimitris Syrivelis, and Gustavo Alonso. Exploiting HBM on FPGAs for data processing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):36:1–36:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3491238.

Sun:2022:BEC

[SKJ22] Gongjin Sun, Seongyoung Kang, and Sang-Woo Jun. BurstZ+: Eliminating the communication bottleneck of scientific computing accelerators via accelerated compression. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (2):21:1–21:34, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3476831.

Sasongko:2021:HCS

[SKW^{+}21] Arif Sasongko, I. M. Narendra Kumara, Arief Wicaksana, Frédéric Rousseau, and Olivier Muller. Hardware context switchbased cryptographic accelerator for handling multiple streams.

ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(3):14:1–14:25, September 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3460941.

Sinnen:2023:ISS

[SLD23] Oliver Sinnen, Qiang Liu, and Azadeh Davoodi. Introduction to special section on FPT’20. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):1:1–1:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3579850.

Sano:2010:FAB

[SLH^{+}10] Kentaro Sano, Wang Luzhou, Yoshiaki Hatsuda, Takanori Iizuka, and Satoru Yamamoto. FPGA-array with bandwidth-reduction mechanism for scalable and power-efficient numerical simulations based on finite difference methods. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):21:1–21:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

She:2025:SLP

[SLH^{+}25] Yuhan She, Jierui Liu, Yanlong Huang, Ray C. C. Cheung, and Hong Yan. A speculative loop pipeline framework with accurate path modeling for high-level synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):17:1–17:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Shao:2020:PGF

[SLL^{+}20] Zhiyuan Shao, Chenhao Liu, Ruoshi Li, Xiaofei Liao, and Hai Jin. Processing grid-format real-world graphs on DRAM-based FPGA accelerators with application-specific caching mechanisms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(3):11:1–11:33, September 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3391920.

Sugiura:2025:FAC

[SM25] Keisuke Sugiura and Hiroki Matsutani. FPGA-accelerated correspondence-free point cloud registration with PointNet features. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):25:1–25:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Singh:2024:ISI

[SML^{+}24] Satwant Singh, Carlos E. M. Marin, Yun (Eric) Liang, Yao Chen, Nele Mentens, and Raymond Nijssen. Introduction to the special issue on FPGA-based embedded systems for industrial and IoT applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):52:1–52:??, December 2024. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3698202.

Suh:2023:AHC

[SMN^{+}23] Han-Sok Suh, Jian Meng, Ty Nguyen, Vijay Kumar, Yu Cao, and Jae-Sun Seo. Algorithm–hardware co-optimization for energy-efficient drone detection on resource-constrained FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):33:1–33:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3583074.

Swierczynski:2015:PSE

[SMOP15] Pawel Swierczynski, Amir Moradi, David Oswald, and Christof Paar. Physical security evaluation of the bitstream encryption mechanism of Altera Stratix II and Stratix III FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):7:1–7:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Serre:2020:DBH

[SP20] François Serre and Markus Püschel. DSL-based hardware generation with Scala: Example Fast Fourier Transforms and sorting networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(1):1:1–1:23, February 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3359754.

Saldana:2010:MPM

[SPM^{+}10] Manuel Saldaña, Arun Patel, Christopher Madill, Daniel Nunes, Danyao Wang, Paul Chow, Ralph Wittig, Henry Styles, and Andrew Putnam. MPI as a programming model for high-performance reconfigurable computers. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):22:1–22:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Siozios:2012:NFE

[SPS12] Kostas Siozios, Vasilis F. Pavlidis, and Dimitrios Soudris. A novel framework for exploring 3-D FPGAs with heterogeneous interconnect fabric. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(1):4:1–4:??, March 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Skubich:2023:IRT

[SRR23] Christian Skubich, Peter Reichel, and Marc Reichenbach. Increasing the robustness of TERO-TRNGs against process variation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):48:1–48:??, September 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3597418.

Stitt:2016:PSW

[SSC16] Greg Stitt, Eric Schwartz, and Patrick Cooke. A parallel slidingwindow generator for high-performance digital-signal processing on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(3):23:1–23:??, July 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sidiropoulos:2013:JFS

[SSF^{+}13] Harry Sidiropoulos, Kostas Siozios, Peter Figuli, Dimitrios Soudris, Michael Hübner, and Jürgen Becker. JITPR: a framework for supporting fast application’s implementation onto FP-GAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2):7:1–7:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Schelten:2023:HTR

[SSK^{+}23] Niklas Schelten, Fritjof Steinert, Justin Knapheide, Anton Schulte, and Benno Stabernack. A high-throughput, resourceefficient implementation of the RoCEv2 remote DMA protocol and its application. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):5:1–5:??, March 2023. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3543176.

Sterpone:2010:NTD

[Ste10] Luca Sterpone. A new timing driven placement algorithm for dependable circuits on SRAM-based FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):7:1–7:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sahoo:2024:ADF

[SUK24] Siva Satyendra Sahoo, Salim Ullah, and Akash Kumar. AxOMaP: Designing FPGA-based approximate arithmetic operators using mathematical programming. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):31:1–31:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3648694.

Seetharaman:2009:ASF

[SV09] G. Seetharaman and B. Venkataramani. Automation schemes for FPGA implementation of wave-pipelined circuits. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2): 11:1–11:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sherwin:2022:MFF

[SWT^{+}22] Krystine Dawn Sherwin, Kevin I-Kai Wang, Prabu Thiagaraj, Ben Stappers, and Oliver Sinnen. Median filters on FPGAs for infinite data and large, rectangular windows. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):50:1–50:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3530273.

Sun:2025:FBL

[SXZ^{+}25a] Mingqian Sun, Guangwei Xie, Fan Zhang, Wei Guo, Xitian Fan, Li Chen, and Jiayu Du. FPGA-based large-scale sorting with optimized bandwidth utilization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):22:1–22:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Sun:2025:PRE

[SXZ^{+}25b] Mingqian Sun, Guangwei Xie, Fan Zhang, Wei Guo, Xitian Fan, Tianyang Li, Li Chen, and Jiayu Du. PTME: a regular expression matching engine based on speculation and enumerative computation on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):3:1–3:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3655626.

Sani:2024:EIU

[SY24] Sajjad Rostami Sani and Andy Ye. Evaluating the impact of using multiple-metal layers on the layout area of switch blocks for tile-based FPGAs in FinFET 7nm. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):17:1–17:??, March 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3639055.

Salamat:2022:NGN

[SZKR22] Sahand Salamat, Hui Zhang, Yang Seok Ki, and Tajana Rosing. NASCENT2: Generic near-storage sort accelerator for data analytics on SmartSSD. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):16:1–16:29, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3472769.

Shi:2023:OGR

[SZZW23] Kaichuang Shi, Xuegong Zhou, Hao Zhou, and Lingli Wang. An optimized GIB routing architecture with bent wires for FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):2:1–2:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

doi/10.1145/3519599.

Takano:2012:DAA

[Tak12] Shigeyuki Takano. Design and analysis of adaptive processor. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(1):5:1–5:??, March 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Takano:2017:PSA

[Tak17] Shigeyuki Takano. Performance scalability of adaptive processor architecture. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(2):16:1–16:??, April 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tian:2010:HPQ

[TB10] Xiang Tian and Khaled Benkrid. High-performance quasi-Monte Carlo financial simulation: FPGA vs. GPP vs. GPU. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3 (4):26:1–26:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tamimi:2025:DDA

[TBS^{+}25] Sajjad Tamimi, Arthur Bernhardt, Florian Stock, Ilia Petrov, and Andreas Koch. DANSEN: Database acceleration on native computational storage by exploiting NDP. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):4:1–4:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3655625.

Tarafdar:2022:AOF

[TDH^{+}22] Naif Tarafdar, Giuseppe Di Guglielmo, Philip C. Harris, Jeffrey D. Krupa, Vladimir Loncar, Dylan S. Rankin, Nhan Tran, Zhenbin Wu, Qianfeng Shen, and Paul Chow. AIgean: an open framework for deploying machine learning on heterogeneous clusters. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):23:1–23:32, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3482854.

Trochatos:2024:CCF

[TES24] Theodoros Trochatos, Anthony Etim, and Jakub Szefer. Covertchannels in FPGA-enabled SmartSSDs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):27:1–27:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3635312.

Thielmann:2012:MLH

[THK12] Benjamin Thielmann, Jens Huthmann, and Andreas Koch. Memory latency hiding by load value speculation for reconfigurable computers. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3):13:1–13:??, October 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Thomas:2015:THG

[Tho15] David B. Thomas. The table-Hadamard GRNG: an area-efficient FPGA Gaussian random number generator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(4):23:1– 23:??, October 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tavva:2025:CCB

[TJCP25] Yaswanth Tavva, Rohan Juneja, Trevor E. Carlson, and Li-Shiuan Peh. CTScan: a CGRA-based platform for the emulation of power side-channel attacks on edge CPUs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):26:1–26:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tang:2016:AKM

[TK16] Qing Y. Tang and Mohammed A. S. Khalid. Acceleration of kmeans algorithm using Altera SDK for OpenCL. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1): 6:1–6:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Trautmann:2024:DCE

[TKB^{+}24] Jens Trautmann, Paul Krüger, Andreas Becher, Stefan Wildermann, and Jürgen Teich. Design, calibration, and evaluation of real-time waveform matching on an FPGA-based digitizer at 10 GS/s. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):24:1–24:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3635719.

Tridgell:2019:UTN

[TKH^{+}19] Stephen Tridgell, Martin Kumm, Martin Hardieck, David Boland, Duncan Moss, Peter Zipf, and Philip H. W. Leong. Unrolling ternary neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(4):22:1–22:??, October 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Thomas:2008:MGR

[TL08] David B. Thomas and Wayne Luk. Multivariate Gaussian random number generation targeting reconfigurable hardware. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(2):12:1–12:??, June 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tai:2011:POA

[TL11] Tzu-Chiang Tai and Yen-Tai Lai. A performance-oriented algorithm with consideration on communication cost for dynamically reconfigurable FPGA partitioning. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):16:1–16:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tang:2024:GOH

[TLC^{+}24] Enhao Tang, Shun Li, Ruiqi Chen, Hao Zhou, Yuhanxiao Ma, Haoyang Zhang, Jun Yu, and Kun Wang. Graph-OPU: a highly flexible FPGA-Based overlay processor for graph neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):58:1–58:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3691636.

Taka:2021:PVA

[TMLS21] Endri Taka, Konstantinos Maragos, George Lentaris, and Dimitrios Soudris. Process variability analysis in interconnect, logic, and arithmetic blocks of 16-nm FinFET FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14 (3):13:1–13:30, September 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3458843.

Tili:2017:RPG

[TOS17] Ilian Tili, Kalin Ovtcharov, and J. Gregory Steffan. Reducing the performance gap between soft scalar CPUs and custom hardware with TILT. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3):22:1–22:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tokuda:2025:DBF

[TTY25] Daichi Tokuda and Shinya Takamaeda-Yamazaki. DF-BETA: an FPGA-based memory locality aware decision forest accelerator via bit-level early termination. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):20:1–20:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tang:2020:PSM

[TWG^{+}20] Qi Tang, Zhe Wang, Biao Guo, Li-Hua Zhu, and Ji-Bo Wei. Partitioning and scheduling with module merging on dynamic partial reconfigurable FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(3):16:1–16:24, September 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3403702.

Tao:2023:LGL

[TWL^{+}23] Zhuofu Tao, Chen Wu, Yuan Liang, Kun Wang, and Lei He. LW-GCN: a lightweight FPGA-based graph convolutional network accelerator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):10:1–10:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3550075.

Tatsumura:2018:EFM

[TYB18] Kosuke Tatsumura, Sadegh Yazdanshenas, and Vaughn Betz. Enhancing FPGAs with magnetic tunnel junction-based block RAMs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):6:1–6:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Tian:2023:SSA

[TYL^{+}23] Xingyu Tian, Zhifan Ye, Alec Lu, Licheng Guo, Yuze Chi, and Zhenman Fang. SASA: a scalable and automatic stencil acceleration framework for optimized hybrid spatial and temporal parallelism on HBM-based FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):28:1–28:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3572547.

Tan:2015:SHP

[TZWZ15] Guangming Tan, Chunming Zhang, Wendi Wang, and Peiheng Zhang. SuperDragon: a heterogeneous parallel system for accelerating 3D reconstruction of cryo-electron microscopy images. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(4):25:1–25:??, October 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ul-Abdin:2016:RCF

[UAS16] Zain Ul-Abdin and Bertil Svensson. A retargetable compilation framework for heterogeneous reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):24:1–24:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Umuroglu:2019:OBS

[UCR^{+}19] Yaman Umuroglu, Davide Conficconi, Lahiru Rasnayake, Thomas B. Preusser, and Magnus Själander. Optimizing bit-serial matrix multiplication for reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(3):15:1– 15:??, September 2019. CODEN ???? ISSN 1936-7406 (print),

1936-7414 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3337929.

Underwood:2009:SSL

[UHU09] Keith D. Underwood, K. Scott Hemmert, and Craig D. Ulmer. From silicon to science: The long road to production reconfigurable supercomputing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(4):26:1–26:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ulusel:2014:FDE

[UNBR14] Onur Ulusel, Kumud Nepal, R. Iris Bahar, and Sherief Reda. Fast design exploration for performance, power and accuracy tradeoffs in FPGA-Based accelerators. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(1):4:1–4:??, February 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ueno:2023:VVC

[US23] Tomohiro Ueno and Kentaro Sano. VCSN: Virtual circuitswitching network for flexible and simple-to-operate communication in HPC FPGA cluster. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(2):25:1–25:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3579848.

Ueno:2017:BCF

[USY17] Tomohiro Ueno, Kentaro Sano, and Satoru Yamamoto. Bandwidth compression of floating-point numerical data streams for FPGA-based high-performance computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3):18:1– 18:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Vestias:2023:EDL

[VDdSN23] Mário Véstias, Rui P. Duarte, José T. de Sousa, and Horácio Neto. Efficient design of low bitwidth convolutional neural networks on FPGA with optimized dot product units. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (1):13:1–13:??, March 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3546182.

Hirtum:2024:CND

[VDG^{+}24] Lennart Van Hirtum, Patrick De Causmaecker, Jens Goemaere, Tobias Kenter, Heinrich Riebler, Michael Lass, and Christian Plessl. A computation of the ninth Dedekind number using FPGA supercomputing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):40:1–40:??, September 2024. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3674147.

Voros:2014:ISI

[VG14] Nikolaos Voros and Guy Gogniat. Introduction to the special issue on the 7th International Workshop on Reconfigurable Communication-centric Systems-on-Chip (ReCoSoC’12). ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):23:1–23:??, August 2014. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Vaidya:2011:NMC

[VL11] Pranav Vaidya and Jaehwan John Lee. A novel multicontext coarse-grained reconfigurable architecture (CGRA) for accelerating column-oriented databases. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(2):13:1–13:??, May 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Vliegen:2015:SRD

[VMV15] Jo Vliegen, Nele Mentens, and Ingrid Verbauwhede. Secure, remote, dynamic reconfiguration of FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(4):8:1–8:??, January 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Vaishnav:2020:FMF

[VPPK20] Anuj Vaishnav, Khoa Dang Pham, Joseph Powell, and Dirk Koch. FOS: a modular FPGA operating system for dynamic workloads. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(4):20:1–20:28, October 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3405794.

Vassiliadis:2009:ADF

[VTN09] Nikolaos Vassiliadis, George Theodoridis, and Spiridon Nikolaidis. An application development framework for ARISE reconfigurable processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(4):24:1–24:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Vakili:2025:DFL

[VVZL25] Shervin Vakili, Mobin Vaziri, Amirhossein Zarei, and J. M. Pierre Langlois. DyRecMul: Fast and low-cost approximate multiplier for FPGAs using dynamic reconfiguration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):7:1–7:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3663480.

Wang:2023:LSL

[WAS^{+}23] Erwei Wang, Marie Auffret, Georgios-Ilias Stavrou, Peter Y. K. Cheung, George A. Constantinides, Mohamed S. Abdelfattah, and James J. Davis. Logic shrinkage: Learned connectivity sparsification for LUT-based neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):57:1–57:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3583075.

Woods:2015:PDP

[WAT15] Louis Woods, Gustavo Alonso, and Jens Teubner. Parallelizing data processing on FPGAs with shifter lists. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(2):7:1– 7:??, April 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Woods:2010:GEA

[WBAM10] Roger Woods, Jürgen Becker, Peter Athanas, and Fearghal Morgan. Guest editorial ARC 2009. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(1):1:1–1:??, December 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Winterstein:2016:SLH

[WBC16] Felix J. Winterstein, Samuel R. Bayliss, and George A. Constantinides. Separation logic for high-level synthesis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(2): 10:1–10:??, February 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wilson:2024:IFT

[WBCW24] Andrew Elbert Wilson, Nathan Baker, Ethan Campbell, and Michael Wirthlin. Improving fault tolerance for FPGA SoCs through post-radiation design analysis. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):50:1–50:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3674841.

Wong:2016:MCM

[WBR16] Henry Wong, Vaughn Betz, and Jonathan Rose. Microarchitecture and circuits for a 200 MHz out-of-order soft processor memory system. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):7:1–7:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wong:2018:HPI

[WBR18] Henry Wong, Vaughn Betz, and Jonathan Rose. Highperformance instruction scheduling circuits for superscalar outof-order soft processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):1:1–1:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wijtvliet:2021:CER

[WCK21] Mark Wijtvliet, Henk Corporaal, and Akash Kumar. CGRA-EAM-Rapid energy and area estimation for coarse-grained reconfigurable architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(4):19:1–19:28, December 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3468874.

Wulf:2016:FEO

[WGGR16] Nicholas Wulf, Alan D. George, and Ann Gordon-Ross. A framework for evaluating and optimizing FPGA-based SoCs for aerospace computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):1:1–1:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wulf:2017:OFP

[WGGR17] Nicholas Wulf, Alan D. George, and Ann Gordon-Ross. Optimizing FPGA performance, power, and dependability with linear programming. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3):23:1–23:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wu:2023:TEA

[WHJ^{+}23] Guiming Wu, Qianwen He, Jiali Jiang, Zhenxiang Zhang, Yuan Zhao, Yinchao Zou, Jie Zhang, Changzheng Wei, Ying Yan, and Hui Zhang. Topgun: an ECC accelerator for private set intersection. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):52:1–52:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3603114.

Wilton:2008:SDO

[WHQ^{+}08] Steven J. E. Wilton, Chun Hok Ho, Bradley Quinton, Philip H. W. Leong, and Wayne Luk. A synthesizable datapath-oriented embedded FPGA fabric for silicon debug applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1 (1):7:1–7:??, March 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wang:2010:VVP

[WL10] Xiaojun Wang and Miriam Leeser. VFloat: a variable precision fixed- and floating-point library for reconfigurable hardware. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(3):16:1–16:??, September 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Williams:2010:CFR

[WMG^{+}10] Jason Williams, Chris Massie, Alan D. George, Justin Richardson, Kunal Gosrani, and Herman Lam. Characterization of fixed and reconfigurable multi-core devices for application acceleration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):19:1–19:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Weng:2024:TAS

[WML^{+}24] Olivia Weng, Gabriel Marcano, Vladimir Loncar, Alireza Khodamoradi, Abarajithan G., Nojan Sheybani, Andres Meza, Farinaz Koushanfar, Kristof Denolf, Javier Mauricio Duarte, and Ryan Kastner. Tailor: Altering skip connections for resourceefficient inference. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(1):11:1–11:??, March 2024. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3624990.

Wijesundera:2018:FRP

[WPSI18] Deshya Wijesundera, Alok Prakash, Thambipillai Srikanthan, and Achintha Ihalage. Framework for rapid performance estimation of embedded soft core processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(2):9:1–9:??, November 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wilson:2016:UAA

[WS16] David Wilson and Greg Stitt. The unified accumulator architecture: a configurable, portable, and extensible floating-point accumulator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(3):21:1–21:??, July 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wong:2009:SMC

[WSC09] Justin S. J. Wong, Pete Sedcole, and Peter Y. K. Cheung. Self-measurement of combinatorial circuit delays in FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2):10:1–10:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wolf:2023:ASE

[WSDH23] Dennis Leander Wolf, Christoph Spang, Daniel Diener, and Christian Hochberger. Advantages of a statistical estimation approach for clock frequency estimation of heterogeneous and irregular CGRAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(1):7:1–7:??, March 2023. CODEN ????

ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3531062.

Wu:2025:TPM

[WSZ^{+}25] Jiajun Wu, Mo Song, Jingmin Zhao, Yizhao Gao, Jia Li, and Hayden Kwok-Hay So. TATAA: Programmable mixed-precision transformer acceleration with a transformable arithmetic architecture. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(1):14:1–14:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/3714416.

Wang:2019:FBA

[WTS19] Haomiao Wang, Prabu Thiagaraj, and Oliver Sinnen. FPGAbased acceleration of FT convolution for pulsar search using OpenCL. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(4):24:1–24:??, January 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3268933.

Wu:2022:LPF

[WWC^{+}22] Chen Wu, Mingyu Wang, Xinyuan Chu, Kun Wang, and Lei He. Low-precision floating-point arithmetic for high-performance FPGA-based CNN acceleration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):6:1–6:21, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3474597.

Wegley:2016:ASD

[WYZ16] Evan Wegley, Yanhua Yi, and Qinhai Zhang. Application of specific delay window routing for timing optimization in FPGA designs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):29:1–29:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Wong:2024:DDF

[WZL24] Linus Y. Wong, Jialiang Zhang, and Jing Li. DONGLE 2.0: Direct FPGA-orchestrated NVMe storage for HLS. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (3):45:1–45:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3650038.

Wu:2024:FAI

[WZY^{+}24] Zi-Ming Wu, Meng-Yuan Zhao, Bin Yan, Jeng-Shyang Pan, and Hong-Mei Yang. FPGA accelerated implementation of 3D mesh secret sharing based on symmetric similarity of model. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(4):60:1–60:??, December 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3689049.

Xu:2009:FAR

[XCG^{+}09] Ning-Yi Xu, Xiong-Fei Cai, Rui Gao, Lei Zhang, and Feng-Hsiung Hsu. FPGA acceleration of RankBoost in Web search engines. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4):19:1–19:??, January 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Xu:2016:CGA

[XJD^{+}16] Jinwei Xu, Jingfei Jiang, Yong Dou, Xiaolong Shen, and Zhiqiang Liu. Coarse-grained architecture for fingerprint matching. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(2):12:1–12:??, February 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Xu:2024:ESA

[XJXQ24] Shiyao Xu, Jingfei Jiang, Jinwei Xu, and Xifu Qian. Efficient SpMM accelerator for deep learning: Sparkle and its automated generator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(3):38:1–38:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3665896.

Xiao:2024:EEH

[XPN^{+}24] Yuanlong Xiao, Dongjoon Park, Zeyu Jason Niu, Aditya Hota, and André Dehon. ExHiPR: Extended high-level partial reconfiguration for fast incremental FPGA compilation. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (2):21:1–21:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3617837.

Xu:2023:FAG

[XRC23] Tiancheng Xu, Scott Rixner, and Alan L. Cox. An FPGA accelerator for genome variant calling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(4):53:1–53:??, December 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3595297.

Yazdanshenas:2018:CAM

[YB18] Sadegh Yazdanshenas and Vaughn Betz. COFFE 2: Automatic modelling and optimization of complex and heterogeneous FPGA architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(1):3:1–3:??, April 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3301298.

Yu:2016:OAH

[YBS16] Ting Yu, Chris Bradley, and Oliver Sinnen. ODoST: Automatic hardware acceleration for biomedical model integration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):27:1–27:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Yasudo:2021:APE

[YCV^{+}21] Ryota Yasudo, José G. F. Coutinho, Ana-Lucia Varbanescu, Wayne Luk, Hideharu Amano, Tobias Becker, and Ce Guo. Analytical performance estimation for large-scale reconfigurable dataflow platforms. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(3):12:1–12:21, September 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3452742.

Yu:2009:VPS

[YEC^{+}09] Jason Yu, Christopher Eagleston, Christopher Han-Yu Chou, Maxime Perreault, and Guy Lemieux. Vector processing as a soft processor accelerator. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2):12:1–12:??, June 2009. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Yang:2017:FSA

[YFW^{+}17] Hsin-Jung Yang, Kermin Fleming, Felix Winterstein, Michael Adler, and Joel Emer. (FPL 2015) Scavenger: Automating the construction of application-optimized memory hierarchies. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(2):13:1–13:??, April 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Yu:2018:IDC

[YGH^{+}18] Jincheng Yu, Guangjun Ge, Yiming Hu, Xuefei Ning, Jiantao Qiu, Kaiyuan Guo, Yu Wang, and Huazhong Yang. Instruction driven cross-layer CNN accelerator for fast detection on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(3):22:1–22:??, December 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Yang:2021:BWB

[YHK^{+}21] Tao Yang, Zhezhi He, Tengchuan Kou, Qingzheng Li, Qi Han, Haibao Yu, Fangxin Liu, Yun Liang, and Li Jiang. BISWSRBS: a Winograd-based CNN accelerator with a fine-grained regular sparsity pattern and mixed precision quantization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14 (4):18:1–18:28, December 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3467476.

Yoo:2010:IRR

[YKBS10] Sang-Kyung Yoo, Deniz Karakoyunlu, Berk Birand, and Berk Sunar. Improving the robustness of ring oscillator TRNGs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(2):9:1–9:??, May 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Yang:2024:HQO

[YLF^{+}24] Geng Yang, Jie Lei, Zhenman Fang, Yunsong Li, Jiaqing Zhang, and Weiying Xie. HyBNN: Quantifying and optimizing hardware efficiency of binary neural networks. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17(2):25:1–25:??, June 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414

(electronic). URL https://dl.acm.org/doi/10.1145/3631610.

Yoshimi:2017:PPJ

[YOY17] Masato Yoshimi, Yasin Oge, and Tsutomu Yoshinaga. Pipelined parallel join and its FPGA-based acceleration. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(4):28:1– 28:??, December 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Yang:2023:HOF

[YSC^{+}23] Gangqiang Yang, Zhengyuan Shi, Cheng Chen, Hailiang Xiong, Fudong Li, Honggang Hu, and Zhiguo Wan. Hardware optimizations of Fruit-80 stream cipher: Smaller than grain. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (2):22:1–22:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3569455.

Yan:2011:FBA

[YXC^{+}11] Jing Yan, Ning-Yi Xu, Xiong-Fei Cai, Rui Gao, Yu Wang, Rong Luo, and Feng-Hsiung Hsu. An FPGA-based accelerator for LambdaRank in Web search engines. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):25:1–25:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Ziener:2016:FBD

[ZBB^{+}16] Daniel Ziener, Florian Bauer, Andreas Becher, Christopher Dennl, Klaus Meyer-Wegener, Ute Schürfeld, Jürgen Teich, Jörg-Stephan Vogt, and Helmut Weber. FPGA-based dynamically reconfigurable SQL query processing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(4):25:1–25:??, September 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhao:2009:TMB

[ZBC^{+}09] Weisheng Zhao, Eric Belhaire, Claude Chappert, Bernard Dieny, and Guillaume Prenat. TAS-MRAM-based low-power high-speed runtime reconfiguration (RTR) FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(2):8:1–8:??, June 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhang:2012:PSF

[ZBR12] Wei Zhang, Vaughn Betz, and Jonathan Rose. Portable and scalable FPGA-based acceleration of a direct linear system solver. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(1):6:1–6:??, March 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhang:2022:RFH

[ZCK22] Niansong Zhang, Xiang Chen, and Nachiket Kapre. RapidLayout: Fast hard block placement of FPGA-optimized systolic arrays using evolutionary algorithm. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(4):38:1–38:??, December 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3501803.

Zhang:2015:EAR

[ZCL15] Jianfeng Zhang, Paul Chow, and Hengzhu Liu. An enhanced adaptive recoding rotation CORDIC. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1):4:1–4:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhang:2016:CBE

[ZCL16] Jianfeng Zhang, Paul Chow, and Hengzhu Liu. CORDICbased enhanced systolic array architecture for QR decomposition. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(2):9:1–9:??, February 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zeng:2022:UFV

[ZDS^{+}22] Shulin Zeng, Guohao Dai, Hanbo Sun, Jun Liu, Shiyao Li, Guangjun Ge, Kai Zhong, Kaiyuan Guo, Yu Wang, and Huazhong Yang. A unified FPGA virtualization framework for generalpurpose deep neural networks in the cloud. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15 (3):24:1–24:31, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3480170.

Zhao:2025:HDP

[ZFCZ25] Chenfeng Zhao, Clayton Faber, Roger Chamberlain, and Xuan Zhang. HLPerf: Demystifying the performance of HLS-based graph neural networks with dataflow architectures. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18 (1):2:1–2:??, March 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3655627.

Zaidi:2016:VSF

[ZG16] Ali Mustafa Zaidi and David Greaves. Value state flow graph: a dataflow compiler IR for accelerating control-intensive code in spatial hardware. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(2):14:1–14:??, February 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zick:2012:LCS

[ZH12] Kenneth M. Zick and John P. Hayes. Low-cost sensing with ring oscillator arrays for healthier reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5 (1):1:1–1:??, March 2012. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhou:2021:SHC

[ZHL^{+}21] Zhen Zhou, Debiao He, Zhe Liu, Min Luo, and Kim-Kwang Raymond Choo. A software/hardware co-design of Crystals-Dilithium signature scheme. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(2):11:1–11:21, July 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/3447812.

Zhuang:2024:CCH

[ZLY^{+}24] Jinming Zhuang, Jason Lau, Hanchen Ye, Zhuoping Yang, Shixin Ji, Jack Lo, Kristof Denolf, Stephen Neuendorffer, Alex Jones, Jingtong Hu, Yiyu Shi, Deming Chen, Jason Cong, and Peipei Zhou. CHARM 2.0: Composing heterogeneous accelerators for deep learning on versal ACAP architecture. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 17 (3):51:1–51:??, September 2024. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3686163.

Zhao:2023:ASC

[ZMH^{+}23] Kang Zhao, Yuchun Ma, Ruining He, Jixing Zhang, Ning Xu, and Jinian Bian. Adaptive selection and clustering of partial reconfiguration modules for modern FPGA design flow. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16 (2):27:1–27:??, June 2023. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/doi/

10.1145/3567427.

Zhou:2022:ROS

[ZML^{+}22] Yun Zhou, Pongstorn Maidee, Chris Lavin, Alireza Kaviani, and Dirk Stroobandt. RWRoute: an open-source timing-driven router for commercial FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(1):8:1–8:27, March 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3491236.

Zhao:2018:FGM

[ZNA^{+}18] Zhuoran Zhao, Nguyen T. H. Nguyen, Dimitris Agiakatsikas, Ganghee Lee, Ediz Cetin, and Oliver Diessel. Fine-grained module-based error recovery in FPGA-based TMR systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):4:1–4:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhang:2019:RAD

[ZQ19] Jiliang Zhang and Gang Qu. Recent attacks and defenses on FPGA-based systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(3):14:1–14:??, September 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3340557.

Zhou:2025:FBB

[ZQBL25] Wenjie Zhou, Haoyan Qi, David Boland, and Philip H. W. Leong. FPGA-based block minifloat training accelerator for a time series prediction network. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):16:1–16:??, June 2025. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhang:2021:CHP

[ZSP^{+}21] Xuzhi Zhang, Xiaozhe Shao, George Provelengios, Naveen Kumar Dumpala, Lixin Gao, and Russell Tessier. CoNFV: a heterogeneous platform for scalable network function virtualization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(1):1:1–1:29, January 2021. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3409113.

Zhou:2020:AFR

[ZVS20] Yun Zhou, Dries Vercruyce, and Dirk Stroobandt. Accelerating FPGA routing through algorithmic enhancements and connectionaware parallelization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(4):18:1–18:26, October 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3406959.

Zuo:2025:ALD

[ZW25] Cheng Zuo and Chang Wu. Algorithmic-level design partitioning for latency minimization in multi-chip and multi-die systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 18(2):30:1–30:??, June 2025. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhou:2019:FAN

[ZWM19] Xuegong Zhou, Lingli Wang, and Alan Mishchenko. Fast adjustable NPN classification using generalized symmetries. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 12(2):7:1–7:??, June 2019. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3313917.

Zhang:2020:MRB

[ZZB^{+}20] Jialiang Zhang, Yue Zha, Nicholas Beckwith, Bangya Liu, and Jing Li. MEG: a RISCV-based system emulation infrastructure for near-data processing using FPGAs and high-bandwidth memory. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(4):19:1–19:24, October 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://dl.acm.

org/doi/10.1145/3409114.

Zhang:2013:FBA

[ZZJB13] Yan Zhang, Fan Zhang, Zheming Jin, and Jason D. Bakos. An FPGA-Based accelerator for frequent itemset mining. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(1):2:1–2:??, May 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).